### v1

# The Coverage Principle: A Framework for Understanding Compositional Generalization

Abstract
Large language models excel at pattern matching, yet often fall short in systematic compositional generalization. We propose the coverage principle: a data-centric framework showing that models relying primarily on pattern matching for compositional tasks cannot reliably generalize beyond substituting fragments that yield identical results when used in the same contexts. We demonstrate that this framework has a strong predictive power for the generalization capabilities of Transformers. First, we derive and empirically confirm that the training data required for two-hop generalization grows at least quadratically with the token set size, and the training data efficiency does not improve with 20x parameter scaling. Second, for compositional tasks with path ambiguity where one variable affects the output through multiple computational paths, we show that Transformers learn context-dependent state representations that undermine both performance and interpretability. Third, Chain-of-Thought supervision improves training data efficiency for multi-hop tasks but still struggles with path ambiguity. Finally, we outline a mechanism-based taxonomy that distinguishes three ways neural networks can generalize: structure-based (bounded by coverage), property-based (leveraging algebraic invariances), and shared-operator (through function reuse). This conceptual lens contextualizes our results and highlights where new architectural ideas are needed to achieve systematic compositionally. Overall, the coverage principle provides a unified lens for understanding compositional reasoning, and underscores the need for fundamental architectural or training innovations to achieve truly systematic compositionality. [^1] All code and resources are available at: https://github.com/kaistAI/coverage-principle
1 Introduction
[…]
Figure 1: Left: In a two-hop task $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, two fragments $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ satisfying $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ consistently yield the same final output when combined with the same context $x_{3}$, supporting their functional equivalence. Right: Among all possible inputs (few shown), we draw an edge between any two inputs that differ only by functionally equivalent fragments to form a substitution graph. Then, coverage is the set of observed inputs (highlighted as blue) and all inputs connected to them. The coverage principle states that a learner relying on pattern matching can reliably generalize only to inputs that are reachable from the observations by substituting functionally equivalent fragments.
Human generalization capability is often attributed to our ability to manipulate symbols compositionally [1, 2]. Despite the remarkable performance of Large Language Models (LLMs) [3, 4, 5, 6, 7, 8], our understanding of how they represent and manipulate compositional structures is still limited. Meanwhile, a growing body of work points to the brittleness of LLMs in performing systematic [^2] We refer to systematicity as an ability to generalize to unseen outcomes by exploiting their structural relations. For a more thorough discussion, see Fodor and Pylyshyn [9], Marcus [10], and Szabó [11]. compositional reasoning [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These observations suggest that LLMs succeed primarily through pattern matching, i.e., by exploiting surface-level statistical correlations between input fragments and outputs, rather than systematic compositional reasoning. However, prior analyses of the pattern-matching behavior remain largely observational and task-specific, leaving us without a unified framework to predict when they will succeed or fail.
We show that Transformers’ [26] generalization on synthetic compositional tasks is predictable through a unified framework grounded in one observation: Transformers generalize by utilizing functionally equivalent fragments supported by input-output relations. This insight allows us to move beyond observing success or failure cases to systematically predicting generalization boundaries. To this end, we first formalize the concept of functional equivalence, where two fragments (i.e., components of input sequences) are regarded as equivalent when they yield identical results in all shared contexts observed in the training data. Then, we introduce coverage, the set of inputs reachable from training data by substituting functionally equivalent fragments. Finally, we propose the coverage principle, which states that coverage is a necessary condition for compositional generalization, i.e., model predictions become unreliable beyond this fundamental boundary, as long as the only generalization mechanism is pattern matching.
From this conceptual formulation, we derive specific predictions about generalization in Transformers. First, we derive and empirically confirm that reliable generalization on a two-hop reasoning task requires training data to scale at least quadratically with the token set size. Remarkably, our experiments suggest that this scaling relationship remains invariant even for 20x larger model, underscoring that the limitation is inherent to data properties rather than model capacity.
Second, our framework also predicts specific complications in tasks with path ambiguities, where the same variable influences outputs through multiple computational paths. We show that Transformers cannot form unified representations of identical intermediate states in such cases, instead developing context-dependent state representations that fragment across different paths. This degrades both performance and interpretability, explaining why seemingly simple compositional tasks can remain challenging even for large models.
Third, while Chain-of-Thought (CoT) supervision [27, 28] improves data efficiency by ‘flattening’ tasks into sequences of single-hop problems, our analysis reveals a fundamental limitation: CoT-trained models still cannot resolve path ambiguities without showing nearly all possible in-domain combinations. This finding provides insights into understanding the Transformers’ failure modes in complex reasoning tasks with path ambiguities (e.g., planning [29, 30, 31]).
We place these findings within a broader, mechanism-based taxonomy that distinguishes three routes to generalization: (i) structure-based generalization via functional equivalence (the focus of this paper), (ii) property-based generalization that exploits algebraic invariances of individual primitive functions, and (iii) shared-operator generalization that reuses the same computation across positions. This taxonomy clarifies how certain model behaviours that occur outside the coverage region can arise from Type-II or Type-III mechanisms, providing a conceptual framework for analysing diverse sources of generalization.
Overall, the coverage principle offers a unified conceptual framework providing valuable insights into compositional generalization, highlighting how compositional reasoning and its mechanistic interpretation remain inherently bounded by coverage limitations. This underscores the need for architectural or training innovations to achieve truly systematic generalization in neural networks.
2 Related work
Compositional generalization and systematicity
Since Fodor and Pylyshyn [9] posed the systematicity challenge, arguing connectionist systems inherently lack symbolic structural capabilities, understanding how neural networks might achieve compositional reasoning has remained a persistent research question [32, 33, 34, 35, 36, 37, 38, 39, 10]. Central to this debate is the variable binding problem, how neural networks can flexibly associate variables with values across contexts [40, 41, 42], which Marcus [37] argues as the fundamental limitation of modern connectionist approaches. Despite AI advances, a growing body of work reveals limitations in neural networks’ compositional capabilities [14, 43, 18, 44, 45, 20, 46, 15, 21, 22, 23, 24, 47, 42]. Dziri et al. [17] observe that Transformers perform linearized subgraph matching to solve compositional tasks, but leave open when and why it occurs. Our coverage principle addresses this gap by formally demarcating what pattern matching models can achieve without true variable binding, showing both the possibilities and fundamental limitations of pattern matching-based approaches through data-based preconditions. This aligns with empirical findings of Andreas [48] and Auersperger and Pecina [49], where compositional data augmentation that implicitly expands coverage significantly improves generalization. Our derived power-law scaling relationship between token set size and required dataset size contributes to scaling laws literature [50, 51], but differs by characterizing a data requirement invariant to model size for compositional tasks. Unlike prior disentanglement work [52, 53, 54, 55] which focuss on latent factors, our data-centric view identifies fundamental boundary conditions for pattern-based generalization. The coverage principle thus complements recent architectural or training innovations [56, 57, 58, 59, 60] by showing what any pattern matching approach must overcome to achieve true systematicity.
Mechanistic interpretability
Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors [61, 62, 63, 64]. Recent work traces how Transformer components coordinate on compositional tasks [65, 66, 67, 68, 69]. Wang et al. [19] demonstrate that in-domain compositional generalization emerges through grokking, but do not identify necessary data conditions for such emergence. Our data-centric framework complements this optimization-centric view by pinpointing when generalizing circuits can exist: only when training data provides sufficient evidence of functional equivalence, essentially formalizing when the pattern matching mechanisms that Fodor and Pylyshyn [9] critique can succeed or must fail. Our findings also explain why standard interpretability techniques like logit lens [70, 71] may fail to identify compositional structure in models trained on tasks with path ambiguities.
3 A formulation of the coverage principle
We now develop a formal framework of the coverage principle. We first provide an intuitive illustration with a two-hop compositional task, then generalize the intuitions to formalize the coverage principle for more general compositional tasks.
Imagine a learner [^3] By “learner”, we mean any system that learns from data (e.g., neural networks). observing input-output pairs determined by a function $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$. The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is sampled from a finite set ${\mathcal{X}}$ [^4] For brevity, we use a shared token set ${\mathcal{X}}$. Position-specific domains ${\mathcal{X}}_{i}$ can be embedded as subsets of an enlarged token set $\tilde{{\mathcal{X}}}={\mathcal{X}}_{1}\cup{\mathcal{X}}_{2}\cup{\mathcal{X}}_{3}$ without loss of generality. Being unknown to the learner, let $f$ be a function that can be factorized as the composition of two primitive functions as $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig. 2(a). How can the learner generalize by only seeing the input-output pairs?
Our key intuition is that a learner exploits the compositional structure only when observations show that two fragments behave identically. For instance, assume that $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ for $(x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{X}}^{2}$. From the underlying structure, these fragments are functionally equivalent, i.e., substituting one for the other never changes the output. If observations consistently support their equivalence, i.e., $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$ for observed $x_{3}$ values, this equivalence can be supported (Fig. 1 (Left)). Intuitively, the learner would be able to harness this equivalence to generalize to an unseen input $(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime\prime})$, provided the learner has seen $f(x_{1},x_{2},x_{3}^{\prime\prime})$.
In short, the learner can utilize the observed functional equivalence to correctly infer the output of an unseen input, if it can reach an observed input by safe substitutions supported by observations (Fig. 1 (Right)). Coverage is a set of such inputs that are reachable from an observed input through chains of functionally equivalent substitutions. The essence of our framework, the coverage principle, is that a learner can only generalize inside the coverage when the only source of compositional generalization is the observation and utilization of functional equivalence.
We now formalize these concepts for an arbitrary fixed-length task with an arbitrary set of observations. We restrict our attention to single-token prediction tasks defined as a deterministic mapping $f:{\mathcal{X}}^{n}\to{\mathcal{X}}$, where ${\mathcal{X}}$ is a finite set of tokens. We also consider a fixed observation set $D\subset{\mathcal{X}}^{n}$, a collection of inputs that are allowed to be observed by the learner. Write ${\bm{x}}=(x_{1},\dots,x_{n})\in{\mathcal{X}}^{n}$ and, for a subset $I\subset[n]:=\{1,\dots,n\}$, let ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ be a subsequence of ${\bm{x}}$. The first step is to formalize what it means for two subsequences to be functionally equivalent.
Definition 3.1 (Functional $k$-equivalence).
Fix a nonempty proper subset of indices $I\subsetneq[n]$. Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{|I|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset D$ to be an $I$-co-occurrence of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ if it satisfies ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[n]\setminus I}={\bm{x}}^{\prime}_{[n]\setminus I}$. Also, we say ${\bm{a}}$ and ${\bm{a}}^{\prime}$ are functionally $k$-equivalent at $I$ in $D$, denoted by ${\bm{a}}\equiv^{I}_{D}{\bm{a}}^{\prime}$, if it satisfies:
  1. 1.
    (Sufficiency of co-occurrences.) There are $k$ or more distinct $I$-co-occurrences of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$;
  2. 2.
    (Consistency.) Every $I$-co-occurrence $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ satisfies $f({\bm{x}})=f({\bm{x}}^{\prime})$.
In other words, two subsequences are functionally $k$-equivalent if they behave identically in the same contexts at least $k$ times. The hyperparameter $k$ represents the strength of evidence required to establish functional equivalence between two subsequences. The minimum value $k=1$ corresponds to the weakest form of evidence, meaning a single shared context is sufficient to establish equivalence, whereas higher values of $k$ demand more robust evidence.
Next, we ask: which inputs are reachable from observed data utilizing functional equivalence? To formalize this, we define the substitution graph: Let ${\mathcal{G}}_{D,k}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{n}$ of all possible inputs. Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[n]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$. With this substitution graph ${\mathcal{G}}_{D,k}$, we formally define the $k$-coverage as a set of inputs which are connected [^5] For an undirected graph ${\mathcal{G}}$, two vertices $u,v$ are connected if ${\mathcal{G}}$ contains a path between $u$ and $v$. to at least one observed input as follows:
Definition 3.2 ($k$-coverage).
The $k$-coverage of $D$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in ${\mathcal{X}}^{n}$ that is connected to an input in $D$ on the substitution graph ${\mathcal{G}}_{D,k}$.
Note that the notion of coverage is a stricter condition of the conventional definition of in-domain, which is obtained by random train/test split [19] or taking combinations of observed internal computations [17]. We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms. Now, we formally phrase the coverage principle:
Claim 3.3 (Coverage principle).
Let a learner output $\hat{f}$ that is consistent with the training set $D$. If generalization is derived solely from observed functional $k$-equivalences and ${\bm{x}}\notin\mathrm{Cover}_{k}(D)$, the value $\hat{f}({\bm{x}})$ is unconstrained by $D$.
This establishes a fundamental boundary: no learner can reliably generalize beyond the coverage as long as the only source of the generalization is functional equivalence. In Sec. 5.1, we show that Transformers indeed fail to generalize beyond coverage when our random mapping dataset construction isolates functional equivalence as the sole generalization source. However, deeper insights come from asking when generalization success occurs in practice. While theory suggests that deep learning methods implicitly favor simpler solutions [72, 73, 74], which may involve leveraging functional equivalence, we still lack a systematic picture of how task structure, dataset, and model size interact to determine when such solutions emerge. We study this through experiments in Sec. 5.
4 Experimental setup
$x_{1}$$x_{2}$$f_{1}$$b$$x_{3}$$f_{2}$$t$
(a) 2-Hop
$x_{1}$$x_{2}$$x_{3}$$x_{4}$$f_{1}$$f_{2}$$b_{1}$$b_{2}$$f_{3}$$t$
(b) Parallel 2-Hop
$x_{1}$$x_{2}$$f_{1}$$b_{1}$$x_{3}$$f_{2}$$b_{2}$$x_{4}$$f_{3}$$t$
(c) 3-Hop
$x_{1}$$x_{2}$$f_{1}$$b$$x_{3}$$f_{2}$$t$
(d) Non-Tree
Figure 2: Four synthetic compositional structures we study.
Dataset construction
We construct four synthetic compositional tasks: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig. 2). We create random mappings from the product space of a token set to control generalization sources not attributable to compositional structures (i.e., commutativity). [^6] Such external generalization sources will be discussed in Sec. 7 and App. H. We explain the dataset construction process using 2-Hop task (Fig. 2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example. We construct training datasets by defining a token set with size $|{\mathcal{X}}|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset. See Sec. B.1 for more details of the dataset construction process.
Training & evaluation
Following Wang et al. [19], we train randomly initialized GPT-2 [75] models with 8 layers, 12 heads, and 768 dimensions (see Sec. B.2 for details). We construct two evaluation sets, each with 2,000 instances: (1) In-Domain (ID) Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{2})$ in 2-Hop task) are observed during training, but the specific combination was unseen. (2) Out-of-Domain (OOD) Test Set: at least one primitive function application is never observed during training.
5 Results
5.1 Predictive power of $k$-coverage on Transformer’s compositional generalization
Figure 3: Left: Percentage of covered ID data depending on $k$ values and dataset size ($N$), for 2-Hop task ($|{\mathcal{X}}|=50$). Right: Test accuracy depending on $k$-cutoff values for 2-Hop task ($|{\mathcal{X}}|$=50, $N$=10k). Each line represents a different training checkpoint. Note that OOD accuracy remains at chance level (=1/50) regardless of training time. The bars below show the number of test data for each $k$-cutoff value.
We now empirically validate the predictive power of the coverage principle by quantifying the correlation between $k$-coverage and model generalization performance. To this end, we implement and release a task-agnostic coverage determination algorithm (see App. C for implementation details) that can be applied to diverse compositional structures. Then, we analyze what fraction of ID test data of 2-Hop task with $|{\mathcal{X}}|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$. Fig. 3 (Left) shows that at $N=5$k, every ID test example is already covered with minimal evidence ($k=1$).
Ideally, a single witness of functional equivalence could suffice. However, experiments show that such minimal coverage alone is insufficient for efficient generalization in practice. To demonstrate this, we define a sample’s $k$-cutoff as the lowest $k$ for which an input lies in $k$-coverage, measuring the strength of evidence for functional equivalence. For example, a $k$-cutoff of 3 means that an example is inside coverage with $k=3$ but not with $k=4$. For 2-Hop dataset with $N=10$k, we classify each ID test instance according to its $k$-cutoff, and track accuracy development for each group across 50k training epochs. As shown in Fig. 3 (Right), generalization success shows a strong relationship with $k$-cutoff values. Test data with low $k$-cutoff values show delayed improvement even after extensive training, while examples with stronger evidence generalize much faster. Notably, OOD test examples (outside coverage) remain at chance-level ($1/50=0.02$) throughout training, empirically validating our coverage principle.
These results yield two important insights. First, successful generalization in practice requires robust coverage so the model can confidently identify and utilize functional equivalence relationships. The $k$ parameter effectively quantifies this evidence strength, directly impacting generalization speed and reliability. Second, while our experiments use uniformly distributed data, the coverage principle explains why models struggle with generalizing long-tail distribution in imbalanced real-world data [76, 77, 78]. Rare combinations naturally receive limited functional equivalence evidence (low $k$), placing them effectively outside practical coverage, despite technically being in-distribution.
5.2 The clustering of latent representation drives generalization on coverage
Figure 4: Left: Heatmap of Intra-Inter Cosine Gap (IICG) across layers and positions, sliced by $k$-cutoff. Higher IICG values indicate stronger clustering of representations that share the same intermediate state. The positions with the highest IICG values are marked with squares. Right: PCA visualization of latent representations at position $x_{2}$ and layer 3. Datapoints are classified by their intermediate states $b=f_{1}(x_{1},x_{2})$.
Next, we investigate how the model internally represents functional equivalence for $k$-covered inputs. Specifically, we inspect a GPT-2 trained on 2-Hop task ($|{\mathcal{X}}|=50$, $N=10\text{k}$) for 50k epochs (corresponding to the yellow line in Fig. 3 (Right)). [^7] Analyses for varying factors including task structures, entity set size ($|{\mathcal{X}}|$), dataset size ($N$), and training steps give consistent results; see App. D. We observe that when a model successfully generalizes to ID test data, it maps functionally equivalent components into tight latent clusters, thereby encoding the equivalence relationships needed for compositional generalization.
To quantify this representation clustering phenomenon, we develop a metric that captures how distinctly the model separates functionally equivalent fragments from others. Specifically, we measure the difference between the average pairwise cosine similarity of latent vectors that share the same intermediate state $b=f_{1}(x_{1},x_{2})$ ($\overline{\cos}_{\text{intra}}$), and those that do not ($\overline{\cos}_{\text{inter}}$), for each position and layer of the model. We term this difference the Intra–Inter Cosine Gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation. Fig. 4 (Left) reveals a clear relationship: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations. Critically, OOD examples show no clustering pattern at all, since they lack any functional equivalence evidence in the training data, as predicted by the coverage principle. The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually. We verify that the representation clusters play a causal role in ID generalization with causal tracing [67, 66], a widely used mechanistic interpretability technique to identify Transformer circuits (see Fig. 8 in App. D).
Our findings extend the insights of Wang et al. [19] in several ways. First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps. Consequently, out-of-coverage inputs lack these unified circuits by construction, explaining their systematic failure on OOD examples. Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens [70] may fail to detect these functional equivalence representations despite their presence. See App. I for a detailed analysis.
5.3 When does coverage lead to full ID generalization? A power-law lower bound
Figure 5: Left: Log-log plot of measured $\hat{N}_{\mathrm{req}}$ vs. token set size ($|{\mathcal{X}}|$) across three compositional tasks. The slope $c$ corresponds to the empirical power-law scaling exponent. Omitted points for 3-Hop are due to prohibitively large dataset requirements. Right: Power-law scaling remains invariant across GPT-2 model sizes (68M to 1.5B parameters) for 2-Hop task. $R^{2}>0.99$ for all linear fitting.
Our analysis of $k$-coverage and representation clustering demonstrates that stronger functional equivalence evidence leads to better generalization. A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data? Intuitively, this requires the training set to support the functional equivalence of every pair of inputs that shares the same intermediate state $b$. Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$. Assuming a learner requires at least $k$ distinct pairs of evidence to establish functional equivalence of two fragments, how does the required dataset size scale with the token set size $|{\mathcal{X}}|$? In practical terms, this question addresses how much data is fundamentally needed for compositional generalization, which is central to understanding data scaling requirements for compositional tasks. For 2-Hop task, we derive the following result (full derivation in App. E):
Result 5.1 (Power-law lower bound).
Let $f:{\mathcal{X}}^{3}\!\to\!{\mathcal{X}}$ be a 2-Hop composition. Assume a learner recognizes functional equivalence of two subsequences only after observing them in at least $k$ distinct pairs of evidence (i.e., functionally $k$-equivalent). [^8] Note that this assumption is equivalent to modeling Fig. 3 (Right) as a step function. Let $N_{\text{req}}(|{\mathcal{X}}|,k)$ be the smallest training dataset size that enables complete generalization to ID examples under this evidence threshold. Then, up to poly-logarithmic factors in $|{\mathcal{X}}|$, $N_{\text{req}}(|{\mathcal{X}}|,k)=\tilde{\Omega}\bigl{(}|{\mathcal{X}}|^{\,\alpha(k)}\bigr{)},\text{where}\;\alpha(k)=2.5-\frac{0.5}{k}.$
Res. 5.1 predicts that a learner relying on pattern matching requires training dataset size scaling at least quadratically with respect to the token set size, to fully generalize on ID test data. To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data. [^9] See App. F measurement details. Fig. 5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs. $|{\mathcal{X}}|$ across different task structures. The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions of at least quadratic scaling. Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well. The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
Strikingly, these exponents remain invariant across three different GPT-2 model sizes spanning a 20x range in parameters (from 68M to 1.5B) for all three tasks (Fig. 5 (Right) and Tab. 2 in App. F). This supports that the scaling relationship is primarily determined by data properties rather than model capacity, reinforcing our framework’s data-centric perspective. The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App. F). Overall, these results suggest that compositional generalization in Transformers requires dataset sizes scaling as a power-law with token set size. This has significant implications for how we approach complex reasoning tasks, suggesting that data curation strategies informed by coverage considerations may be more effective than raw parameter scaling.
5.4 Path ambiguity hinders generalization and interpretability in Transformers
Identical Intermediate StatesFunctional Equivalence$(x_{1},x_{2})$$(x_{1}^{\prime},x_{2})$$(x_{1},x_{2}^{\prime})$$(x_{1}^{\prime},x_{2}^{\prime})$$(x_{1},x_{2}^{\prime\prime})$$(x_{1}^{\prime},x_{2}^{\prime\prime})$
Figure 6: Left: In Non-tree task, the coverage principle predicts that representations of input subsequences with the same intermediate state $b=f_{1}(x_{1},x_{2})$ split into multiple context-dependent state representations, conditioned on $x_{2}$ value. Middle: ID test accuracy after standard training (100 epochs post training accuracy$>$0.99), showing Non-tree task significantly underperforms 2-Hop task ($|{\mathcal{X}}|=50$). Right: IICG heatmap from a model that achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|{\mathcal{X}}|=50$, $N=50\text{k}$).
Many real-world reasoning tasks involve computational structures where a single variable affects the output through multiple paths, creating what we call path ambiguity. Our framework predicts that in such cases, Transformers struggle with forming unified representations of intermediate states that are theoretically equivalent, instead developing context-dependent state representations that vary based on the input context. In this section, we analyze Non-tree task (Fig. 2(d)) as a case study, where $x_{2}$ affects the output through two paths: as input to $f_{1}$ and directly to $f_{2}$. Following the definition of functional equivalence (Def. 3.1), this path ambiguity prevents the model from establishing the functional equivalence of two subsequences $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ that produce the same intermediate state $b$, unless they also share the same $x_{2}$ value ($x_{2}=x_{2}^{\prime}$). Consequently, the model creates context-dependent state representations rather than unifying them to represent true intermediate state $b$, as illustrated in Fig. 6 (Left).
In consequence, the path ambiguity hinders both generalization performance and efficiency, as the model must establish functional equivalence for each $x_{2}$-conditioned equivalent pair. Fig. 6 (Middle) shows that GPT-2 can fully generalize on ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data. [^10] For $|\mathcal{X}|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times|\mathcal{X}|^{3}\;\approx\;61\text{k}$ distinct ID triples). Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig. 17 in App. G). Extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96. However, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig. 6 (Right)). In contrast, grouping by $x_{2}$-conditioned intermediate state ($(b,x_{2})$) leads to high IICG scores, suggesting the formation of context-dependent state representations.
This context-dependence raises an interpretability concern, as standard linear probing-based techniques like logit lens [70, 71] would not reliably identify intermediate states. Moreover, our analysis provides insights into why LLMs fall short in planning tasks [29, 31, 30], as planning often requires correct tracking of intermediate states, which can influence outcomes through multiple paths.
5.5 CoT supervision improves data efficiency but still struggles with path ambiguity
Figure 7: Left: Power-law scaling of required dataset size vs. token set size for tasks with CoT supervision. $R^{2}>0.98$ for all linear fits. Middle: Comparison of ID test Accuracy of Non-tree task ($|{\mathcal{X}}|=50$) with and without CoT supervision (100 epochs post training accuracy$>$0.99). Right: IICG score comparison for Non-tree and 2-Hop task with CoT supervision ($|\mathcal{X}|=50$, $N=10\text{k}$). The scores are measured at each layer of intermediate state position $b$, based on two grouping strategies: $b$ and $(b,x_{2})$. Models are trained for 100 epochs after reaching training accuracy$>$0.99.
CoT supervision [27, 28] dramatically improves performance on multi-step reasoning tasks. We investigate how CoT interacts with the coverage principle and whether it can address the challenges observed in Sections 5.3 and 5.4. Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example. This explicit supervision substantially improves data efficiency (Fig. 7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in 3-Hop task, aligning with previous studies on the sample efficiency of CoT [79, 80, 81]. Remarkably, the scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision. We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
However, we find models trained with CoT supervision still struggle with the path ambiguity in Non-tree task. Despite showing improvements with CoT supervision, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig. 7 (Middle). IICG analysis (Right) reveals that the model’s representations remain partially context-dependent. For 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve. In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model. We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$. Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec. 5.4. Our analysis likely explains why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data [82].
6 A taxonomy for understanding generalization mechanisms
Our coverage analysis reveals fundamental limits of pattern matching for compositional generalization. This raises a broader question: What other mechanisms allow neural networks to go beyond coverage boundaries? We sketch a preliminary taxonomy that distinguishes three complementary mechanisms and offers hypotheses about why models sometimes succeed beyond coverage [^11] The taxonomy is not exhaustive; Appendix H details limitations and open questions..
Type-I: Structure-based generalization occurs when models identify and utilize functionally equivalent components based on how primitive functions are composed. This is precisely what we formalize through the coverage principle: models learn that different input fragments yield identical results in shared contexts, enabling generalization to new fragment combinations. Crucially, this generalization remains bounded by coverage, and reliable generalization fails without sufficient functional equivalence evidence. Type-I describes the ceiling of pattern matching without explicit variable binding.
Type-II: Function property-based generalization exploits intrinsic properties of individual primitive functions (e.g., algebraic invariances such as commutativity), group-theoretic structure in modular arithmetic [63], or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present [81]). Unlike Type-I, this can transcend coverage limitations by leveraging properties that hold across all inputs of a primitive. The Reversal Curse exemplifies the layered nature of compositional challenges across multiple generalization types. The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. Property-based approaches like bidirectional training partially succeed [83] by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data. However, recent evidence [42] shows that models still struggle once entities switch syntactic roles, indicating a remaining challenge in variable binding. [^12] See [42] for controlled concept–surface dissociation experiments.
Type-III: Shared-operator generalization emerges through reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$). Recurrent architectures [84] exemplify this through weight sharing across time steps, enabling processing of variable-length sequences [85]. In Transformers, inductive biases towards reuse of the same computation through parameter sharing [86, 87, 19] can improve generalization on compositional tasks beyond coverage. We interpret this mechanism as exploiting structural repetition without requiring Type-II’s algebraic property understanding.
Distinguishing mechanisms from phenomena
Prior categorizations focus on observed phenomena, for example, “systematic” versus “mix-and-match” [12]. We instead categorize the underlying mechanisms, as our coverage principle offers quantitative predictions for Type-I boundaries while identifying when Type-II or Type-III mechanisms become necessary. Many influential studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure. We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.
Implications and future directions
Real compositional tasks typically involve combinations of all three types, making it difficult to isolate why models succeed or fail. We suggest diagnostic approaches based on the proposed taxonomy: coverage ablation tests for Type-I, algebraic identity probes for Type-II, and parameter-sharing ablations for Type-III. While preliminary, this taxonomy provides a conceptual scaffold for understanding the fundamental capabilities and limitations of neural networks on compositional reasoning. This work isolates and formalizes structure-based generalization (Type-I) to clarify its specific boundaries. When models succeed beyond our coverage predictions, e.g., through ignoring distractors without seeing all combinations or leveraging parameter sharing, we view these as exploiting Type-II or Type-III mechanisms rather than contradicting our framework. Our focus on Type-I suggests that fundamental systematicity challenges remain as long as models rely primarily on pattern matching, requiring architectural innovations that harness all three mechanisms and, crucially, incorporate explicit variable binding. We hope this preliminary taxonomy serves as a conversation starter, and confirming or refuting its utility is an empirical matter that we invite the community to explore.
7 Discussion
Systematicity without variable binding
The coverage principle pinpoints a fundamental question in the systematicity debate: How far can a neural network generalize compositionally when it relies only on pattern matching and has no mechanism for explicit variable binding? By formalizing the exact conditions under which pattern matching succeeds, we draw a principled boundary around what can be learned from input–output pairs alone. This boundary clarifies the limitations that Fodor and Pylyshyn [9] and Marcus [37] highlighted decades ago. Our results therefore complement approaches that encode compositional structure directly, whether through symbolic interfaces, slot-based representations, or neural modules [56, 57, 58, 59, 60, 88], and delineate what must still be overcome to reach truly systematic generalization.
Practical implications
Our framework provides valuable insights into various observed LLM behaviors. First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization [43]. Next, our framework explains why models struggle with generalizing long-tail knowledge [76, 77, 78], as low-frequency combinations naturally receive limited functional equivalence evidence. Similarly, failures in complex planning tasks [30, 82, 31, 89], even with CoT techniques, might be attributed to path ambiguities. Moreover, our coverage principle predicts the reversal curse phenomenon [16, 42, 83], as pattern-matching models fundamentally cannot generalize to reversed relationships without explicit functional equivalence evidence in training data.
Beyond explaining failures, our framework helps understand when and why standard interpretability techniques like linear probing or logit lens can fail. In addition, our analysis suggests targeted data augmentation strategies that maximize coverage by ensuring diverse shared contexts for functionally equivalent components. This can explain the success of strategic data augmentation methods [90, 91, 49, 48] and offers principled guidance. Finally, although our empirical study focuses on GPT-2–style Transformers, the coverage principle itself makes no architectural assumptions, and it applies to any learning system whose generalization mechanism is primarily pattern matching. Extending the analysis to recurrent (e.g., LSTM [84]), state-space [92], or convolutional [93] architectures, and more realistic data where multiple computational structures coexist within the same dataset [12, 15, 94] remains promising future work.
8 Conclusion
We introduce the coverage principle, a data-centric framework that specifies when pattern-matching learners can and cannot generalize compositionally. Our theoretical analysis and controlled experiments show that a Transformer’s success is tightly predicted by coverage: two-hop tasks obey a superquadratic data scaling law that does not improve with 20x parameter scaling, path ambiguity fragments internal representations, and CoT supervision helps only inside the coverage boundary.
These results isolate Type I structure-based generalization and reveal its limits. Our taxonomy then situates two additional mechanisms, Type II property exploitation and Type III shared-operator reuse, that help explain coverage “violations” reported in the literature. Yet all three mechanisms may still rely on sophisticated pattern matching rather than explicit variable binding, leaving the fundamental systematicity challenge unresolved.
Hence, the systematicity challenge posed by Fodor and Pylyshyn [9] and by Marcus [37] remains open. Matching human-like compositionality will likely require architectures that can bind and manipulate symbols independently of surface form, not merely scale up or refine current pattern-matching strategies. We hope the coverage principle and the accompanying taxonomy provide a clear target and a roadmap for the next stage of innovation.
Acknowledgments and Disclosure of Funding
We would like to thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
Appendix A Limitations
We deliberately restrict to synthetic tasks to isolate structure-based limits without confounds from lexical or domain priors. We leave extending the coverage analysis to natural data as a future work. Additionally, our experiments focus on autoregressive Transformer architectures, and the applicability of the coverage principle to other architectures (e.g., state-space models) remains to be validated.
Appendix B Detailed experimental setup
B.1 Dataset construction details
We now provide detailed information about our dataset construction process. While we primarily explain this process for the 2-Hop task, we follow similar procedures for the other compositional structures.
Vocabulary and Token Representation
For a task with token set size $|\mathcal{X}|$, we create $|\mathcal{X}|$ special tokens of the form <t_0>, <t_1>, $\ldots$, <t_$(|\mathcal{X}|-1)$>, which we append to the standard GPT-2 vocabulary. We also add special tokens </a> to mark the end of sequences. For Chain-of-Thought (CoT) experiments, intermediate computations are represented in the target sequence as the actual intermediate token.
Function Construction
For the 2-Hop task, we construct two primitive functions $f_{1}:\mathcal{X}^{2}\rightarrow\mathcal{X}$ and $f_{2}:\mathcal{X}^{2}\rightarrow\mathcal{X}$ by randomly mapping from their respective domains to the codomain $\mathcal{X}$. We create the domain by taking the Cartesian product of the token set with itself. For each function, we randomly designate a fraction $p_{\text{seen}}=0.7$ of its domain as the "seen" portion, resulting in sets $S_{f_{1}}$ and $S_{f_{2}}$.
Dataset Generation Algorithm
To generate the training dataset, we first identify all possible combinations where both primitive operations come from their respective "seen" domains. Specifically, we find all valid tuples $(x_{1},x_{2},x_{3},t)$ such that:
$(x_{1},x_{2})$ $\in\text{domain}(S_{f_{1}})$ (1)
$(f_{1}(x_{1},x_{2}),x_{3})$ $\in\text{domain}(S_{f_{2}})$ (2)
$t$ $=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$ (3)
From this set of all possible in-domain combinations, we uniformly sample $N$ examples to form our training dataset. When the number of possible combinations exceeds $N$, this sampling ensures the model sees only a subset of possible in-domain combinations.
Test Set Construction
We carefully construct test sets to evaluate the model’s generalization capabilities across different coverage conditions. Our test sets contain:
  • In-Domain (ID) Test Set: Combinations not seen during training but where both primitive operations were observed in other contexts. These examples may lie within the coverage as defined by our framework.
  • Out-of-Domain (OOD) Test Set: Examples where at least one primitive operation was never observed in training. These fall outside the coverage boundary.
Input-Output Format
The dataset is formatted for auto-regressive token prediction. For the standard 2-Hop task, inputs comprise three tokens representing $x_{1}$, $x_{2}$, and $x_{3}$, while the target includes these input tokens followed by the prediction $t$ and an end marker. Below are the examples of the dataset format for different settings.
  • Standard Format:
    • Input: <t_5><t_12><t_3>
    • Target Completion: <t_17></a>
    • The model must predict the final output token followed by the end marker.
  • Chain-of-Thought Format:
    • Input: <t_5><t_12><t_3>
    • Target Completion: <t_9><t_17></a>
    • The model must first predict the intermediate computation result <t_9> (where <t_9> = $f_{1}$(<t_5>, <t_12>)), followed by the final output.
  • Partial Computation Format ($f_{1}$):
    • Input: <t_5><t_12>
    • Target Completion: <t_9></a>
    • These examples represent the primitive function applications used to construct the full compositional task.
For the other compositional tasks, we follow analogous construction procedures, adjusting the number of input tokens and the composition structure based on the specific task’s requirements. For example, Parallel 2-Hop requires four input tokens, while 3-Hop follows a three-step composition chain requiring appropriate modifications to the function construction and sampling procedures.
B.2 Training details
Table 1: Model configurations for different GPT-2 variants used in our experiments
Configuration GPT-2-Small GPT-2 GPT-2-XL
Number of Attention Heads 6 12 25
Number of Layers 4 8 48
Hidden Dimension 768 768 1600
Total Parameters 68M 96M 1.5B
For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters). As shown in Tab. 1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768. The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768. Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600. The implementation follows the codebase from [19].
We train all models using the AdamW optimizer with beta values of (0.9, 0.999) and epsilon of 1e-8. We set the learning rate to 8e-4 with a weight decay of 0.1. A batch size of 16,384 is used, with full gradient descent applied for datasets smaller than the batch size. All training is conducted with mixed precision (fp16) on 4 NVIDIA A100 GPUs with 80GB memory each. We employ a constant learning rate schedule with a linear warmup period of 2,000 steps. This standardized training configuration is maintained across all experiments to ensure fair comparisons between different task structures and dataset sizes, unless explicitly varied in specific ablation studies.
Appendix C Implementation details for the coverage determination algorithm
Input: Training examples $\mathcal{D}=\{(x_{i},f(x_{i}))\}$, where $x_{i}\in\mathcal{X}^{n}$ and $f(x_{i})\in\mathcal{X}$
Minimum evidence threshold $k\geq 1$
Output: Coverage set $\text{Cover}(\mathcal{D})$
/* STEP 1: Build behavior maps for each subsequence pattern */
foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
       $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in\mathcal{D}\}$
end foreach
/* STEP 2: Identify functionally equivalent subsequences */
foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
       $UF_{I}\leftarrow\text{new UnionFind}()$
      foreach pair of subsequences $(\alpha,\beta)$ in $\text{Behavior}_{I}$ do
[…]
            if No contradictions in SharedComplements and matching evidence $\geq k$ then
                   $UF_{I}.\text{Union}(\alpha,\beta)$ ;
                    // Mark as functionally equivalent
             end if
       end foreach
      $\text{EquivClasses}_{I}\leftarrow UF_{I}$
end foreach
/* STEP 3: Build substitution graph */
$G\leftarrow\text{empty graph}$ with nodes for all $x\in\mathcal{X}^{n}$
foreach pair of inputs $(x,y)$ with $f(x)=f(y)$ do
       foreach subset $I$ where $x$ and $y$ differ only on indices in $I$ do
             if $\text{EquivClasses}_{I}.\text{Find}(x_{I})=\text{EquivClasses}_{I}.\text{Find}(y_{I})$ then
                   Add edge $(x,y)$ to $G$
                  break
             end if
       end foreach
end foreach
/* STEP 4: Determine coverage */
$\text{Cover}(\mathcal{D})\leftarrow\bigcup_{x\in\mathcal{D}}$ ConnectedComponent$(G,x)$
return $\text{Cover}(\mathcal{D})$
Algorithm 1 $k$-Coverage Determination Algorithm
Algorithm 1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
Stage 1: Behavior mapping
We first analyze the training data to create a mapping of behaviors for each possible subsequence of the input. For each subset of indices $I$, we record how different subsequences $x_{I}$ behave when paired with their complements $x_{[n]\setminus I}$, essentially mapping each subsequence to a function from complements to outputs.
Stage 2: Equivalence class construction
For each subset of indices $I$, we build equivalence classes of subsequences that exhibit functionally identical behavior. Two subsequences are considered equivalent only if: (1) they share at least $k$ distinct complements where they produce the same output, and (2) they never produce different outputs when given the same complement (no contradictions). We use a Union-Find data structure to efficiently track and merge these equivalence classes. The Union-Find (or Disjoint-Set) data structure efficiently maintains a collection of disjoint sets, supporting two key operations: (1) Find - determine which set an element belongs to, and (2) Union - merge two sets.
Stage 3: Substitution Graph Construction
We construct a graph where nodes represent input sequences from our training and test sets, rather than the entire domain space (which would be computationally prohibitive for large token sets). We add an edge between two inputs $x$ and $y$ if and only if: (1) they produce the same output, (2) they differ only in one subsequence position set $I$, and (3) their differing subsequences belong to the same equivalence class. This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output. Our implementation uses parallel processing to efficiently construct this graph even for large datasets.
Stage 4: Coverage computation
Finally, we compute the coverage set by taking the union of all connected components in the substitution graph that contain at least one training example. This set comprises all inputs that are reachable from the training data through chains of equivalent subsequence substitutions.
Appendix D Detailed analysis for representation unification experiments
D.1 Causal Tracing Methodology
To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference [67, 66]. Specifically, we measure the causal effect using the indirect effect metric defined in [95]. This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization. We illustrate the measurement with 2-Hop task.
Our analysis begins by collecting three types of computational traces:
  1. 1.
    Clean run ($G$): We run the model on a compositional task with input $(x_{1},x_{2},x_{3})$ where the corresponding output is $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$.
  2. 2.
    Corrupted run ($G^{*}$): We replace the original input with a corrupted version by changing the first two tokens $(x_{1},x_{2})$ to $(x_{1}^{\prime},x_{2}^{\prime})$, where $f_{1}(x_{1}^{\prime},x_{2}^{\prime})\neq f_{1}(x_{1},x_{2})$. This ensures that the model produces a different final output $t^{*}\neq t$. During this run, we cache all hidden states ${h^{*}}_{i}^{(\ell)}$ for each token position $i$ and layer $\ell$.
  3. 3.
    Patched run ($G[\leftarrow{h^{*}}_{i}^{(\ell)}]$): We run the model on the input from the clean run, but at a specific token position $i$ and layer $\ell$, we replace the hidden state with the corresponding state from the corrupted run.
To quantify the causal effect of a specific hidden state $h_{i}^{(\ell)}$ on the model’s prediction, we measure the Indirect Effect (IE):
$\text{IE}_{h_{i}^{(\ell)}}=\frac{p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})-p(t^{*})}{p^{*}(t^{*})-p(t^{*})}$ (4)
where:
  • $p(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the clean run $G$
  • $p^{*}(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the corrupted run $G^{*}$
  • $p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the patched run $G[\leftarrow{h^{*}}_{i}^{(\ell)}]$
This metric quantifies how much corruption in a particular state affects the overall outcome. An IE value close to 1 indicates that the corruption of the state $h_{i}^{(\ell)}$ to ${h^{*}}_{i}^{(\ell)}$ alone almost completely changes the prediction to that of the corrupted run, suggesting that this state is causally important for the computation. Conversely, an IE value close to 0 indicates that the state has minimal causal impact on the prediction.
In our experiments, we apply causal tracing to analyze different subsets of test data categorized by their $k$-cutoff values, where $k$ represents the minimum evidence threshold required for functional equivalence (as defined in Sec. 3 of the main text). This allows us to correlate the strength of functional equivalence evidence with the formation of unified internal representations.
D.2 Causal tracing results for each $k$-cutoff value in 2-Hop task
Figure 8 displays the causal tracing results for the 2-Hop task, broken down by different $k$-cutoff values. We observe that the causal patterns are similar across different $k$-cutoff values, with slight differences in where and how strongly the causal effects manifest in the model. This suggests that once an example falls within coverage (even with minimal evidence, $k=1$), the model forms internal representations that play similar causal roles in prediction.
Figure 8: Causal tracing results for the 2-Hop task across different $k$-cutoff values, showing Indirect Effect (IE) scores at each layer and position.
D.3 Token set size ablation
We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$. For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99. Figure Fig. 9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases. The causal tracing results in Fig. 10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
Figure 9: IICG heatmap across different token set sizes, showing consistent representation clustering patterns.
Figure 10: Causal tracing results showing indirect effect heatmaps for different token set sizes $|{\mathcal{X}}|$.
D.4 Task ablation
We show that GPT-2 models trained on Parallel-2-Hop and 3-Hop tasks exhibit the same patterns: clustered functional equivalence representations of intermediate states at specific layers and positions, confirmed through cosine similarity analysis, with causal tracing analysis verifying their role in model predictions. For both tasks, we analyze with $|{\mathcal{X}}|=50$ and examine model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
Figures 11 and 12 show the results for the Parallel-2-Hop task. The IICG patterns reveal strong representation clustering at mid-layers: at positions $x_{2}$ and $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{4}$ when grouped by $b_{2}=f_{2}(x_{3},x_{4})$. Causal tracing confirms the causal role of these clustered representations in the model’s predictions.
Figure 11: IICG heatmap for Parallel-2-Hop task with grouping strategies based on $b_{1}=f_{1}(x_{1},x_{2})$ (Left), $b_{2}=f_{2}(x_{3},x_{4})$ (Middle), and $t=f_{3}(b_{1},b_{2})$ (Right).
Figure 12: Causal tracing results showing indirect effect heatmap for Parallel-2-Hop task. Left: perturbation with different $(x_{1},x_{2})$ pair leading to a different $b_{1}$ value. Right: perturbation with different $(x_{3},x_{4})$ pair leading to a different $b_{2}$ value.
Similarly, Figures 13 and 14 show results for the 3-Hop task. The IICG patterns exhibit strong representation clustering at mid-layers: at position $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{3}$ when grouped by $b_{2}=f_{2}(b_{1},x_{3})$. Causal tracing again confirms the causal importance of these representations.
Figure 13: IICG heatmap for 3-Hop task with grouping strategies based on $b_{1}=f_{1}(x_{1},x_{2})$ (Left), $b_{2}=f_{2}(b_{1},x_{3})$ (Middle), and $t=f_{3}(b_{2},x_{4})$ (Right).
Figure 14: Causal tracing results showing indirect effect heatmap for 3-Hop task. Left: perturbation with different $(x_{1},x_{2})$ pair leading to different $b_{1}$ value. Right: perturbation leading to different $b_{2}=f_{2}(b_{1},x_{3})$ value.
These results demonstrate that the formation of clustered intermediate state representations and their causal role in compositional generalization is a consistent phenomenon across different task structures, supporting the generality of our findings beyond the 2-Hop task.
Appendix E Derivation of Result 5.1 (power–law lower bound)
Problem setting
Let ${\mathcal{X}}$ be a finite token set with cardinality $|{\mathcal{X}}|$. The target mapping
$f:{\mathcal{X}}^{3}\longrightarrow{\mathcal{X}},\qquad f(x_{1},x_{2},x_{3})=f_{2}\bigl{(}f_{1}(x_{1},x_{2}),\,x_{3}\bigr{)}$
is a two-hop composition of unknown primitives $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. Write $b=f_{1}(x_{1},x_{2})\in{\mathcal{X}}$ for the intermediate state. Throughout the proof we impose two assumptions.
  1. (A1)
    Balanced classes: Each intermediate value $b\in{\mathcal{X}}$ is realized by exactly $|{\mathcal{X}}|$ first-hop pairs, i.e., the sets $E_{b}:=\{(x_{1},x_{2})\in{\mathcal{X}}^{2}:f_{1}(x_{1},x_{2})=b\}$ all have size $|{\mathcal{X}}|$ and form a partition of ${\mathcal{X}}^{2}$.
  2. (A2)
    Uniform training sampler: The training set $D$ contains $N$ triples drawn uniformly with replacement from the domain ${\mathcal{X}}^{3}$.
Functional $k$-equivalence and learner model
For two first-hop fragments $a,a^{\prime}\in{\mathcal{X}}^{2}$ define
$a\sim a^{\prime}\quad:\Longleftrightarrow\quad f_{1}(a)=f_{1}(a^{\prime}).$
They are functionally $k$-equivalent w.r.t. $D$ (denoted $a\equiv^{k}_{D}a^{\prime}$) when there exist $k$ distinct contexts $c_{1},\dots,c_{k}\in{\mathcal{X}}$ such that for every $r\leq k$ both $(a,c_{r})$ and $(a^{\prime},c_{r})$ appear in $D$ and $f(a,c_{r})=f(a^{\prime},c_{r})$. We call each $(a,c_{r}),(a^{\prime},c_{r})$ an evidence pair.
The coverage principle by itself is only a necessity statement: outside the $k$-coverage region a purely pattern matching learner’s predictions are unconstrained. To convert this into a sufficient data condition we adopt an explicit inductive bias, matching the premise of Result 5.1:
Learner assumption: Whenever two fragments become linked by $k$ independent (i.e., pairwise from distinct contexts) evidence pairs, the learner treats them as functionally equivalent; the learner also propagates equivalence transitively along chains of such links.
With this rule in place the relevant structure inside each class $E_{b}$ is the $k$-evidence graph: vertices are the $|{\mathcal{X}}|$ first-hop pairs in the class and an edge connects two vertices whenever the pair is observed at least $k$ times in shared contexts. If that graph is connected (every vertex reachable from every other), then every fragment in $E_{b}$ is linked by a chain of $k$-evidence steps and is therefore recognized as equivalent by the learner. Hence
Data sufficiency criterion: If the $k$-evidence graph of each class $E_{b}$ is connected, the learner generalizes perfectly to all in-domain (ID) inputs.
This criterion requires connectivity, and we aim to derive the condition to yield the connectivity with high probability using Erdős–Rényi model [96, 97]. The minimal dataset size achieving this with high probability is denoted $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$.
Note that the $k$-evidence graph on $E_{b}$ is the restriction of the substitution graph $\mathcal{G}_{D,k}$ (Def. 3.2) to the vertex set $E_{b}\times\{x_{3}\}$ for any fixed $x_{3}$. Connectivity of every class therefore implies that every first-hop fragment lies in the same connected component as some training input, i.e., the entire in-domain set is contained in $k$-coverage. Under the learner assumption, this is both necessary and sufficient for perfect ID generalization, yielding Res. 5.1.
E.1 Step 1: probability of a single evidence pair
Evidence pair probability
Fix two distinct first–hop fragments $i,j\in E_{b}$ and a context $c\in{\mathcal{X}}$. We want to find $p_{1}$, the probability that context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$:
$p_{1}:=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
Let $q:=1/|{\mathcal{X}}|^{3}$ denote the probability of drawing any specific triple in a single draw. Using the inclusion–exclusion principle:
$p_{1}$ $=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
$=1-\Pr[(i,c)\not\in D]-\Pr[(j,c)\not\in D]+\Pr[(i,c)\not\in D\text{ and }(j,c)\not\in D]$
$=1-(1-q)^{N}-(1-q)^{N}+(1-2q)^{N}$ (S1.1)
For $q\ll 1$ (which holds when $|{\mathcal{X}}|$ is large), we can use Taylor expansion:
$(1-q)^{N}$ $=1-Nq+\frac{N(N-1)}{2}q^{2}+O(q^{3})$
$(1-2q)^{N}$ $=1-2Nq+\frac{N(N-1)}{2}(2q)^{2}+O(q^{3})$ (S1.2)
Substituting these approximations:
$p_{1}$ $=1-2\left(1-Nq+\frac{N(N-1)}{2}q^{2}\right)+\left(1-2Nq+2N(N-1)q^{2}\right)+O(q^{3})$
$=1-2+2Nq-N(N-1)q^{2}+1-2Nq+2N(N-1)q^{2}+O(q^{3})$
$=N(N-1)q^{2}+O(q^{3})$
$=\frac{N(N-1)}{|{\mathcal{X}}|^{6}}+O\left(\frac{1}{|{\mathcal{X}}|^{9}}\right)$
$=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))$ (S1.3)
Therefore, in the regime of interest ($N\gg|{\mathcal{X}}|$ but $N\ll|{\mathcal{X}}|^{3}$):
$\boxed{p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))}$
Equation (E.1) gives an exact expression for the probability (up to lower-order terms) that the single context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$.
Remark on “lucky coincidences”
Because functional $k$-equivalence demands consistency across all $k$ evidences, a single coincidental equality $f(i,c)=f(j,c)$ with $i\not\sim j$ can only masquerade as evidence when $k=1$ and the dataset lacks any contradicting context $c^{\prime}$. For $k\geq 2$ the joint probability that two independent contexts simultaneously produce such coincidences is $|{\mathcal{X}}|^{-k}$ per fragment pair and hence negligible relative to the isolate probability once $N=\Omega(|{\mathcal{X}}|^{2})$. Consequently, the effects of such “lucky coincidences” on the lower bound in Res. 5.1 can be neglected.
E.2 Step 2: probability of observing $k$ evidences for one fixed pair
Having established the probability $p_{1}$ for a single evidence pair in Step 1, we now derive $p_{k}$, the probability that a dataset provides at least $k$ distinct contexts as evidence for the functional equivalence of two fragments $i$ and $j$. This probability will determine the edge probability in the random graph model analyzed in Step 3.
Indicators for one fragment pair
Fix two distinct first-hop fragments $i,j\in E_{b}$ and, for each context (third token) $c\in{\mathcal{X}}$, set
[…]
Thus $Z_{ij}(c)=1$ exactly when the single context $c$ supplies an evidence pair for the functional equivalence of $i$ and $j$.
Single-context success probability
From (E.1),
$p_{1}:=\Pr[Z_{ij}(c)=1]=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)},\qquad(|{\mathcal{X}}|\to\infty).$
Negatively correlated counts and an i.i.d. surrogate
Because all $N$ draws come from a single multinomial $\bigl{(}N,1/|{\mathcal{X}}|^{3}\bigr{)}$, the indicators $\{Z_{ij}(c)\}_{c\in{\mathcal{X}}}$ are negatively correlated: drawing many triples with one context leaves fewer draws for the others. Negative correlation decreases the probability that several $Z_{ij}(c)$ equal 1 simultaneously. Hence the tail probability for the true count
$Y_{ij}:=\sum_{c\in{\mathcal{X}}}Z_{ij}(c)$
is upper-bounded by the tail of an i.i.d. binomial variable with the same single-trial success probability $p_{1}$. Concretely, define
$Y_{ij}^{\star}\sim\mathrm{Binom}\bigl{(}|{\mathcal{X}}|,p_{1}\bigr{)},\qquad p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)}.$
Then for every real $t$
$\Pr\bigl{[}Y_{ij}\geq t\bigr{]}\leq\Pr\bigl{[}Y_{ij}^{\star}\geq t\bigr{]}.$
Using $Y_{ij}^{\star}$ therefore overestimates the chance of obtaining $k$ or more distinct evidence contexts, which is conservative for our goal of deriving a lower bound on the required dataset size $N$.
Since $Y_{ij}^{\star}$ is binomial with mean
$\mu:=\mathbb{E}[Y_{ij}^{\star}]=|{\mathcal{X}}|\,p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{5}}\bigl{(}1+o(1)\bigr{)},$
we may work henceforth with $Y_{ij}^{\star}$ alone; the resulting bounds apply verbatim to the original $Y_{ij}$.
Poisson tail via Le Cam
[…]
$\Pr[Y_{ij}\geq k]\leq\Pr[Y_{ij}^{\star}\geq k]=\sum_{r=k}^{\infty}\frac{e^{-\mu}\mu^{r}}{r!}\leq\frac{\mu^{k}}{k!}\bigl{(}1+o(1)\bigr{)}.$
This upper bound $p_{k}:=\mu^{k}/k!(1+o(1))$ will be the edge probability used in the connectivity threshold of Step 3.
E.3 Step 3: connectivity inside each equivalence class
With the edge probability $p_{k}$ from Step 2, we now analyze when the $k$-evidence graphs become connected. Recall that under our learner assumption, perfect generalization requires every equivalence class to form a connected component in the $k$-evidence graph. We model this as a random graph connectivity problem.
Random graph construction
Fix one balanced class $E_{b}$ of size $n:=|{\mathcal{X}}|$ (Assumption (A1)). Create a graph $G_{b}$ whose vertices are the first-hop pairs in $E_{b}$, and place an undirected edge $\{i,j\}$ exactly when the pair $(i,j)$ has been observed in at least $k$ distinct shared contexts. By Step E.2, each potential edge appears with probability
$p_{k}=\Pr\bigl{[}Y_{ij}\geq k\bigr{]}=\frac{\mu^{k}}{k!}\,,\qquad\mu=\frac{N^{2}}{|{\mathcal{X}}|^{5}}.$
Why $G_{b}$ can be approximated as Erdős–Rényi
As mentioned earlier, the dependence among distinct edges in $G_{b}$ arises from the constraint that the total sample size is $N$, which induces a negative correlation. Such negative dependence reduces the likelihood of simultaneously creating many edges. Consequently, viewing $G_{b}$ as an independent Erdős–Rényi graph $G(n,p_{k})$ provides a conservative model, and any threshold we derive for connectivity under independence remains valid (or becomes easier to satisfy).
Classwise connectivity threshold
For Erdős–Rényi graphs, the classical result of Erdős–Rényi [97] states
$\Pr\bigl{[}G(n,p)\text{ is connected}\bigr{]}\xrightarrow[n\to\infty]{}1\quad\Longleftrightarrow\quad p\geq\frac{\log n+\omega(1)}{n}.$
Setting $n=|{\mathcal{X}}|$ and $p=p_{k}$ yields the requirement
$p_{k}=\frac{\mu^{k}}{k!}\geq\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|}(1+o(1)).$
Substituting $\mu=N^{2}/|{\mathcal{X}}|^{5}$ and rearranging gives
$\frac{N^{2k}}{|{\mathcal{X}}|^{5k}\,k!}\gtrsim\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|},\qquad\Longrightarrow\qquad N\gtrsim|{\mathcal{X}}|^{\frac{5k-1}{2k}}\,\bigl{(}k!\log|{\mathcal{X}}|\bigr{)}^{1/(2k)}.$
Because $(k!\log|{\mathcal{X}}|)^{1/(2k)}=(\log|{\mathcal{X}}|)^{O(1/k)}$ grows only poly-logarithmically, we hide it inside a $\tilde{\Omega}(\,\cdot\,)$:
[…]
Since every class must be connected to achieve the data–sufficiency criterion in the problem setting, we conclude
[…]
Appendix F Additional results for power-law scaling analysis
F.1 Measurement protocol for $N_{\mathrm{req}}$
To empirically determine the minimum dataset size required for reliable compositional generalization ($N_{\mathrm{req}}$), we develop a measurement protocol that accounts for practical computational constraints while ensuring robustness. For each token set size $|\mathcal{X}|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
Specifically, our criterion for “reliable generalization” on ID is defined as:
  • The model must reach ID test accuracy of 0.99 within 100 epochs after achieving training accuracy > 0.99.
This protocol balances several considerations:
  1. 1.
    Training-to-generalization delay: Larger datasets naturally require more iterations to fit training data. By measuring epochs after reaching training accuracy > 0.99, we focus on the generalization gap rather than conflating it with initial training difficulty.
  2. 2.
    Epoch-based measurement: Using epochs rather than raw training steps ensures that the model sees each functional equivalence evidence approximately the same number of times, regardless of dataset size. This provides a fairer comparison across different dataset sizes.
  3. 3.
    Practical time constraints: While indefinite training might eventually yield generalization with smaller datasets, we established a reasonable upper bound (100 epochs post-training convergence) to reflect practical limitations.
  4. 4.
    Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion. This establishes that our measurement error is at most $-\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
F.2 Measured power-law scaling constants across task structures and model sizes
Using our measurement protocol, we measure the required dataset size $N_{\mathrm{req}}$ across three different compositional structures (2-Hop, Parallel-2-Hop, and 3-Hop) and three model scales (68M, 96M, and 1.5B parameters). For each task structure, we vary the token set size $|\mathcal{X}|$ from 50 to 200, allowing us to observe the scaling relationship.
Table 2 presents the power-law exponents obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$. The consistency of exponents across model sizes suggests that the observed power-law scaling relates to properties of the compositional tasks themselves, rather than model capacity. This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
Table 2: Power law exponents for different tasks and GPT-2 sizes, obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots. $R^{2}>0.99$ for all linear fitting.
Model Size 2-hop Parallel-2-hop 3-hop
68M 2.13 2.47 2.61
96M 2.26 2.35 2.50
1.5B 2.28 2.17 2.60
F.3 Robustness to hyperparameter variations
Figure 15: Robustness of power-law scaling relationship to hyperparameter variations in the 2-Hop task with $|\mathcal{X}|=50$. Each line shows the training and test accuracy curves for a different configuration: (1) baseline, (2) reduced learning rate (4e-4, half of baseline), (3) reduced weight decay (0.01, one-tenth of baseline), and (4) changed generalization criteria (test accuracy > 0.95 within 10 epochs after training accuracy > 0.95). $R^{2}>0.99$ for all linear fitting.
To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations. Figure 15 demonstrates that for the 2-Hop task with $|\mathcal{X}|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
  1. 1.
    Learning rate reduction: Halving the learning rate from 8e-4 to 4e-4
  2. 2.
    Weight decay reduction: Decreasing weight decay by a factor of 10 (from 0.1 to 0.01)
  3. 3.
    Generalization criteria modification: Requiring test accuracy > 0.95 within 10 epochs after training accuracy > 0.95
This robustness to hyperparameter variations suggests that the power-law relationship between token set size and required dataset size is primarily a property of the compositional generalization process, rather than an artifact of specific optimization settings.
Appendix G Detailed analysis for Non-tree task
This section provides additional analyses that support our findings in Sec. 5.4 regarding the challenges of path ambiguity in the Non-tree task.
G.1 Coverage analysis
Figure 16: Coverage analysis for Non-tree task with $|\mathcal{X}|=50$. The graph shows the percentage of ID test data covered at different $k$ values across various dataset sizes ($N$). Compared to the 2-Hop task (Fig. 3, left), Non-tree has significantly lower coverage at equivalent dataset sizes, indicating that path ambiguity impedes the formation of functional equivalence relationships.
Fig. 16 demonstrates that with equivalent training dataset sizes, a smaller percentage of ID test examples fall inside $k$-coverage for the Non-tree task compared to the 2-Hop task shown in Fig. 3 (Left). This aligns with our theoretical analysis in Sec. 5.4, which predicts that path ambiguity limits the establishment of functional equivalence relationships between input subsequences, as the model cannot generalize across different $x_{2}$ values in the Non-tree structure even when they produce the same intermediate state $b=f_{1}(x_{1},x_{2})$.
G.2 Effect of model scaling
Figure 17: ID test accuracy comparison between GPT-2 (96M parameters) and GPT-2-XL (1.5B parameters) on the Non-tree task with $|\mathcal{X}|=50$, measured 100 epochs after training accuracy exceeds 0.99. Despite the 15x increase in parameter count, the accuracy does not increase.
Fig. 17 shows that scaling up the model size to GPT-2-XL (1.5B parameters) does not significantly improve generalization performance on the Non-tree task, even when measured 100 epochs after reaching training accuracy > 0.99. This suggests that the challenges posed by path ambiguity cannot be overcome simply by increasing model capacity, supporting our claim that the limitation is structural rather than related to model capacity.
G.3 Representation analysis in successful generalization
For a model that eventually achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|\mathcal{X}|=50$, $N=50\text{k}$), we conduct causal tracing analysis to understand how it achieves generalization despite path ambiguity.
Figure 18: Causal tracing analysis for the Non-tree model after extended training. The heatmap shows indirect effect values across different layer-token positions. Left: perturbation leading to different intermediate state $b=f_{1}(x_{1},x_{2})$. Middle: same $b$ but different $x_{2}$. Right: different $b$ and $x_{2}$.
The causal tracing results in Fig. 18 reveal how the model achieves generalization in the presence of path ambiguity. Across all perturbation strategies, the model’s predictions show strong causal dependence on representations at both the $x_{1}$ and $x_{2}$ positions, indicating reliance on direct access to both input tokens rather than an abstracted intermediate computation. This pattern contrasts sharply with the 2-Hop task, where causal effects concentrate primarily at positions corresponding to clustered functional equivalence representations.
This analysis demonstrates that even models achieving high accuracy on Non-tree tasks do so by developing context-dependent representations rather than unified abstractions of intermediate states. The model forms separate computational pathways conditioned on the $x_{2}$ value, rather than learning a single unified representation of the intermediate state $b=f_{1}(x_{1},x_{2})$. This represents a fundamentally different solution strategy compared to the 2-Hop task, with implications for both generalization capability and interpretability.
Appendix H Detailed discussion on the taxonomy of generalization
In Section 7, we introduce a taxonomy of generalization that extends beyond the coverage principle. We propose this taxonomy as an initial conceptual framework to disentangle different mechanisms of generalization that are often conflated in the literature. This categorization offers hypotheses for why neural networks succeed or fail at compositional tasks, connecting empirical observations to theoretical principles. We emphasize that the taxonomy is preliminary and that Sec. H.5 outlines its current limits, open questions, and future directions.
H.1 Type-I: Structure-based generalization
Type-I generalization occurs when a model succeeds by identifying and utilizing functionally equivalent components based on how primitive functions are composed. This form of generalization is precisely what we have formalized and studied via the coverage principle. Crucially, this kind of generalization remains strictly bounded by coverage as defined in Section 3.
In Type-I generalization, a model learns to recognize when different input fragments yield identical results in shared contexts, allowing it to generalize to new combinations of these fragments. However, without exposure to sufficient functional equivalence evidence during training, the model cannot reliably generalize to novel compositions.
Type-I generalization effectively describes what a pattern matching learner can achieve through the functional equivalence observed by input-output relationships, without requiring abstract variable binding mechanisms. This limitation aligns with the classic argument by Fodor and Pylyshyn [9] and Marcus [37] that connectionist networks lacking explicit variable binding capabilities [^13] We expand on the variable binding concept defined in Sec. 2 beyond flexibly associating variables with values, and we distinguish between self-replicative expansion (e.g., 2-hop $\to$ 3-hop via Type-III) and complex compositional binding (e.g., substituting noun phrases for nouns), where the latter may require explicit variable binding mechanisms beyond parameter sharing. cannot exhibit fully systematic generalization. We introduce this perspective as a conceptual framing of the coverage principle: it describes the ceiling of pattern-based generalization in neural networks that rely on pattern matching rather than symbolic variable binding.
H.2 Type-II: Function property-based generalization
Type-II generalization occurs when a model exploits specific properties intrinsic to individual primitive functions, enabling generalization beyond patterns observed during training. Unlike Type-I generalization which requires explicit evidence of functional equivalence, Type-II generalization leverages invariant properties that hold for all inputs of the primitive, making the function’s output fully determined by those invariances even in unobserved contexts.
We identify several sub-families of property-based generalization:
  • Algebraic invariances: commutativity, associativity, and similar properties
  • Group-theoretic structure: as in Fourier basis representations for modular arithmetic [63]
  • Input irrelevance: attention sparsity that ignores distractors [81]
The reversal curse [16] exemplifies the layered nature of compositional challenges across multiple generalization types. The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. Property-based approaches like bidirectional training [83] partially succeed by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data. However, recent work [42] reveals deeper representational binding constraints: even though models can exploit inverse mapping properties in idealized conditions (abstract concept-level tasks), they struggle to maintain consistent concept representations when moving to realistic surface-level predictions, where entities switch roles between subject and object positions. This demonstrates how real compositional tasks involve multiple mechanism layers, where achieving one capability may be insufficient due to representational binding limitations.
Type-II generalization fundamentally differs from Type-I in that it can, in principle, transcend the limitations imposed by the coverage principle, allowing models to generalize correctly to inputs that lie outside the coverage of their training data by leveraging invariant properties of the underlying functions. Systematic characterization of pure Type-II extrapolation remains limited and is an important direction for future work.
H.3 Type-III: Shared-operator generalization
Type-III generalization emerges through reuse of identical primitive functions across different computational positions, e.g., when $f_{1}=f_{2}$ in a multi-hop reasoning task. Unlike Type-I (bounded by coverage) and Type-II (exploiting algebraic properties), Type-III generalization leverages the structural repetition of identical operations.
This mechanism is not unique to Transformers. Recurrent architectures [84] exemplify Type-III generalization through their recurrent weight sharing: the same transformation is applied at each time step, enabling the network to process sequences of varying lengths despite training on fixed-length examples. This architectural inductive bias directly encodes the assumption that the same operation should apply across temporal positions, allowing recurrent architectures to extrapolate to longer sequences [85].
In the Transformer context, Wang et al. [19] demonstrate that Universal Transformer layers [86] enable generalization beyond in-domain data in two-hop tasks, while [87] show improved compositional reasoning with weight-shared layers. These findings suggest parameter sharing as a key enabling mechanism, where a single learned implementation of a function can be applied across multiple contexts.
A prominent manifestation of Type-III generalization appears in depth extrapolation and length generalization tasks, where models generalize to deeper or longer compositional structures than seen during training [99, 100, 101, 102]. While length generalization tasks fundamentally involve applying the same computation to longer inputs (Type-III), successful generalization often requires additional mechanisms like specialized positional encodings that facilitate computation reuse across positions. The observed brittleness of current approaches [103] suggests that such inductive biases are necessary but may not be sufficient for robust systematicity.
In summary, Type-III generalization extends beyond the coverage-bound limitations of Type-I by exploiting identical operations, without requiring the algebraic property understanding characteristic of Type-II. Type-III therefore complements Types I and II, but may still operate without explicit variable binding.
H.4 Relationship to prior taxonomies of generalization
Prior taxonomies of generalization in neural networks have primarily categorized the phenomena of compositional generalization rather than the underlying mechanisms. Lake and Baroni [12] distinguish between “mix-and-match” generalization (for small training-test differences) and systematic compositional generalization (requiring abstract rule application). Other frameworks focus on empirical evaluations through benchmarks [23] or decomposing compositional abilities into constituent skills [38].
Our taxonomy differs by categorizing the mechanisms that enable generalization rather than categorizing based on the phenomena. By distinguishing between structure-based pattern matching (Type-I), algebraic property exploitation (Type-II), and shared-operator transfer (Type-III), our framework provides a lens for interpreting why models succeed or fail at certain compositional tasks. This mechanistic approach not only explains observed behaviors but also offers hypotheses about when Types II or III become necessary for certain generalization challenges.
H.5 Open questions and future directions
To operationalize this taxonomy, future work could develop diagnostics: coverage ablation tests for Type-I (systematically removing training examples to test coverage boundaries), algebraic identity probes for Type-II (testing whether models recognize properties like commutativity without explicit training), and parameter-sharing ablations for Type-III (comparing shared vs. independent weights across positions). Current examples in this paper are illustrative rather than definitive classifications; real tasks often involve combinations of mechanisms, and our taxonomy provides a scaffold for disentangling them rather than enforcing mutually exclusive categories.
Several fundamental questions remain. First, practical computations often involve mixed ingredients of generalization, spanning multiple types. When classification is ambiguous, we suggest identifying the dominant mix of types that contribute, as most realistic tasks involve combinations rather than pure instances of any single type. Second, identifying the true compositional structure underlying a task can be ambiguous, as Locatello et al. [52] demonstrate regarding multiple valid decompositions. Third, real-world datasets contain heterogeneous computational structures, and it remains unclear how neural networks accommodate this diversity. Fourth, none of the three mechanisms solves the variable binding problem, and devising diagnostics and architectures for binding remains crucial.
Future work could explore connections to complementary frameworks:
  • Minimum Description Length [73]: Type-II may compress regularities that lower description length. Similarly, Type-III can be viewed as compression through parameter reuse.
  • Formal language learning [104, 105]: Connections to grammatical inference and automata learning
  • Cognitive science [106]: Human compositional learning strategies that may inspire architectural innovations
This taxonomy is not intended as a complete classification but rather as a conceptual framework to organize our understanding of generalization mechanisms and stimulate further research into the fundamental capabilities and limitations of neural networks on compositional tasks. Ultimately, developing architectural and optimization strategies that enable neural networks to leverage all three types more robustly and ultimately addressing the variable binding problem represents an important frontier for addressing the longstanding challenge of systematicity in connectionist systems.
Appendix I Partial computation observation drives the alignment of functional equivalence representation and vocabulary space
In this section, we investigate how exposure to partial computations affects the interpretability of intermediate state representations through vocabulary space alignment. We compare two training conditions on a modified 2-Hop task with $|\mathcal{X}|=50$ and $N=10\text{k}$, after 40k epochs of training:
  1. 1.
    Standard Training: $f_{1}\neq f_{2}$, model only sees complete two-hop examples $(x_{1},x_{2},x_{3})\mapsto t$.
  2. 2.
    With Partial Computation: $f_{1}=f_{2}$, model additionally sees all possible partial computations $(x_{1},x_{2})\mapsto b$ where $b=f_{1}(x_{1},x_{2})$ (2,500 partial examples, not counted toward the $N=10\text{k}$ two-hop training data).
To assess interpretability, we measure the Mean Reciprocal Rank (MRR) of intermediate state representations when projected to vocabulary space using the unembedding matrix. Low MRR indicates that the model’s internal representation of intermediate state $b$ aligns with the corresponding vocabulary token.
Fig. 19 shows a striking contrast between the two conditions. Under standard training, the MRR score remains very high throughout training, indicating that intermediate representations are not aligned with vocabulary space despite the model successfully learning the compositional task. However, when partial computations are included, the MRR score becomes very high, demonstrating clear vocabulary alignment.
Figure 19: MRR scores for intermediate state representations projected to vocabulary space. Left: Standard training ($f_{1}\neq f_{2}$, no partial computation) shows very high MRR regardless of position and layer. Right: Training with partial computation ($f_{1}=f_{2}$, with partial examples) shows MRR of 0 in layers 3 to 8 at position $x_{2}$, indicating strong vocabulary alignment.
This experiment suggests that logit lens interpretability is orthogonal to functional equivalence representation formation. A model can develop functionally correct intermediate representations that enable compositional generalization while remaining completely uninterpretable through standard vocabulary projection techniques. Interpretability via logit lens requires explicit vocabulary anchoring through exposure to partial computations that map intermediate states to vocabulary tokens.
This finding has important implications for mechanistic interpretability research: the absence of interpretable representations through logit lens does not indicate the absence of structured internal computation. Furthermore, it suggests that interpretability techniques may need to account for how training data shapes the alignment between internal representations and vocabulary space, rather than assuming such alignment emerges naturally from task performance.

### v2

@@
-The Coverage Principle: A Framework for Understanding Compositional Generalization
+Characterizing Pattern Matching and Its Limits on Compositional Task Structures
 Abstract
-Large language models excel at pattern matching, yet often fall short in systematic compositional generalization.
-We propose the coverage principle: a data-centric framework showing that models relying primarily on pattern matching for compositional tasks cannot reliably generalize beyond substituting fragments that yield identical results when used in the same contexts.
-We demonstrate that this framework has a strong predictive power for the generalization capabilities of Transformers.
-First, we derive and empirically confirm that the training data required for two-hop generalization grows at least quadratically with the token set size, and the training data efficiency does not improve with 20x parameter scaling.
-Second, for compositional tasks with path ambiguity where one variable affects the output through multiple computational paths, we show that Transformers learn context-dependent state representations that undermine both performance and interpretability.
-Third, Chain-of-Thought supervision improves training data efficiency for multi-hop tasks but still struggles with path ambiguity.
-Finally, we outline a mechanism-based taxonomy that distinguishes three ways neural networks can generalize: structure-based (bounded by coverage), property-based (leveraging algebraic invariances), and shared-operator (through function reuse).
-This conceptual lens contextualizes our results and highlights where new architectural ideas are needed to achieve systematic compositionally.
-Overall, the coverage principle provides a unified lens for understanding compositional reasoning, and underscores the need for fundamental architectural or training innovations to achieve truly systematic compositionality.
-[^1] All code and resources are available at: https://github.com/kaistAI/coverage-principle
+Despite impressive capabilities, LLMs’ successes often rely on pattern-matching behaviors, yet these are also linked to OOD generalization failures in compositional tasks.
+However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), obscuring a precise and testable account of how well LLMs perform generalization through pattern matching and their limitations.
+To address this ambiguity, we first formalize pattern matching as functional equivalence, i.e., identifying pairs of subsequences of inputs that consistently lead to identical results when the rest of the input is held constant.
+Then, we systematically study how decoder-only Transformer and Mamba behave in controlled tasks with compositional structures that isolate this mechanism.
+Our formalism yields predictive and quantitative insights: (1) Instance-wise success of pattern matching is well predicted by the number of contexts witnessing the relevant functional equivalence.
+(2) We prove a tight sample complexity bound of learning a two-hop structure by identifying the exponent of the data scaling law for perfect in-domain generalization.
+Our empirical results align with the theoretical prediction, under 20× parameter scaling and across architectures.
+(3) Path ambiguity is a structural barrier: when a variable influences the output via multiple paths, models fail to form unified intermediate state representations, impairing accuracy and interpretability.
+(4) Chain-of-Thought reduces data requirements yet does not resolve path ambiguity.
+Hence, we provide a predictive, falsifiable boundary for pattern matching and a foundational diagnostic for disentangling mixed generalization mechanisms.
 1 Introduction
+Despite the remarkable performance of Large Language Models (LLMs) (brown2020language; touvron2023llama), compositional generalization studies suggest that the core mechanism of generalization might be “pattern matching”, i.e., models learning local statistical regularities between input fragments and outputs in some cases (loula2018rearranging; johnson2017clevr; berglund2023reversal; wang2024grokked; mirzadeh2024gsm; keysers2019measuring; csordas2022ctl++).
+However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), discussing pattern matching without a precise definition and diagnosing it post-hoc rather than characterizing it predictively.
+As a result, it remains unclear which behaviors should count as pattern matching and which should not, obscuring a constructive and testable account of its boundary.
+To make this notion precise, we (1) introduce a model-agnostic, data-centric formalism for pattern matching and (2) systematically study how modern architectures like decoder-only Transformers (vaswani2017attention) and Mamba (gu2023mamba) perform generalization through pattern matching.
+Specifically, we first propose a model-agnostic and data-centric definition of pattern matching by formalizing the substitution of input patterns observed to result in identical outputs in shared contexts as functional-equivalence (Sec.˜3; henceforth, we use pattern matching as equivalent to functional-equivalence-based generalization).
+This induces a coverage boundary: if learning relies only on such evidence, reliable prediction is expected only for test inputs reachable by these substitutions.
+Moreover, to isolate and study pattern-matching behaviors, we use controlled setups that deliberately remove other generalization sources and make functional equivalence the primary available mechanism.
+With this setting, our formalism yields predictive and quantitative insights about the limitations of pattern matching that, to our knowledge, have not been well characterized in prior works:
+  • Generalization success is well predicted by the number of supporting contexts that witness the relevant functional equivalence.
+Mechanistically, Transformers implement functional equivalence via clustered intermediate representations at specific layers/positions, with clustering strength aligning with evidence strength (Sec.˜5).
+  • We prove a tight sample complexity bound of pattern matching on a two-hop structure by identifying the exponent of the data scaling law (in terms of the token set size) which is necessary and sufficient for perfect in-domain generalization (Theorem˜6.1).
+The measured power-law exponent agrees with our theoretical bounds and remains stable under roughly $20\times$ parameter increase (from 68M to 1.5B) for GPT-2 (radford2019language), and also holds for Mamba (gu2023mamba) architecture (Sec.˜6).
+  • When the same variable influences the output along multiple computational paths, models fail to form unified intermediate state representations.
+Analysis reveals that they instead develop context-dependent state representations, impairing both generalization and interpretability (Sec.˜7).
+  • Chain-of-Thought (CoT) supervision (wei2022chain) reduces data requirements yet does not resolve path ambiguity without seeing nearly exhaustive in-domain combinations (Sec.˜8).
+Finally, we situate this characterization of pattern matching within a mechanism-based taxonomy of generalization mechanisms, proposing two additional distinguishable mechanisms of generalization in compositional tasks: property-based and shared-operator generalization (Sec.˜9.2 and I).
+Our formalism opens several research directions with practical implications (e.g., targeted data augmentation to maximize coverage) and motivates expansion to broader tasks and architectures, as well as systematic studies of how pattern matching interacts with other generalization mechanisms.
+Overall, our study provides a predictive, falsifiable boundary for what can be achieved through pattern matching alone and a foundational diagnostic for disentangling mixed mechanisms in modern neural networks.
 […]
-Figure 1: Left: In a two-hop task $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, two fragments $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ satisfying $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ consistently yield the same final output when combined with the same context $x_{3}$, supporting their functional equivalence.
+Figure 1: Illustration of functional equivalence.
+Left: In a two-hop task $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, two fragments $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ satisfying $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ consistently yield the same final output when combined with the same context $x_{3}$, supporting their functional equivalence.
 Right: Among all possible inputs (few shown), we draw an edge between any two inputs that differ only by functionally equivalent fragments to form a substitution graph.
 Then, coverage is the set of observed inputs (highlighted as blue) and all inputs connected to them.
-The coverage principle states that a learner relying on pattern matching can reliably generalize only to inputs that are reachable from the observations by substituting functionally equivalent fragments.
-Human generalization capability is often attributed to our ability to manipulate symbols compositionally [1, 2].
-Despite the remarkable performance of Large Language Models (LLMs) [3, 4, 5, 6, 7, 8], our understanding of how they represent and manipulate compositional structures is still limited.
-Meanwhile, a growing body of work points to the brittleness of LLMs in performing systematic [^2] We refer to systematicity as an ability to generalize to unseen outcomes by exploiting their structural relations.
-For a more thorough discussion, see Fodor and Pylyshyn [9], Marcus [10], and Szabó [11].
-compositional reasoning [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
-These observations suggest that LLMs succeed primarily through pattern matching, i.e., by exploiting surface-level statistical correlations between input fragments and outputs, rather than systematic compositional reasoning.
-However, prior analyses of the pattern-matching behavior remain largely observational and task-specific, leaving us without a unified framework to predict when they will succeed or fail.
-We show that Transformers’ [26] generalization on synthetic compositional tasks is predictable through a unified framework grounded in one observation: Transformers generalize by utilizing functionally equivalent fragments supported by input-output relations.
-This insight allows us to move beyond observing success or failure cases to systematically predicting generalization boundaries.
-To this end, we first formalize the concept of functional equivalence, where two fragments (i.e., components of input sequences) are regarded as equivalent when they yield identical results in all shared contexts observed in the training data.
-Then, we introduce coverage, the set of inputs reachable from training data by substituting functionally equivalent fragments.
-Finally, we propose the coverage principle, which states that coverage is a necessary condition for compositional generalization, i.e., model predictions become unreliable beyond this fundamental boundary, as long as the only generalization mechanism is pattern matching.
-From this conceptual formulation, we derive specific predictions about generalization in Transformers.
-First, we derive and empirically confirm that reliable generalization on a two-hop reasoning task requires training data to scale at least quadratically with the token set size.
-Remarkably, our experiments suggest that this scaling relationship remains invariant even for 20x larger model, underscoring that the limitation is inherent to data properties rather than model capacity.
-Second, our framework also predicts specific complications in tasks with path ambiguities, where the same variable influences outputs through multiple computational paths.
-We show that Transformers cannot form unified representations of identical intermediate states in such cases, instead developing context-dependent state representations that fragment across different paths.
-This degrades both performance and interpretability, explaining why seemingly simple compositional tasks can remain challenging even for large models.
-Third, while Chain-of-Thought (CoT) supervision [27, 28] improves data efficiency by ‘flattening’ tasks into sequences of single-hop problems, our analysis reveals a fundamental limitation: CoT-trained models still cannot resolve path ambiguities without showing nearly all possible in-domain combinations.
-This finding provides insights into understanding the Transformers’ failure modes in complex reasoning tasks with path ambiguities (e.g., planning [29, 30, 31]).
-We place these findings within a broader, mechanism-based taxonomy that distinguishes three routes to generalization: (i) structure-based generalization via functional equivalence (the focus of this paper), (ii) property-based generalization that exploits algebraic invariances of individual primitive functions, and (iii) shared-operator generalization that reuses the same computation across positions.
-This taxonomy clarifies how certain model behaviours that occur outside the coverage region can arise from Type-II or Type-III mechanisms, providing a conceptual framework for analysing diverse sources of generalization.
-Overall, the coverage principle offers a unified conceptual framework providing valuable insights into compositional generalization, highlighting how compositional reasoning and its mechanistic interpretation remain inherently bounded by coverage limitations.
-This underscores the need for architectural or training innovations to achieve truly systematic generalization in neural networks.
+We define pattern matching as a type of generalization that occurs inside the coverage, harnessing functional equivalence.
 2 Related Work
-Compositional generalization and systematicity
-Since Fodor and Pylyshyn [9] posed the systematicity challenge, arguing connectionist systems inherently lack symbolic structural capabilities, understanding how neural networks might achieve compositional reasoning has remained a persistent research question [32, 33, 34, 35, 36, 37, 38, 39, 10].
-Central to this debate is the variable binding problem, how neural networks can flexibly associate variables with values across contexts [40, 41, 42], which Marcus [37] argues as the fundamental limitation of modern connectionist approaches.
-Despite AI advances, a growing body of work reveals limitations in neural networks’ compositional capabilities [14, 43, 18, 44, 45, 20, 46, 15, 21, 22, 23, 24, 47, 42].
-Dziri et al.
-[17] observe that Transformers perform linearized subgraph matching to solve compositional tasks, but leave open when and why it occurs.
-Our coverage principle addresses this gap by formally demarcating what pattern matching models can achieve without true variable binding, showing both the possibilities and fundamental limitations of pattern matching-based approaches through data-based preconditions.
-This aligns with empirical findings of Andreas [48] and Auersperger and Pecina [49], where compositional data augmentation that implicitly expands coverage significantly improves generalization.
-Our derived power-law scaling relationship between token set size and required dataset size contributes to scaling laws literature [50, 51], but differs by characterizing a data requirement invariant to model size for compositional tasks.
-Unlike prior disentanglement work [52, 53, 54, 55] which focuss on latent factors, our data-centric view identifies fundamental boundary conditions for pattern-based generalization.
-The coverage principle thus complements recent architectural or training innovations [56, 57, 58, 59, 60] by showing what any pattern matching approach must overcome to achieve true systematicity.
-Mechanistic interpretability
-Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors [61, 62, 63, 64].
-Recent work traces how Transformer components coordinate on compositional tasks [65, 66, 67, 68, 69].
-Wang et al.
-[19] demonstrate that in-domain compositional generalization emerges through grokking, but do not identify necessary data conditions for such emergence.
-Our data-centric framework complements this optimization-centric view by pinpointing when generalizing circuits can exist: only when training data provides sufficient evidence of functional equivalence, essentially formalizing when the pattern matching mechanisms that Fodor and Pylyshyn [9] critique can succeed or must fail.
-Our findings also explain why standard interpretability techniques like logit lens [70, 71] may fail to identify compositional structure in models trained on tasks with path ambiguities.
-3 A formulation of the coverage principle
-We now develop a formal framework of the coverage principle.
-We first provide an intuitive illustration with a two-hop compositional task, then generalize the intuitions to formalize the coverage principle for more general compositional tasks.
-Imagine a learner [^3] By “learner”, we mean any system that learns from data (e.g., neural networks).
-observing input-output pairs determined by a function $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$.
-The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is sampled from a finite set ${\mathcal{X}}$ [^4] For brevity, we use a shared token set ${\mathcal{X}}$.
-Position-specific domains ${\mathcal{X}}_{i}$ can be embedded as subsets of an enlarged token set $\tilde{{\mathcal{X}}}={\mathcal{X}}_{1}\cup{\mathcal{X}}_{2}\cup{\mathcal{X}}_{3}$ without loss of generality.
-Being unknown to the learner, let $f$ be a function that can be factorized as the composition of two primitive functions as $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig.
-2(a).
-How can the learner generalize by only seeing the input-output pairs?
-Our key intuition is that a learner exploits the compositional structure only when observations show that two fragments behave identically.
-For instance, assume that $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$ for $(x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{X}}^{2}$.
-From the underlying structure, these fragments are functionally equivalent, i.e., substituting one for the other never changes the output.
-If observations consistently support their equivalence, i.e., $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$ for observed $x_{3}$ values, this equivalence can be supported (Fig.
-1 (Left)).
-Intuitively, the learner would be able to harness this equivalence to generalize to an unseen input $(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime\prime})$, provided the learner has seen $f(x_{1},x_{2},x_{3}^{\prime\prime})$.
-In short, the learner can utilize the observed functional equivalence to correctly infer the output of an unseen input, if it can reach an observed input by safe substitutions supported by observations (Fig.
-1 (Right)).
+Pattern matching behaviors of LLMs on compositional tasks.
+It is well perceived that pattern matching alone is inadequate for systematic generalization (fodor1988connectionism), and modern LLMs display generalization abilities that seem to be far beyond what pattern matching alone can do, as measured by their remarkable performance on complex benchmarks (achiam2023gpt).
+However, a growing body of work has consistently reported that LLMs still fall short on benchmarks designed to test compositionality (hupkes2020compositionality), including mathematical reasoning (mirzadeh2024gsm), multi-hop reasoning (yang2024large; wang2024grokked), and more (lake2018generalization; kim2020cogs; csordas2022ctl++; dziri2024faith).
+This gap between their capabilities and pattern-matching behaviors on compositional tasks calls for a principled framework to define what pattern matching is and to what extent a model’s behavior can be attributed to pattern matching, but it is mostly discussed with behavioral studies under the context of a specifically designed benchmark.
+Our work addresses this gap by formally defining pattern matching, and we systematically analyze models’ behaviors with controlled tasks that are designed to isolate pattern-matching regimes grounded in our framework.
+Mechanistic interpretability.
+Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors (elhage2021mathematical; olsson2022context; nanda2023progress; elhage2022toy).
+Recent work analyzes how Transformer components are causally related to certain behaviors (meng2022locating; hanna2023does; goldowsky2023localizing).
+In particular, it is reported that in-domain compositional generalization can emerge through grokking, with identifiable intermediate state representations inside Transformers (wang2024grokked).
+Our framework complements these works by providing mechanistic insights about pattern matching.
+Our findings also explain why standard interpretability techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) may fail to identify state representations in models trained on tasks with path ambiguities.
+3 Formalizing Pattern Matching with Functional Equivalence
+We now develop a formal framework for pattern matching. We first provide an intuitive illustration with a two-hop structure, then generalize to arbitrary fixed-length discrete-sequence tasks.
+Imagine a learner observing data determined by $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$.
+The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is chosen from a finite set $\mathcal{X}$}.
+[^1] For brevity, we use a shared token set $\mathcal{X}$} in the main text.
+A more general notion using position-specific domains (e.g., ${\mathcal{X}}_{i}$’) can be used; see Sec.
+F.1.
+Suppose (unknown to the learner) that $f$ factorizes as the composition of two primitive functions, $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig.˜2(a).
+How can the learner generalize by only seeing the input-output patterns?
+Our key intuition is that a learner exploits the underlying patterns only when two fragments of inputs are observed to behave identically.
+For instance, assume that two fragments $(x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime})\in{\mathcal{X}}^{2}$ give the same implicit intermediate state upon the application of $f_{1}$, i.e., $f_{1}(x_{1},x_{2})=f_{1}(x_{1}^{\prime},x_{2}^{\prime})=b$.
+These fragments behave identically regardless of context, i.e., they are functionally equivalent: for all $x_{3}\in{\mathcal{X}}$, $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$.
+If observations consistently support their equivalence, i.e., $f(x_{1},x_{2},x_{3})=f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$ for observed $x_{3}$ values, this equivalence can be supported (Fig.˜1 Left).
+Intuitively, the learner would harness this equivalence pattern to predict $f(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime\prime})$, provided the training set contains $f(x_{1},x_{2},x_{3}^{\prime\prime})$.
+Equivalently, the learner can utilize the observed functional equivalence to correctly infer the output of an unseen input, if it can reach an observed input by ‘safe substitutions’ (edges in the substitution graph) supported by observations (Fig.˜1 Right), which we define as a pattern matching.
 Coverage is a set of such inputs that are reachable from an observed input through chains of functionally equivalent substitutions.
-The essence of our framework, the coverage principle, is that a learner can only generalize inside the coverage when the only source of compositional generalization is the observation and utilization of functional equivalence.
-We now formalize these concepts for an arbitrary fixed-length task with an arbitrary set of observations.
-We restrict our attention to single-token prediction tasks defined as a deterministic mapping $f:{\mathcal{X}}^{n}\to{\mathcal{X}}$, where ${\mathcal{X}}$ is a finite set of tokens.
-We also consider a fixed observation set $D\subset{\mathcal{X}}^{n}$, a collection of inputs that are allowed to be observed by the learner.
-Write ${\bm{x}}=(x_{1},\dots,x_{n})\in{\mathcal{X}}^{n}$ and, for a subset $I\subset[n]:=\{1,\dots,n\}$, let ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ be a subsequence of ${\bm{x}}$.
+Then, coverage sets a boundary for what can be achieved by solely relying on substituting observed, equivalently behaving patterns.
+In other words, a learner can only generalize inside the coverage when it relies on functional equivalence, which we will define as pattern matching.
+We now formalize these concepts for an arbitrary fixed-length task with an arbitrary set of discrete sequence observations.
+We restrict our attention to single‑token prediction tasks defined as a deterministic mapping $f:{\mathcal{X}}^{\ell}\to{\mathcal{X}}$, where $\mathcal{X}$} is a finite set of tokens.
+We also consider a fixed observation set $D\subset{\mathcal{X}}^{\ell}$, a collection of inputs that are allowed to be observed by the learner.
+Write ${\bm{x}}=(x_{1},\dots,x_{\ell})\in{\mathcal{X}}^{\ell}$ and, for a subset $I\subset[\ell]:=\{1,\dots,\ell\}$, let ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ be a subsequence of $\bm{x}$}.
 The first step is to formalize what it means for two subsequences to be functionally equivalent.
 Definition 3.1 (Functional $k$-equivalence).
-Fix a nonempty proper subset of indices $I\subsetneq[n]$.
-Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{|I|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset D$ to be an $I$-co-occurrence of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ if it satisfies ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[n]\setminus I}={\bm{x}}^{\prime}_{[n]\setminus I}$.
-Also, we say ${\bm{a}}$ and ${\bm{a}}^{\prime}$ are functionally $k$-equivalent at $I$ in $D$, denoted by ${\bm{a}}\equiv^{I}_{D}{\bm{a}}^{\prime}$, if it satisfies:
+Fix a nonempty proper subset $I$ of indices in $\ell$].
+Consider any set $S\subset{\mathcal{X}}^{\ell}$ of input sequences.
+[^2] The set $S$ can be any subset of the whole domain, e.g., $\mathcal{X}$$^{\ell}$ itself, the train dataset $D$, or whatever else.
+Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{\left|I\right|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ to be an $I$-co-occurrence of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ if it satisfies $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset S$ and ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[\ell]\setminus I}={\bm{x}}^{\prime}_{[\ell]\setminus I}$.
+Also, the subsequences $\bm{a}$} and $\bm{a}$$^{\prime}$ are said to be functionally $k$-equivalent at $I$ in $S$ and denoted by ${\bm{a}}\equiv^{I}_{S}{\bm{a}}^{\prime}$, if it satisfies:
   1. 1.
-    (Sufficiency of co-occurrences.) There are $k$ or more distinct $I$-co-occurrences of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$;
+    (Sufficiency of co-occurrences.) There are $k$ or more distinct $I$-co-occurrences of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$;
   2. 2.
-    (Consistency.) Every $I$-co-occurrence $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ of ${\bm{a}}$ and ${\bm{a}}^{\prime}$ in $D$ satisfies $f({\bm{x}})=f({\bm{x}}^{\prime})$.
+    (Consistency.) Every $I$-co-occurrence $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ satisfies $f({\bm{x}})=f({\bm{x}}^{\prime})$.
 In other words, two subsequences are functionally $k$-equivalent if they behave identically in the same contexts at least $k$ times.
@@
 Next, we ask: which inputs are reachable from observed data utilizing functional equivalence?
-To formalize this, we define the substitution graph: Let ${\mathcal{G}}_{D,k}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{n}$ of all possible inputs.
-Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[n]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$.
-With this substitution graph ${\mathcal{G}}_{D,k}$, we formally define the $k$-coverage as a set of inputs which are connected [^5] For an undirected graph ${\mathcal{G}}$, two vertices $u,v$ are connected if ${\mathcal{G}}$ contains a path between $u$ and $v$.
+To formalize this, we define substitution graph: Let ${\mathcal{G}}^{(D,k)}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{\ell}$ of all possible inputs.
+Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[\ell]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$.
+This process is illustrated on the right side of Fig.˜1, as a special case where $k=1$.
+With this substitution graph ${\mathcal{G}}^{(D,k)}$, we formally define the $k$-coverage as a set of inputs which are connected [^3] For an undirected graph $\mathcal{G}$}, two vertices $u,v$ are connected if $\mathcal{G}$} contains a path between $u$ and $v$.
 to at least one observed input as follows:
 Definition 3.2 ($k$-coverage).
-The $k$-coverage of $D$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in ${\mathcal{X}}^{n}$ that is connected to an input in $D$ on the substitution graph ${\mathcal{G}}_{D,k}$.
-Note that the notion of coverage is a stricter condition of the conventional definition of in-domain, which is obtained by random train/test split [19] or taking combinations of observed internal computations [17].
-We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms.
-Now, we formally phrase the coverage principle:
-Claim 3.3 (Coverage principle).
-Let a learner output $\hat{f}$ that is consistent with the training set $D$.
-If generalization is derived solely from observed functional $k$-equivalences and ${\bm{x}}\notin\mathrm{Cover}_{k}(D)$, the value $\hat{f}({\bm{x}})$ is unconstrained by $D$.
-This establishes a fundamental boundary: no learner can reliably generalize beyond the coverage as long as the only source of the generalization is functional equivalence.
-In Sec.
-5.1, we show that Transformers indeed fail to generalize beyond coverage when our random mapping dataset construction isolates functional equivalence as the sole generalization source.
-However, deeper insights come from asking when generalization success occurs in practice.
-While theory suggests that deep learning methods implicitly favor simpler solutions [72, 73, 74], which may involve leveraging functional equivalence, we still lack a systematic picture of how task structure, dataset, and model size interact to determine when such solutions emerge.
-We study this through experiments in Sec.
-5.
+The $k$-coverage of $D\subset{\mathcal{X}}^{\ell}$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in $\mathcal{X}$$^{\ell}$ connected to an ${\bm{x}}\in D$ with a path in the substitution graph ${\mathcal{G}}^{(D,k)}$.
+Note that the notion of coverage is a stricter condition of the canonical definition of in-domain (ID), which is obtained by random train/test split (wang2024grokked) or taking combinations of observed internal computations (dziri2024faith).
+In Sec.˜5, we demonstrate that learners may not necessarily generalize on data that are classified as ID in a canonical sense, but coverage can precisely explain when and why this occurs.
+We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms, and we demonstrate that the predictions made by our framework are invariant across model architecture and scale in Sec.˜6 and 7.
+Finally, $k$-coverage can be algorithmically determined for any fixed-length discrete sequence tasks (Alg.
+1), which we use for the analyses in the following sections.
+Now, we formally define pattern matching as a kind of generalization that is done by substituting functionally $k$-equivalent fragments of inputs, whose boundary is precisely the $k$-coverage defined above.
+This formalization enables us to predict, before testing, which inputs will be reliably handled through pattern matching and which require additional mechanisms.
+In other words, we view pattern matching as possible only within $k$-coverage, and generalization outside the coverage requires generalization mechanisms other than pattern matching, which we discuss in Sec.˜10 and I.
+In the following sections, we draw a systematic picture of how task structure, dataset, and model size interact to determine the success and failure of pattern matching through controlled setups, leading us to important and nontrivial insights.
 4 Experimental Setup
-$x_{1}$$x_{2}$$f_{1}$$b$$x_{3}$$f_{2}$$t$
+[…]
 (a) 2-Hop
@@
 (c) 3-Hop
-$x_{1}$$x_{2}$$f_{1}$$b$$x_{3}$$f_{2}$$t$
+[…]
 (d) Non-Tree
-Figure 2: Four synthetic compositional structures we study.
-Dataset construction
-We construct four synthetic compositional tasks: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig.
-2).
-We create random mappings from the product space of a token set to control generalization sources not attributable to compositional structures (i.e., commutativity).
-[^6] Such external generalization sources will be discussed in Sec.
-7 and App.
-H.
-We explain the dataset construction process using 2-Hop task (Fig.
-2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example.
-We construct training datasets by defining a token set with size $|{\mathcal{X}}|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$.
+Figure 2: Four synthetic task structures we study.
+Dataset construction.
+We construct four synthetic tasks with different structures: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig.˜2).
+To isolate functional-equivalence-based generalizations, we create random mappings from a product space of token sets to control the generalization sources not attributable to compositional structures (i.e., commutativity).
+We explain the dataset construction process using 2-Hop task (Fig.˜2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example.
+We construct training datasets by defining a token set with size $\left|{\mathcal{X}}\right|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$.
 We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset.
-See Sec.
-B.1 for more details of the dataset construction process.
-Training & evaluation
-Following Wang et al.
-[19], we train randomly initialized GPT-2 [75] models with 8 layers, 12 heads, and 768 dimensions (see Sec.
-B.2 for details).
-We construct two evaluation sets, each with 2,000 instances: (1) In-Domain (ID) Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{2})$ in 2-Hop task) are observed during training, but the specific combination was unseen.
-(2) Out-of-Domain (OOD) Test Set: at least one primitive function application is never observed during training.
-5 Results
-5.1 Predictive power of $k$-coverage on Transformer’s compositional generalization
-Figure 3: Left: Percentage of covered ID data depending on $k$ values and dataset size ($N$), for 2-Hop task ($|{\mathcal{X}}|=50$).
-Right: Test accuracy depending on $k$-cutoff values for 2-Hop task ($|{\mathcal{X}}|$=50, $N$=10k).
+See Sec.˜B.1 for more details of the dataset construction process.
+Training & evaluation.
+Following wang2022interpretability, we train randomly initialized GPT-2 (radford2019language) models with 8 layers, 12 heads, and 768 dimensions as a base model (see Sec.˜B.2 for details).
+We construct two evaluation sets, each with 2,000 instances: (1) ID Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{3})$ in 2-Hop task) are observed during training, but their specific combination was unseen.
+(2) Out-of-coverage (canonical OOD) Test Set: at least one primitive function application is never observed during training, which is used as a control group.
+5 Quantitative Analysis of Pattern Matching in Transformers
+5.1 Evidence Strength is Tightly Aligned to Pattern-Matching Success
+Figure 3: Left: Percentage of covered ID data depending on $k$ values and dataset size ($N$), for 2-Hop task ($\left|{\mathcal{X}}\right|=50$).
+Right: Test accuracy depending on $k$-cutoff values for 2-Hop task ($\left|{\mathcal{X}}\right|=50$, NN=10k).
 Each line represents a different training checkpoint.
-Note that OOD accuracy remains at chance level (=1/50) regardless of training time.
+Note that out-of-coverage ($k=0$) accuracy remains at chance level ($\approx 1/50$) regardless of training time.
 The bars below show the number of test data for each $k$-cutoff value.
-We now empirically validate the predictive power of the coverage principle by quantifying the correlation between $k$-coverage and model generalization performance.
-To this end, we implement and release a task-agnostic coverage determination algorithm (see App.
-C for implementation details) that can be applied to diverse compositional structures.
-Then, we analyze what fraction of ID test data of 2-Hop task with $|{\mathcal{X}}|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$.
-Fig.
-3 (Left) shows that at $N=5$k, every ID test example is already covered with minimal evidence ($k=1$).
-Ideally, a single witness of functional equivalence could suffice.
-However, experiments show that such minimal coverage alone is insufficient for efficient generalization in practice.
-To demonstrate this, we define a sample’s $k$-cutoff as the lowest $k$ for which an input lies in $k$-coverage, measuring the strength of evidence for functional equivalence.
+We first analyze the correlation between $k$-coverage and ID generalization performance of the GPT-2 model.
+To this end, we implement and release a task-agnostic coverage determination algorithm (see App.˜C) that can be applied to diverse compositional structures.
+Then, we analyze what fraction of ID test data of 2-Hop task with $\left|{\mathcal{X}}\right|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$.
+Fig.˜3 (Left) shows that at N=5N=5k, every ID test example is already covered with minimal evidence ($k=1$).
+Hence, in an ideal scenario where a single witness of functional equivalence suffices, training with the dataset as small as N=5N=5k will lead to perfect ID generalization.
+However, we demonstrate in our experiments that minimal coverage (i.e., $k=1$) alone is practically insufficient for ID generalization.
+To demonstrate this, let us fix a training dataset $D$ and define the $k$-cutoff of each input sequence as the lowest value of $k$ for which an input lies in $k$-coverage, measuring the strength of evidence for functional equivalence.
 For example, a $k$-cutoff of 3 means that an example is inside coverage with $k=3$ but not with $k=4$.
-For 2-Hop dataset with $N=10$k, we classify each ID test instance according to its $k$-cutoff, and track accuracy development for each group across 50k training epochs.
-As shown in Fig.
-3 (Right), generalization success shows a strong relationship with $k$-cutoff values.
+For out-of-coverage data, we define $k$-cutoff as 0.
+Then, for the 2-Hop dataset with N=10N=10k, we classify each ID test instance according to its $k$-cutoff, and track the accuracy development of the GPT-2 model for each group across 50k training epochs.
+As shown in Fig.˜3 (Right), ID test accuracy shows a strongly positive correlation with $k$-cutoff values.
 Test data with low $k$-cutoff values show delayed improvement even after extensive training, while examples with stronger evidence generalize much faster.
-Notably, OOD test examples (outside coverage) remain at chance-level ($1/50=0.02$) throughout training, empirically validating our coverage principle.
 These results yield two important insights.
-First, successful generalization in practice requires robust coverage so the model can confidently identify and utilize functional equivalence relationships.
-The $k$ parameter effectively quantifies this evidence strength, directly impacting generalization speed and reliability.
-Second, while our experiments use uniformly distributed data, the coverage principle explains why models struggle with generalizing long-tail distribution in imbalanced real-world data [76, 77, 78].
-Rare combinations naturally receive limited functional equivalence evidence (low $k$), placing them effectively outside practical coverage, despite technically being in-distribution.
-5.2 The clustering of latent representation drives generalization on coverage
+First, successful ID generalization in practice requires a robust coverage so the model can confidently identify and utilize functional equivalence relationships.
+The parameter $k$ effectively quantifies this evidence strength, directly impacting generalization speed and reliability.
+Second, while our experiments use uniformly sampled datasets, the results can explain why models struggle with generalizing long-tail distributions in imbalanced real-world data (mallen2022not; kandpal2023large; chang2024large).
+Despite technically being in-distribution, rare combinations naturally receive limited evidence of functional equivalence (low $k$), effectively and practically placing them outside the coverage.
+We believe that our insights will motivate future research on targeted data augmentation strategies to maximize $k$-coverage.
+5.2 Latent Representation Clusters Drive Pattern Matching on Coverage
 Figure 4: Left: Heatmap of Intra-Inter Cosine Gap (IICG) across layers and positions, sliced by $k$-cutoff.
@@
 Next, we investigate how the model internally represents functional equivalence for $k$-covered inputs.
-Specifically, we inspect a GPT-2 trained on 2-Hop task ($|{\mathcal{X}}|=50$, $N=10\text{k}$) for 50k epochs (corresponding to the yellow line in Fig.
-3 (Right)).
-[^7] Analyses for varying factors including task structures, entity set size ($|{\mathcal{X}}|$), dataset size ($N$), and training steps give consistent results; see App.
+Specifically, we inspect a GPT-2 trained on 2-Hop task ($\left|{\mathcal{X}}\right|=50$, $N=10\text{k}$) for 50k epochs (corresponding to the yellow line in Fig.˜3 (Right)).
+[^4] The analyses for varying factors including task structures, entity set size ($\left|{\mathcal{X}}\right|$), dataset size ($N$), and training steps give consistent results; see App.
 D.
@@
 We term this difference the intra–inter cosine gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation.
-Fig.
-4 (Left) reveals a clear relationship: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations.
-Critically, OOD examples show no clustering pattern at all, since they lack any functional equivalence evidence in the training data, as predicted by the coverage principle.
+Fig.˜4 (Left) reveals a positive correlation: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations.
+In contrast, out-of-coverage ($k=0$) examples exhibit no clustering pattern, as they lack evidence of functional equivalence in the training data.
 The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually.
-We verify that the representation clusters play a causal role in ID generalization with causal tracing [67, 66], a widely used mechanistic interpretability technique to identify Transformer circuits (see Fig.
-8 in App.
-D).
-Our findings extend the insights of Wang et al.
-[19] in several ways.
+We verify that the representation clusters play a causal role in pattern matching with causal tracing (goldowsky2023localizing; hanna2023does), a widely used technique to identify Transformer circuits (Fig.˜8).
+Our findings extend the previous insights from mechanistic interpretability literature (wang2024grokked) in several ways.
 First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps.
-Consequently, out-of-coverage inputs lack these unified circuits by construction, explaining their systematic failure on OOD examples.
-Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens [70] may fail to detect these functional equivalence representations despite their presence.
-See App.
-I for a detailed analysis.
-5.3 When does coverage lead to full ID generalization? A power-law lower bound
+Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens nostalgebraist2020logitlens may fail to detect these functional equivalence representations despite their presence (see App.˜E).
+6 Data Scaling Law of Pattern Matching Behaviors
+Our analysis in the previous section demonstrates that stronger functional equivalence evidence leads to better generalization.
+A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data?
+Intuitively, this requires the training set to support (strongly enough) the functional equivalence of every pair of inputs that shares the same intermediate state $b$.
+Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$.
+Assuming that generalization is constrained by the $k$-coverage, how should the train data size scale in the token set size $\left|{\mathcal{X}}\right|$ to achieve it completely?
+In practical terms, this question seeks to determine the amount of data necessary and/or sufficient to cover all ID combinations with $k$-coverage, which is crucial for understanding the data cost for pattern-matching generalization.
+To quantify the data cost, we establish and prove the following sample-complexity upper bound for learning a 2-Hop task (the complete statement and proof are provided in App.˜F):
 Figure 5: Left: Log-log plot of measured $\hat{N}_{\mathrm{req}}$ vs.
-token set size ($|{\mathcal{X}}|$) across three compositional tasks.
+token set size ($\left|{\mathcal{X}}\right|$) across three compositional tasks.
 The slope $c$ corresponds to the empirical power-law scaling exponent.
 Omitted points for 3-Hop are due to prohibitively large dataset requirements.
-Right: Power-law scaling remains invariant across GPT-2 model sizes (68M to 1.5B parameters) for 2-Hop task.
+Right: Power-law scaling behavior on 2-Hop task across varying GPT-2 model sizes (68M to 1.5B parameters) and Mamba model (For Mamba, we used 4 layers, a hidden dimension of 256, and a learning rate of 0.008, and $\hat{N}_{\mathrm{req}}$ is measured for only $\left|{\mathcal{X}}\right|\leq 100$, since a larger token set size led to training instability).
 $R^{2}>0.99$ for all linear fitting.
-Our analysis of $k$-coverage and representation clustering demonstrates that stronger functional equivalence evidence leads to better generalization.
-A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data?
-Intuitively, this requires the training set to support the functional equivalence of every pair of inputs that shares the same intermediate state $b$.
-Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$.
-Assuming a learner requires at least $k$ distinct pairs of evidence to establish functional equivalence of two fragments, how does the required dataset size scale with the token set size $|{\mathcal{X}}|$?
-In practical terms, this question addresses how much data is fundamentally needed for compositional generalization, which is central to understanding data scaling requirements for compositional tasks.
-For 2-Hop task, we derive the following result (full derivation in App.
-E):
-Result 5.1 (Power-law lower bound).
-Let $f:{\mathcal{X}}^{3}\!\to\!{\mathcal{X}}$ be a 2-Hop composition.
-Assume a learner recognizes functional equivalence of two subsequences only after observing them in at least $k$ distinct pairs of evidence (i.e., functionally $k$-equivalent).
-[^8] Note that this assumption is equivalent to modeling Fig.
-3 (Right) as a step function.
-Let $N_{\text{req}}(|{\mathcal{X}}|,k)$ be the smallest training dataset size that enables complete generalization to ID examples under this evidence threshold.
-Then, up to poly-logarithmic factors in $|{\mathcal{X}}|$, $N_{\text{req}}(|{\mathcal{X}}|,k)=\tilde{\Omega}\bigl{(}|{\mathcal{X}}|^{\,\alpha(k)}\bigr{)},\text{where}\;\alpha(k)=2.5-\frac{0.5}{k}.$
-Res.
-5.1 predicts that a learner relying on pattern matching requires training dataset size scaling at least quadratically with respect to the token set size, to fully generalize on ID test data.
-To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data.
-[^9] See App.
-F measurement details.
-Fig.
-5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs.
-$|{\mathcal{X}}|$ across different task structures.
-The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions of at least quadratic scaling.
+Theorem 6.1 (Informal; Corollaries˜F.9 and F.17).
+Consider a 2-Hop task with a token set of size $n$.
+For a uniformly randomly sampled train dataset $D$ of size $N$, consider a learner that generalizes within the $k$-coverage of $D$.
+Then, for large enough $n$, the learner achieves perfect ID generalization with high probability if $N\!\gtrsim\!n^{c}$ with $c=2.5-\frac{0.5}{k}$.
+In contrast, the learner (with $k\!\geq\!2$) does not achive perfect ID generalization with high probability for some 2-Hop task if $n^{2}\!\lesssim\!N\!\lesssim\!n^{c}$.
+Here, we ignore the polylogarithmic factors in $n$.
+Theorem 6.1 presents a tight sample complexity bound $\tilde{\Theta}(n^{c})$ for pattern-matching-based generalization, showing that the training dataset required to ensure full ID generalization grows polynomially in the token set size $n$, with an exponent $c\in[2,2.5)$.
+To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(\left|{\mathcal{X}}\right|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data (see App.˜G for the measurement details).
+Fig.˜5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs.
+$\left|{\mathcal{X}}\right|$ across different task structures.
+The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions.
 Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well.
 The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
-Strikingly, these exponents remain invariant across three different GPT-2 model sizes spanning a 20x range in parameters (from 68M to 1.5B) for all three tasks (Fig.
-5 (Right) and Tab.
-2 in App.
-F).
-This supports that the scaling relationship is primarily determined by data properties rather than model capacity, reinforcing our framework’s data-centric perspective.
-The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App.
-F).
-Overall, these results suggest that compositional generalization in Transformers requires dataset sizes scaling as a power-law with token set size.
-This has significant implications for how we approach complex reasoning tasks, suggesting that data curation strategies informed by coverage considerations may be more effective than raw parameter scaling.
-5.4 Path ambiguity hinders generalization and interpretability in Transformers
-Identical Intermediate StatesFunctional Equivalence$(x_{1},x_{2})$$(x_{1}^{\prime},x_{2})$$(x_{1},x_{2}^{\prime})$$(x_{1}^{\prime},x_{2}^{\prime})$$(x_{1},x_{2}^{\prime\prime})$$(x_{1}^{\prime},x_{2}^{\prime\prime})$
-Figure 6: Left: In Non-tree task, the coverage principle predicts that representations of input subsequences with the same intermediate state $b=f_{1}(x_{1},x_{2})$ split into multiple context-dependent state representations, conditioned on $x_{2}$ value.
-Middle: ID test accuracy after standard training (100 epochs post training accuracy$>$0.99), showing Non-tree task significantly underperforms 2-Hop task ($|{\mathcal{X}}|=50$).
-Right: IICG heatmap from a model that achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|{\mathcal{X}}|=50$, $N=50\text{k}$).
-Many real-world reasoning tasks involve computational structures where a single variable affects the output through multiple paths, creating what we call path ambiguity.
-Our framework predicts that in such cases, Transformers struggle with forming unified representations of intermediate states that are theoretically equivalent, instead developing context-dependent state representations that vary based on the input context.
-In this section, we analyze Non-tree task (Fig.
-2(d)) as a case study, where $x_{2}$ affects the output through two paths: as input to $f_{1}$ and directly to $f_{2}$.
-Following the definition of functional equivalence (Def.
-3.1), this path ambiguity prevents the model from establishing the functional equivalence of two subsequences $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ that produce the same intermediate state $b$, unless they also share the same $x_{2}$ value ($x_{2}=x_{2}^{\prime}$).
-Consequently, the model creates context-dependent state representations rather than unifying them to represent true intermediate state $b$, as illustrated in Fig.
-6 (Left).
-In consequence, the path ambiguity hinders both generalization performance and efficiency, as the model must establish functional equivalence for each $x_{2}$-conditioned equivalent pair.
-Fig.
-6 (Middle) shows that GPT-2 can fully generalize on ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data.
-[^10] For $|\mathcal{X}|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times|\mathcal{X}|^{3}\;\approx\;61\text{k}$ distinct ID triples).
-Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig.
-17 in App.
+These exponents remain invariant across three different GPT-2 model sizes spanning a 20x range in parameters (from 68M to 1.5B) for all three tasks (Fig.˜5 Right and Tab.˜2 in App.˜G).
+We also show that the exponent measured with a Mamba model (4 layers and a hidden dimension of 256) falls inside the boundary predicted by the theory (same figure).
+Interestingly, the result in Fig.˜6 (Middle) demonstrates that with increasing training dataset size $N$, there is a sharp phase transition from ID generalization failure to complete success near $N=20\text{k}$.
+Overall, the results support that the data scaling law is primarily determined by data properties rather than model capacity or architectures, and additional generalization mechanisms will be required to achieve milder scaling laws on such compositional tasks.
+[^5] The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App.
 G).
-Extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96.
-However, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig.
-6 (Right)).
-In contrast, grouping by $x_{2}$-conditioned intermediate state ($(b,x_{2})$) leads to high IICG scores, suggesting the formation of context-dependent state representations.
-This context-dependence raises an interpretability concern, as standard linear probing-based techniques like logit lens [70, 71] would not reliably identify intermediate states.
-Moreover, our analysis provides insights into why LLMs fall short in planning tasks [29, 31, 30], as planning often requires correct tracking of intermediate states, which can influence outcomes through multiple paths.
-5.5 CoT supervision improves data efficiency but still struggles with path ambiguity
+We note that our result aligns with the practical observation that parameter scaling does not significantly improve the multi-hop reasoning capability of LLMs (yang2024large) and the data-hungry nature of compositional tasks (lake2018generalization), suggesting that these could be partly attributed to pattern-matching behaviors.
+We leave further analysis of the connection between scaling behavior and pattern matching as an exciting future research direction.
+7 Path Ambiguity Problem as a Failure Case of Pattern Matching
+We identify a path ambiguity problem with our framework, a previously uncharacterized failure mode that pattern matching struggles with task structures where a single variable affects the output through multiple paths.
+In this section, we analyze Non-tree task (Fig.˜2(d)) as a case study, where $x_{2}$ affects the output through two paths, as input to $f_{1}$ and directly to $f_{2}$.
+Unlike in the 2-Hop case, one cannot establish the functional equivalence of two subsequences $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ that produce the same intermediate state $b$, unless they also share the same $x_{2}$ value ($x_{2}=x_{2}^{\prime}$).
+It is because $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ are not guaranteed to behave identically (i.e., $f(x_{1},x_{2},x_{3})$ is not necessarily equal to $f(x_{1}^{\prime},x_{2}^{\prime},x_{3})$) when $x_{1}\neq x_{2}$.
+Consequently, we can predict that Transformers trained on the Non-tree will create context-dependent state representations that are conditioned on $x_{2}$ values, failing to unify them to represent the true intermediate state $b$ (Fig.˜6 Left).
+[…]
+Figure 6: Left: In Non-tree task, the representations of input subsequences with the same intermediate state $b=f_{1}(x_{1},x_{2})$ are split into multiple context-dependent state representations, conditioned on $x_{2}$ value.
+Middle: ID test accuracy after standard training with varying training dataset size ($\left|{\mathcal{X}}\right|=50$, evaluated 100 epochs after training accuracy reaches 0.99).
+Observe a sharp transition from ID generalization failure to complete success near $N=20\text{k}$ for 2-Hop, which does not occur in Non-tree task.
+Right: IICG heatmap from a model that achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $\left|{\mathcal{X}}\right|=50$, $N=50\text{k}$).
+Experiments show that the path ambiguity indeed hinders both generalization on the ID test set and the interpretability of intermediate state representations, as the model now establishes functional equivalence for each $x_{2}$-conditioned equivalent pair.
+Fig.˜6 (Middle) shows that GPT-2 can fully generalize on the ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data.
+[^6] For $\left|{\mathcal{X}}\right|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times\left|{\mathcal{X}}\right|^{3}\;\approx\;61\text{k}$ distinct ID triples).
+Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig.˜18), and the Mamba model used in Sec.˜6 shows the same trend of generalization failure (Fig.˜19).
+In addition, extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96; however, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig.˜6 Right).
+In contrast, grouping by $x_{2}$-conditioned intermediate state $(b,x_{2})$ leads to high IICG scores, showing the formation of context-dependent state representations.
+This context-dependence due to path ambiguity raises an interpretability concern, as standard linear probing-based techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) would not reliably identify intermediate states when a model relies on pattern matching.
+Hence, a generalization mechanism other than pattern matching will be required for a robust ID generalization on complex task structures that require the access and update of intermediate states through multiple paths (e.g., planning tasks (ruis2020benchmark; kambhampati2024position; valmeekam2023planbench)), where further characterization of this problem remains as an exciting future direction.
+8 CoT Improves Data Efficiency but Path Ambiguity Persists
 Figure 7: Left: Power-law scaling of required dataset size vs.
@@
 $R^{2}>0.98$ for all linear fits.
-Middle: Comparison of ID test Accuracy of Non-tree task ($|{\mathcal{X}}|=50$) with and without CoT supervision (100 epochs post training accuracy$>$0.99).
-Right: IICG score comparison for Non-tree and 2-Hop task with CoT supervision ($|\mathcal{X}|=50$, $N=10\text{k}$).
+Middle: Comparison of ID test Accuracy of Non-tree task ($\left|{\mathcal{X}}\right|=50$) with and without CoT supervision.
+Right: IICG score comparison for Non-tree and 2-Hop task with CoT supervision ($\left|{\mathcal{X}}\right|=50$, $N=10\text{k}$).
 The scores are measured at each layer of intermediate state position $b$, based on two grouping strategies: $b$ and $(b,x_{2})$.
 Models are trained for 100 epochs after reaching training accuracy > 0.99.
-CoT supervision [27, 28] dramatically improves performance on multi-step reasoning tasks.
-We investigate how CoT interacts with the coverage principle and whether it can address the challenges observed in Sections 5.3 and 5.4.
+CoT supervision (wei2022chain; kojima2022large) dramatically improves performance on multi-step reasoning tasks.
+We investigate how CoT interacts with our framework and whether it can address the challenges observed in Sections 6 and 7.
 Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example.
-This explicit supervision substantially improves data efficiency (Fig.
-7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in 3-Hop task, aligning with previous studies on the sample efficiency of CoT [79, 80, 81].
-Remarkably, the scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision.
+This substantially improves data efficiency (Fig.˜7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in the 3-Hop task, aligning with previous studies on the sample efficiency of CoT (srivastava2022beyond; kim2024transformers; wen2024sparse).
+The scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision.
 We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
-However, we find models trained with CoT supervision still struggle with the path ambiguity in Non-tree task.
-Despite showing improvements with CoT supervision, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig.
-7 (Middle).
+However, we find the path ambiguity problem persists even with CoT supervision.
+Despite showing improvements, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig.˜7 (Middle).
 IICG analysis (Right) reveals that the model’s representations remain partially context-dependent.
-For 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve.
+For the 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve.
 In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model.
 We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$.
-Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec.
-5.4.
-Our analysis likely explains why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data [82].
-6 A taxonomy for understanding generalization mechanisms
-Our coverage analysis reveals fundamental limits of pattern matching for compositional generalization.
-This raises a broader question: What other mechanisms allow neural networks to go beyond coverage boundaries?
-We sketch a preliminary taxonomy that distinguishes three complementary mechanisms and offers hypotheses about why models sometimes succeed beyond coverage [^11] The taxonomy is not exhaustive; Appendix H details limitations and open questions..
-Type-I: Structure-based generalization occurs when models identify and utilize functionally equivalent components based on how primitive functions are composed.
-This is precisely what we formalize through the coverage principle: models learn that different input fragments yield identical results in shared contexts, enabling generalization to new fragment combinations.
-Crucially, this generalization remains bounded by coverage, and reliable generalization fails without sufficient functional equivalence evidence.
-Type-I describes the ceiling of pattern matching without explicit variable binding.
-Type-II: Function property-based generalization exploits intrinsic properties of individual primitive functions (e.g., algebraic invariances such as commutativity), group-theoretic structure in modular arithmetic [63], or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present [81]).
-Unlike Type-I, this can transcend coverage limitations by leveraging properties that hold across all inputs of a primitive.
-The Reversal Curse exemplifies the layered nature of compositional challenges across multiple generalization types.
-The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”.
-Property-based approaches like bidirectional training partially succeed [83] by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data.
-However, recent evidence [42] shows that models still struggle once entities switch syntactic roles, indicating a remaining challenge in variable binding.
-[^12] See [42] for controlled concept–surface dissociation experiments.
-Type-III: Shared-operator generalization emerges through reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$).
-Recurrent architectures [84] exemplify this through weight sharing across time steps, enabling processing of variable-length sequences [85].
-In Transformers, inductive biases towards reuse of the same computation through parameter sharing [86, 87, 19] can improve generalization on compositional tasks beyond coverage.
-We interpret this mechanism as exploiting structural repetition without requiring Type-II’s algebraic property understanding.
-Distinguishing mechanisms from phenomena
-Prior categorizations focus on observed phenomena, for example, “systematic” versus “mix-and-match” [12].
-We instead categorize the underlying mechanisms, as our coverage principle offers quantitative predictions for Type-I boundaries while identifying when Type-II or Type-III mechanisms become necessary.
-Many influential studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure.
-We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.
-Implications and future directions
-Real compositional tasks typically involve combinations of all three types, making it difficult to isolate why models succeed or fail.
-We suggest diagnostic approaches based on the proposed taxonomy: coverage ablation tests for Type-I, algebraic identity probes for Type-II, and parameter-sharing ablations for Type-III.
-While preliminary, this taxonomy provides a conceptual scaffold for understanding the fundamental capabilities and limitations of neural networks on compositional reasoning.
-This work isolates and formalizes structure-based generalization (Type-I) to clarify its specific boundaries.
-When models succeed beyond our coverage predictions, e.g., through ignoring distractors without seeing all combinations or leveraging parameter sharing, we view these as exploiting Type-II or Type-III mechanisms rather than contradicting our framework.
-Our focus on Type-I suggests that fundamental systematicity challenges remain as long as models rely primarily on pattern matching, requiring architectural innovations that harness all three mechanisms and, crucially, incorporate explicit variable binding.
-We hope this preliminary taxonomy serves as a conversation starter, and confirming or refuting its utility is an empirical matter that we invite the community to explore.
+Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec.˜7.
+Our analysis may explain why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data (stechly2024chain), where we leave further analysis as future work.
 9 Discussion
-Systematicity without variable binding
-The coverage principle pinpoints a fundamental question in the systematicity debate: How far can a neural network generalize compositionally when it relies only on pattern matching and has no mechanism for explicit variable binding?
-By formalizing the exact conditions under which pattern matching succeeds, we draw a principled boundary around what can be learned from input–output pairs alone.
-This boundary clarifies the limitations that Fodor and Pylyshyn [9] and Marcus [37] highlighted decades ago.
-Our results therefore complement approaches that encode compositional structure directly, whether through symbolic interfaces, slot-based representations, or neural modules [56, 57, 58, 59, 60, 88], and delineate what must still be overcome to reach truly systematic generalization.
-Practical implications
-Our framework provides valuable insights into various observed LLM behaviors.
-First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization [43].
-Next, our framework explains why models struggle with generalizing long-tail knowledge [76, 77, 78], as low-frequency combinations naturally receive limited functional equivalence evidence.
-Similarly, failures in complex planning tasks [30, 82, 31, 89], even with CoT techniques, might be attributed to path ambiguities.
-Moreover, our coverage principle predicts the reversal curse phenomenon [16, 42, 83], as pattern-matching models fundamentally cannot generalize to reversed relationships without explicit functional equivalence evidence in training data.
-Beyond explaining failures, our framework helps understand when and why standard interpretability techniques like linear probing or logit lens can fail.
-In addition, our analysis suggests targeted data augmentation strategies that maximize coverage by ensuring diverse shared contexts for functionally equivalent components.
-This can explain the success of strategic data augmentation methods [90, 91, 49, 48] and offers principled guidance.
-Finally, although our empirical study focuses on GPT-2–style Transformers, the coverage principle itself makes no architectural assumptions, and it applies to any learning system whose generalization mechanism is primarily pattern matching.
-Extending the analysis to recurrent (e.g., LSTM [84]), state-space [92], or convolutional [93] architectures, and more realistic data where multiple computational structures coexist within the same dataset [12, 15, 94] remains promising future work.
+9.1 Practical Implications
+Although the experiments in this work have focused on synthetic setups to dissect purely pattern-matching behaviors, our results provide valuable practical implications for understanding and improving LLMs on natural language tasks.
+First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization (lake2017building).
+For instance, our scaling-law analysis (Theorem˜6.1) provides a quantitative explanation for the observed behavior that the data demand for generalization on multi-hop natural language data dramatically increases with the number of hops (yao-etal-2025-language-models).
+Such a data-hungry nature of compositional tasks is also well observed in semantic parsing (dong-lapata-2016-language), where it has been shown that the diversity of data (i.e., the same component is shown in diverse contexts) is more important than the sheer size of the dataset for better generalization (keysers2019measuring).
+Based on these insights, we believe our results will motivate strategic data augmentation methods that seek to maximize coverage by ensuring diverse shared contexts for functionally equivalent components (andreas2019good).
+Second, our framework helps interpret various real-world failure cases observed in LLMs as originating from their reliance on pattern matching.
+For example, our framework aligns with the reversal curse phenomenon, which is the failure to automatically infer that ‘$A$ is a child of $B$’ by observing that ‘$B$ is a parent of $A$’ (berglund2023reversal).
+This is because pattern matching is fundamentally incapable of generalizing to the reversing relation without explicit functional equivalence evidence for that in the training dataset.
+Similarly, our framework can explain the notoriously hard problem of logical negation in LLMs (truong-etal-2023-language).
+A purely pattern-matching learner cannot deduce that the negation rule (i.e., if a statement ‘$p$’ is true, then ‘not $p$’ is false and vice versa) should apply to any well-formed statement, including the ones that aren’t observed, through finite observations.
+Finally, difficulties in solving complex planning tasks with LLMs (valmeekam2023planbench; stechly2024chain; kambhampati2024position; wang2024planning) might be partially attributed to path ambiguities, since such tasks likely require the correct tracking of intermediate states, which can be affected by various computational paths.
+Hence, we believe our framework and analyses can provide important practical insights, which future work can extend to investigate LLMs’ pattern-matching behaviors under more realistic setups.
+9.2 Towards a Taxonomy of Generalization Mechanisms
+Natural language tasks possess algebraic and structural properties that differ from the random mappings we studied.
+For example, the same knowledge can be used in any hop of the multi-hop reasoning in practice, unlike our 2-Hop task that used $f_{1}\neq f_{2}$.
+In practice, a learner may reasonably harness such properties of a given task to generalize beyond the boundary of pattern-matching defined by coverage.
+Therefore, it is natural to ask: What generalization mechanisms enable generalization beyond the coverage boundary?
+While a complete answer requires future work, we outline a mechanism-based taxonomy as a starting point for a constructive categorization of distinct generalization mechanisms beyond pattern matching:
+  • Functional equivalence-based generalization, the main focus of this work.
+  • Function property-based generalization leverages algebraic invariances of individual primitive functions, e.g., commutativity or input irrelevance, where certain arguments never affect the output.
+This distinguishes it from pattern matching, as it leverages a primitive function’s global property that holds across all inputs, not only those observed.
+  • Shared-operator generalization leverages the reuse of the same computation across positions (e.g., when $f_{1}=f_{2}$ in a two-hop task), which may be important in compositional generalization.
+For example, it is known that Transformers with inductive biases towards computation reuse can improve generalization on compositional tasks (csordas2021devil).
+We envision this taxonomy as a foundational diagnostic that quantifies when pattern matching suffices and when other mechanisms are required.
+See App.˜I for a complete discussion on the categorization of generalization mechanisms.
 10 Conclusion
-We introduce the coverage principle, a data-centric framework that specifies when pattern-matching learners can and cannot generalize compositionally.
-Our theoretical analysis and controlled experiments show that a Transformer’s success is tightly predicted by coverage: two-hop tasks obey a superquadratic data scaling law that does not improve with 20x parameter scaling, path ambiguity fragments internal representations, and CoT supervision helps only inside the coverage boundary.
-These results isolate Type I structure-based generalization and reveal its limits.
-Our taxonomy then situates two additional mechanisms, Type II property exploitation and Type III shared-operator reuse, that help explain coverage “violations” reported in the literature.
-Yet all three mechanisms may still rely on sophisticated pattern matching rather than explicit variable binding, leaving the fundamental systematicity challenge unresolved.
-Hence, the systematicity challenge posed by Fodor and Pylyshyn [9] and by Marcus [37] remains open.
-Matching human-like compositionality will likely require architectures that can bind and manipulate symbols independently of surface form, not merely scale up or refine current pattern-matching strategies.
-We hope the coverage principle and the accompanying taxonomy provide a clear target and a roadmap for the next stage of innovation.
-Acknowledgments and Disclosure of Funding
-We would like to thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
+In this work, we formalized a framework for characterizing pattern matching.
+Our theoretical and experimental analyses yielded quantitative and predictive insights into modern neural networks’ pattern-matching behaviors, moving beyond post-hoc accounts of a model’s behavior on compositional tasks: (i) the alignment of instance-wise success with the strength of functional-equivalence evidence (Sec.˜5), (ii) the theoretical identification and empirical verification of a sharp sample complexity bound for complete ID generalization on the 2-Hop task through pattern matching (Sec.˜6), and (iii) the identification of the path ambiguity problem that impairs accuracy and interpretability even under high coverage and CoT supervision (Sec.˜7 and 8).
+We anticipate that future work will build on this foundation, towards a more complete and constructive understanding of compositional generalization and its failures.
+Reproducibility Statement
+All codes for dataset generation, training, and analysis can be found at: https://github.com/kaistAI/coverage-principle.
+The Use of Large Language Models (LLMs)
+This work deployed LLMs to proofread for grammatical errors and improve the quality of writing.
+Acknowledgements
+We thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
+Contents of the Appendix
 Appendix A Limitations
 We deliberately restrict to synthetic tasks to isolate structure-based limits without confounds from lexical or domain priors.
-We leave extending the coverage analysis to natural data as a future work.
-Additionally, our experiments focus on autoregressive Transformer architectures, and the applicability of the coverage principle to other architectures (e.g., state-space models) remains to be validated.
+We leave extending the coverage analysis to discrete sequence tasks with variable lengths and more natural data as future work.
+Additionally, our experiments focus on autoregressive architectures, and the applicability of the coverage principle to broader architectures remains to be validated.
 Appendix B Detailed Experimental Setup
-B.1 Dataset construction details
+B.1 Dataset Construction
 We now provide detailed information about our dataset construction process.
@@
 Vocabulary and Token Representation
-For a task with token set size $|\mathcal{X}|$, we create $|\mathcal{X}|$ special tokens of the form <t_0>, <t_1>, $\ldots$, <t_$(|\mathcal{X}|-1)$>, which we append to the standard GPT-2 vocabulary.
-We also add special tokens </a> to mark the end of sequences.
-For Chain-of-Thought (CoT) experiments, intermediate computations are represented in the target sequence as the actual intermediate token.
+[…]
 Function Construction
@@
 We carefully construct test sets to evaluate the model’s generalization capabilities across different coverage conditions. Our test sets contain:
-  • In-Domain (ID) Test Set: Combinations not seen during training but where both primitive operations were observed in other contexts.
+  • In-Domain (ID) Test Set: Combinations unseen during training but where both primitive operations were observed in other contexts.
 These examples may lie within the coverage as defined by our framework.
-  • Out-of-Domain (OOD) Test Set: Examples where at least one primitive operation was never observed in training. These fall outside the coverage boundary.
+  • Out-of-coverage (canonical OOD) Test Set: Examples where at least one primitive operation was never observed in training. These fall outside the coverage.
 Input-Output Format
@@
 For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters).
-As shown in Tab.
-1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768.
+As shown in Tab.˜1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768.
 The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768.
 Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600.
-The implementation follows the codebase from [19].
+The implementation follows the codebase from (wang2024grokked).
 We train all models using the AdamW optimizer with beta values of (0.9, 0.999) and epsilon of 1e-8.
@@
 Appendix C Implementation Details for the Coverage Determination Algorithm
-Input: Training examples $\mathcal{D}=\{(x_{i},f(x_{i}))\}$, where $x_{i}\in\mathcal{X}^{n}$ and $f(x_{i})\in\mathcal{X}$
+Input: Training examples $D=\{(x_{i},f(x_{i}))\}_{i=1}^{N}$, where $x_{i}\in{\mathcal{X}}^{n}$ and $f(x_{i})\in{\mathcal{X}}$
 Minimum evidence threshold $k\geq 1$
-Output: Coverage set $\text{Cover}(\mathcal{D})$
+Output: Coverage set $\text{Cover}(D)$
 /* STEP 1: Build behavior maps for each subsequence pattern */
-foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
-       $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in\mathcal{D}\}$
+foreach nonempty subset $I\subsetneq[n]$ do
+ $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in D\}$
 end foreach
 /* STEP 2: Identify functionally equivalent subsequences */
-foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
+foreach nonempty subset $I\subsetneq[n]$ do
  $UF_{I}\leftarrow\text{new UnionFind}()$
  foreach pair of subsequences $(\alpha,\beta)$ in $\text{Behavior}_{I}$ do
-[…]
+    $\text{SharedComplements}\leftarrow$ complements observed with both $\alpha$ and $\beta$
     if No contradictions in SharedComplements and matching evidence $\geq k$ then
@@
 foreach pair of inputs $(x,y)$ with $f(x)=f(y)$ do
-       foreach subset $I$ where $x$ and $y$ differ only on indices in $I$ do
+ foreach subset II where $x$ and $y$ differ only on indices in II do
     if $\text{EquivClasses}_{I}.\text{Find}(x_{I})=\text{EquivClasses}_{I}.\text{Find}(y_{I})$ then
@@
 /* STEP 4: Determine coverage */
-$\text{Cover}(\mathcal{D})\leftarrow\bigcup_{x\in\mathcal{D}}$ ConnectedComponent$(G,x)$
-return $\text{Cover}(\mathcal{D})$
+$\text{Cover}(D)\leftarrow\bigcup_{x\in D}$ ConnectedComponent(G,x)(G,x)
+return $\text{Cover}(D)$
 Algorithm 1 $k$-Coverage Determination Algorithm
-Algorithm 1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
+Alg.˜1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
 Stage 1: Behavior mapping
@@
 This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output.
-Our implementation uses parallel processing to efficiently construct this graph even for large datasets.
+Our implementation uses parallel processing to efficiently construct this graph, even for large datasets.
 Stage 4: Coverage computation
@@
 D.1 Causal Tracing Methodology
-To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference [67, 66].
-Specifically, we measure the causal effect using the indirect effect metric defined in [95].
+To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference (goldowsky2023localizing; hanna2023does).
+Specifically, we measure the causal effect using the indirect effect metric defined in (sharma2024locating).
 This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization.
@@
 Conversely, an IE value close to 0 indicates that the state has minimal causal impact on the prediction.
-In our experiments, we apply causal tracing to analyze different subsets of test data categorized by their $k$-cutoff values, where $k$ represents the minimum evidence threshold required for functional equivalence (as defined in Sec.
-3 of the main text).
+In our experiments, we apply causal tracing to analyze different subsets of test data categorized by their $k$-cutoff values, where $k$ represents the minimum evidence threshold required for functional equivalence (as defined in Sec.˜3 of the main text).
 This allows us to correlate the strength of functional equivalence evidence with the formation of unified internal representations.
@@
 For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
-Figure Fig.
-9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases.
-The causal tracing results in Fig.
-10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
+Figure Fig.˜9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases.
+The causal tracing results in Fig.˜10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
 Figure 9: IICG heatmap across different token set sizes, showing consistent representation clustering patterns.
@@
 These results demonstrate that the formation of clustered intermediate state representations and their causal role in compositional generalization is a consistent phenomenon across different task structures, supporting the generality of our findings beyond the 2-Hop task.
-Appendix E Derivation of Result 5.1 (power–law lower bound)
-Problem setting
-Let ${\mathcal{X}}$ be a finite token set with cardinality $|{\mathcal{X}}|$. The target mapping
-$f:{\mathcal{X}}^{3}\longrightarrow{\mathcal{X}},\qquad f(x_{1},x_{2},x_{3})=f_{2}\bigl{(}f_{1}(x_{1},x_{2}),\,x_{3}\bigr{)}$
-is a two-hop composition of unknown primitives $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$.
-Write $b=f_{1}(x_{1},x_{2})\in{\mathcal{X}}$ for the intermediate state.
-Throughout the proof we impose two assumptions.
-  1. (A1)
-    Balanced classes: Each intermediate value $b\in{\mathcal{X}}$ is realized by exactly $|{\mathcal{X}}|$ first-hop pairs, i.e., the sets $E_{b}:=\{(x_{1},x_{2})\in{\mathcal{X}}^{2}:f_{1}(x_{1},x_{2})=b\}$ all have size $|{\mathcal{X}}|$ and form a partition of ${\mathcal{X}}^{2}$.
-  2. (A2)
-    Uniform training sampler: The training set $D$ contains $N$ triples drawn uniformly with replacement from the domain ${\mathcal{X}}^{3}$.
-Functional $k$-equivalence and learner model
-For two first-hop fragments $a,a^{\prime}\in{\mathcal{X}}^{2}$ define
-$a\sim a^{\prime}\quad:\Longleftrightarrow\quad f_{1}(a)=f_{1}(a^{\prime}).$
-They are functionally $k$-equivalent w.r.t.
-$D$ (denoted $a\equiv^{k}_{D}a^{\prime}$) when there exist $k$ distinct contexts $c_{1},\dots,c_{k}\in{\mathcal{X}}$ such that for every $r\leq k$ both $(a,c_{r})$ and $(a^{\prime},c_{r})$ appear in $D$ and $f(a,c_{r})=f(a^{\prime},c_{r})$.
-We call each $(a,c_{r}),(a^{\prime},c_{r})$ an evidence pair.
-The coverage principle by itself is only a necessity statement: outside the $k$-coverage region a purely pattern matching learner’s predictions are unconstrained.
-To convert this into a sufficient data condition we adopt an explicit inductive bias, matching the premise of Result 5.1:
-Learner assumption: Whenever two fragments become linked by $k$ independent (i.e., pairwise from distinct contexts) evidence pairs, the learner treats them as functionally equivalent; the learner also propagates equivalence transitively along chains of such links.
-With this rule in place the relevant structure inside each class $E_{b}$ is the $k$-evidence graph: vertices are the $|{\mathcal{X}}|$ first-hop pairs in the class and an edge connects two vertices whenever the pair is observed at least $k$ times in shared contexts.
-If that graph is connected (every vertex reachable from every other), then every fragment in $E_{b}$ is linked by a chain of $k$-evidence steps and is therefore recognized as equivalent by the learner.
-Hence
-Data sufficiency criterion: If the $k$-evidence graph of each class $E_{b}$ is connected, the learner generalizes perfectly to all in-domain (ID) inputs.
-This criterion requires connectivity, and we aim to derive the condition to yield the connectivity with high probability using Erdős–Rényi model [96, 97].
-The minimal dataset size achieving this with high probability is denoted $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$.
-Note that the $k$-evidence graph on $E_{b}$ is the restriction of the substitution graph $\mathcal{G}_{D,k}$ (Def.
-3.2) to the vertex set $E_{b}\times\{x_{3}\}$ for any fixed $x_{3}$.
-Connectivity of every class therefore implies that every first-hop fragment lies in the same connected component as some training input, i.e., the entire in-domain set is contained in $k$-coverage.
-Under the learner assumption, this is both necessary and sufficient for perfect ID generalization, yielding Res.
-5.1.
-E.1 Step 1: probability of a single evidence pair
-Evidence pair probability
-Fix two distinct first–hop fragments $i,j\in E_{b}$ and a context $c\in{\mathcal{X}}$.
-We want to find $p_{1}$, the probability that context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$:
-$p_{1}:=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
-Let $q:=1/|{\mathcal{X}}|^{3}$ denote the probability of drawing any specific triple in a single draw. Using the inclusion–exclusion principle:
-$p_{1}$ $=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
-$=1-\Pr[(i,c)\not\in D]-\Pr[(j,c)\not\in D]+\Pr[(i,c)\not\in D\text{ and }(j,c)\not\in D]$
-$=1-(1-q)^{N}-(1-q)^{N}+(1-2q)^{N}$ (S1.1)
-For $q\ll 1$ (which holds when $|{\mathcal{X}}|$ is large), we can use Taylor expansion:
-$(1-q)^{N}$ $=1-Nq+\frac{N(N-1)}{2}q^{2}+O(q^{3})$
-$(1-2q)^{N}$ $=1-2Nq+\frac{N(N-1)}{2}(2q)^{2}+O(q^{3})$ (S1.2)
-Substituting these approximations:
-$p_{1}$ $=1-2\left(1-Nq+\frac{N(N-1)}{2}q^{2}\right)+\left(1-2Nq+2N(N-1)q^{2}\right)+O(q^{3})$
-$=1-2+2Nq-N(N-1)q^{2}+1-2Nq+2N(N-1)q^{2}+O(q^{3})$
-$=N(N-1)q^{2}+O(q^{3})$
-$=\frac{N(N-1)}{|{\mathcal{X}}|^{6}}+O\left(\frac{1}{|{\mathcal{X}}|^{9}}\right)$
-$=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))$ (S1.3)
-Therefore, in the regime of interest ($N\gg|{\mathcal{X}}|$ but $N\ll|{\mathcal{X}}|^{3}$):
-$\boxed{p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))}$
-Equation (E.1) gives an exact expression for the probability (up to lower-order terms) that the single context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$.
-Remark on “lucky coincidences”
-Because functional $k$-equivalence demands consistency across all $k$ evidences, a single coincidental equality $f(i,c)=f(j,c)$ with $i\not\sim j$ can only masquerade as evidence when $k=1$ and the dataset lacks any contradicting context $c^{\prime}$.
-For $k\geq 2$ the joint probability that two independent contexts simultaneously produce such coincidences is $|{\mathcal{X}}|^{-k}$ per fragment pair and hence negligible relative to the isolate probability once $N=\Omega(|{\mathcal{X}}|^{2})$.
-Consequently, the effects of such “lucky coincidences” on the lower bound in Res.
-5.1 can be neglected.
-E.2 Step 2: probability of observing $k$ evidences for one fixed pair
-Having established the probability $p_{1}$ for a single evidence pair in Step 1, we now derive $p_{k}$, the probability that a dataset provides at least $k$ distinct contexts as evidence for the functional equivalence of two fragments $i$ and $j$.
-This probability will determine the edge probability in the random graph model analyzed in Step 3.
-Indicators for one fragment pair
-Fix two distinct first-hop fragments $i,j\in E_{b}$ and, for each context (third token) $c\in{\mathcal{X}}$, set
+Appendix E Partial Computation Observation Drives the Alignment of Functional Equivalence Representation and Vocabulary Space
+In this section, we investigate how exposure to partial computations affects the interpretability of intermediate state representations through vocabulary space alignment.
+We compare two training conditions on a modified 2-Hop task with $\left|{\mathcal{X}}\right|=50$ and $N=10\text{k}$, after 40k epochs of training:
+  1. 1.
+    Standard Training: $f_{1}\neq f_{2}$, model only sees complete two-hop examples $(x_{1},x_{2},x_{3})\mapsto t$.
+  2. 2.
+    With Partial Computation: $f_{1}=f_{2}$, model additionally sees all possible partial computations $(x_{1},x_{2})\mapsto b$ where $b=f_{1}(x_{1},x_{2})$ (2,500 partial examples, not counted toward the $N=10\text{k}$ two-hop training data).
+To assess interpretability, we measure the Mean Reciprocal Rank (MRR) of intermediate state representations when projected to vocabulary space using the unembedding matrix.
+The low MRR indicates that the model’s internal representation of intermediate state $b$ aligns with the corresponding vocabulary token.
+Fig.˜15 shows a striking contrast between the two conditions.
+Under standard training, the MRR score remains very high throughout training, indicating that intermediate representations are not aligned with vocabulary space despite the model successfully learning the compositional task.
+However, when partial computations are included, the MRR score becomes very high, demonstrating clear vocabulary alignment.
+Figure 15: MRR scores for intermediate state representations projected to vocabulary space.
+Left: Standard training ($f_{1}\neq f_{2}$, no partial computation) shows very high MRR regardless of position and layer.
+Right: Training with partial computation ($f_{1}=f_{2}$, with partial examples) shows MRR of 0 in layers 3 to 8 at position $x_{2}$, indicating strong vocabulary alignment.
+This experiment suggests that logit lens interpretability is orthogonal to functional equivalence representation formation.
+A model can develop functionally correct intermediate representations that enable compositional generalization while remaining completely uninterpretable through standard vocabulary projection techniques.
+Interpretability via logit lens requires explicit vocabulary anchoring through exposure to partial computations that map intermediate states to vocabulary tokens.
+This finding has important implications for mechanistic interpretability research: the absence of interpretable representations through the logit lens does not indicate the absence of structured internal computation.
+Furthermore, it suggests that interpretability techniques may need to account for how training data shapes the alignment between internal representations and vocabulary space, rather than assuming such alignment emerges naturally from task performance.
+Appendix F Proofs: Asymptotically Perfect Coverage in 2-Hop Task
+We establish a scaling-law analysis for the 2-Hop task (as illustrated in Fig.˜2(a)) within our formal framework for understanding pattern matching (Sec.˜3).
+While a simplified statement (Theorem˜6.1) is already presented in the main text, we provide the complete results of our analyses here in great detail, along with their rigorous and self-contained proofs.
+This appendix is long; we have organized it into the following structure.
+  • In Sec.˜F.1, we first review and (re)define the problem setting for 2-Hop task.
+Since the functional ($k$-)equivalence at indices 1 and 2 matters for the 2-Hop task, we argue that it is enough for us to focus on $I=\{1,2\}$, thereby simplifying some key definitions (e.g., substitution graph, $k$-coverage) accordingly.
+  • In Sec.˜F.2, we state and prove our main theorems that characterize the number of training data sufficient for a learner to achieve a perfect in-domain generalization with high probability.
+There are two main theorems: one is for $k\geq 2$ (Theorem˜F.7) and another is for $k=1$ (Theorem˜F.8).
+As a consequence, we further assume a specific regime of set cardinalities—namely, where all token sets have $\Theta(n)$ elements—and provide Corollary˜F.9, combining both cases for $k\geq 2$ and $k=1$.
+    • We also claim that our results are nearly tight, especially for $k\geq 2$.
+It is done by constructing a worst-case subclass of 2-Hop problems in which, given a dataset slightly smaller than our upper bound, the learner fails in achieving a perfect ID generalization with high probability (Theorems˜F.16 and F.17).
+    Theorem˜6.1 in the main text is indeed a simplified combination of Corollaries˜F.9 and F.17.
+  • In Sec.˜F.3, we provide (almost) all postponed proofs of lemmas used in Sec.˜F.2.
+We first prove that the connectivity of all $b$-evidence graphs (Def.˜F.5), which are the intermediate-state-specific induced subgraphs of the substitution graph, is a sufficient condition for the perfect in-domain generalization we want to show: see Lemma˜F.10.
+    • Sec.˜F.3.1 is dedicated to explaining the renowned Poissonization technique, rigorously proving one of our main lemmas, Lemma˜F.12.
+Under a usual fixed-sized dataset sampling scheme, the edge connections of the substitution graph are not guaranteed to be independent.
+Nevertheless, the Poissonization technique enables us to disentangle these dependencies, thereby simplifying our analysis.
+    • Sec.˜F.3.2 reviews the binomial random $k$-intersection graphs, a topic in the random graph theory literature.
+We bring some results on the threshold functions for their connectivity (without bringing their involved proofs).
+It is crucial because we show that the $b$-evidence graphs can be regarded as binomial random $k$-intersection graphs under the Poissonization.
+F.1 Preliminaries: Definitions, Assumptions, and Facts
+Ground-truth mapping.
+Consider a ground-truth mapping $f:{\mathbb{X}}={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\times{\mathcal{X}}_{3}\to Y$ defined as $f(x_{1},x_{2},x_{3})=f_{2}\left(f_{1}(x_{1},x_{2}),x_{3}\right)$, a two-hop composition of primitive functions $f_{1}:{\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\to B$ and $f_{2}:B\times{\mathcal{X}}_{3}\to Y$.
+Here, the set $B:=f_{1}({\mathcal{X}}_{1}\times{\mathcal{X}}_{2})$ is a collection of intermediate states, implicitly assuming surjectivity of $f_{1}$ without loss of generality.
+Train dataset.
+We take a train dataset $D$ by independently sampling input sequences $N$ times from the uniform distribution over ${\mathbb{X}}={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\times{\mathcal{X}}_{3}$.
+For the sake of simplicity, we allow duplicates/replacements of train samples in $D$, thereby considering $D$ as a multiset.
+We find it helpful in our analysis to define a random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ that has a one-to-one correspondence with a dataset constructed by sampling with replacement as described below.
+Proposition F.1.
+Given a train dataset $D$ and for each ${\bm{x}}\in{\mathbb{X}}$, define $Z_{\bm{x}}$ as the sampled count of $\bm{x}$} (i.e., the number of identical duplicates equal to $\bm{x}$}) in $D$.
+Then, if $D$ is generated by with-replacement sampling $N$ times from $\mathbb{X}$}, the random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ follows a multinomial distribution with parameters $n=N$ and $p_{\bm{x}}=\frac{1}{\left|{\mathbb{X}}\right|}$ ($\forall{\bm{x}}\in{\mathbb{X}}$).
+That is, for all collections of integers $z_{\bm{x}}\geq 0$ ($\forall{\bm{x}}\in{\mathbb{X}}$) such that $\sum_{{\bm{x}}\in{\mathbb{X}}}z_{\bm{x}}=N$,
+${\mathbb{P}}(Z_{\bm{x}}=z_{\bm{x}}\;\forall{\bm{x}}\in{\mathbb{X}})=\frac{N!}{|{\mathbb{X}}|^{N}\cdot\prod_{{\bm{x}}\in{\mathbb{X}}}(z_{\bm{x}}!)}.$
+The one-to-one correspondence between $D$ and $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ is straightforward: their realizations uniquely determine each other (up to the order of sampling).
+A learner making predictions using functional $k$-equivalence.
+Let us consider an artificial learner that is perfectly trained on $D$ (i.e., always predicts correctly on $D$).
+We suppose that the learner confidently predicts the output for the elements in $k$-coverage of $D$ (defined with functional $k$-equivalence), while producing a random (or unconfident) prediction outside of it.
+However, the confident prediction is not necessarily correct in general.
+Luckily, since we suppose a clear two-hop compositional structure of the task, the confident predictions based on functional equivalences at indices $I=\{1,2\}$ are guaranteed to be correct.
+Hence, to study the scope of pattern-matching-based in-domain (ID) generalization as confident and correct predictions on unseen inputs, we restrict the scope of our analysis to the index set $\{1,2\}$ for our analysis in this appendix.
+To formalize how far the learner can generalize (Def.˜F.4) and how far we hope the learner to generalize (Def.˜F.2), let us (re)define the key terms for our analysis of the two-hop task, based on the restriction with respect to the index set $I=\{1,2\}$.
+For brevity, let us write $x_{12}:=(x_{1},x_{2})\in{\mathcal{X}}_{12}:={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}$.
+Definition F.2 (In-domain closure, in terms of $I=\{1,2\}$).
+For a train dataset $D$, its in-domain closure $\overline{D}$ is defined as
+$\overline{D}=\left\{(x_{12},x_{3})\!\in\!{\mathbb{X}}:\exists(x^{\prime}_{12},x^{\prime}_{3})\!\in\!{\mathbb{X}}\text{ such that }f_{1}(x_{12})\!=\!f_{1}(x^{\prime}_{12}),(x^{\prime}_{12},x_{3})\!\in\!D,\text{and }(x_{12},x^{\prime}_{3})\!\in\!D\right\}.$
+Also, its element $(x_{12},x_{3})\in\overline{D}$ is said to be an in-domain data.
+According to the definition above, the in-domain closure $\overline{D}$ contains all input sequences $(x_{12},x_{3})$ such that (i) its subsequences $x_{12}$(${\bm{x}}_{I}$) and $x_{3}$(${\bm{x}}_{I^{\mathsf{c}}}$) has already been observed in the train dataset $D$, and (ii) its output can be inferred using functional equivalence at indices $I=\{1,2\}$, namely, $f(x_{12},x_{3})=f(x^{\prime}_{12},x_{3})$ for a training data $(x^{\prime}_{12},x_{3})\in D$.
+Hence, the set $\overline{D}$ is the largest set of input sequences whose output can be correctly predicted using the exact task structure (e.g., the functional equivalence at $\{1,2\}$).
+We also remark that it is obvious to see that $D\subset\overline{D}$.
+We wish the learner would acquire such prediction capabilities from a training dataset.
+Nevertheless, the task structure is not directly accessible to our learner since it can only identify the necessary functional relationships through functional $k$-equivalences (Def.˜3.1).
+On the other hand, the $k$-coverage quantifies how far the assumed learner can actually make correct predictions using the observed functional $k$-equivalences from the training dataset $D$.
+The definition (Def.˜F.4) resembles that of the connectivity of a graph with a vertex set ${\mathcal{X}}_{12}$.
+Recall that in our main text, we defined a substitution graph with the entire input space $\mathbb{X}$} as its vertex set since we aim to check all functional equivalences across all index sets.
+However, since we are particularly interested in the functional equivalences at $I=\{1,2\}$, we can simplify its definition as below.
+Definition F.3 (Substitution graph, in terms of $I=\{1,2\}$).
+For a train dataset $D$ and a positive integer $k$, we define a substitution graph (in terms of $I=\{1,2\}$) as ${\mathcal{G}}_{\bullet}^{(D,k)}=(V_{\bullet},E_{\bullet}^{(D,k)})$.
+Here, the vertex set is $V_{\bullet}:={\mathcal{X}}_{12}$ and the edge set is
+$E_{\bullet}^{(D,k)}:=\left\{\left\{x_{12},x^{\prime}_{12}\right\}\subset V_{\bullet}:x_{12}\neq x^{\prime}_{12}\text{ and }\left|S_{f}(x_{12},x^{\prime}_{12}\mid D)\right|\geq k\right\},$
+where the set $S_{f}(x_{12},x^{\prime}_{12}\mid D)$ is defined as
+$S_{f}(x_{12},x^{\prime}_{12}\mid D):=\left\{\bar{x}_{3}\in{\mathcal{X}}_{3}:f(x_{12},\bar{x}_{3})=f(x^{\prime}_{12},\bar{x}_{3}),(x_{12},\bar{x}_{3})\in D,(x^{\prime}_{12},\bar{x}_{3})\in D\right\}.$
+In other words, in the substitution graph ${\mathcal{G}}_{\bullet}^{(D,k)}$, two vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if they have at least $k$ distinct co-occerences in $D$.
+With this definition, we define the $k$-coverage as follows.
+Definition F.4 ($k$-coverage, in terms of $I=\{1,2\}$).
+For a train dataset $D$ and a positive integer $k$, the $k$-coverage of $D$ (in terms of $I=\{1,2\}$) is a subset of $\mathbb{X}$} defined as
+$\mathrm{Cover}_{k}(D)$
 […]
-Thus $Z_{ij}(c)=1$ exactly when the single context $c$ supplies an evidence pair for the functional equivalence of $i$ and $j$.
-Single-context success probability
-From (E.1),
-$p_{1}:=\Pr[Z_{ij}(c)=1]=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)},\qquad(|{\mathcal{X}}|\to\infty).$
-Negatively correlated counts and an i.i.d. surrogate
-Because all $N$ draws come from a single multinomial $\bigl{(}N,1/|{\mathcal{X}}|^{3}\bigr{)}$, the indicators $\{Z_{ij}(c)\}_{c\in{\mathcal{X}}}$ are negatively correlated: drawing many triples with one context leaves fewer draws for the others.
-Negative correlation decreases the probability that several $Z_{ij}(c)$ equal 1 simultaneously.
-Hence the tail probability for the true count
-$Y_{ij}:=\sum_{c\in{\mathcal{X}}}Z_{ij}(c)$
-is upper-bounded by the tail of an i.i.d. binomial variable with the same single-trial success probability $p_{1}$. Concretely, define
-$Y_{ij}^{\star}\sim\mathrm{Binom}\bigl{(}|{\mathcal{X}}|,p_{1}\bigr{)},\qquad p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)}.$
-Then for every real $t$
-$\Pr\bigl{[}Y_{ij}\geq t\bigr{]}\leq\Pr\bigl{[}Y_{ij}^{\star}\geq t\bigr{]}.$
-Using $Y_{ij}^{\star}$ therefore overestimates the chance of obtaining $k$ or more distinct evidence contexts, which is conservative for our goal of deriving a lower bound on the required dataset size $N$.
-Since $Y_{ij}^{\star}$ is binomial with mean
-$\mu:=\mathbb{E}[Y_{ij}^{\star}]=|{\mathcal{X}}|\,p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{5}}\bigl{(}1+o(1)\bigr{)},$
-we may work henceforth with $Y_{ij}^{\star}$ alone; the resulting bounds apply verbatim to the original $Y_{ij}$.
-Poisson tail via Le Cam
+$\!=\!\left\{\!(u_{0},x_{3})\!\in\!{\mathbb{X}}:\exists\ell\!\geq\!0,\,\exists u_{0},u_{1},\cdots,u_{\ell}\!\in\!{\mathcal{X}}_{\!12}\text{ such that}\left\{u_{i\!-\!1},u_{i}\right\}\!\in\!E_{\bullet}^{(D,k)}\text{ ($\forall i\!\in\![\ell]$) \& }(u_{\ell},\!x_{3})\!\in\!D\!\right\}\!,$
+where $\ell]=\{1,2,3,\cdots,\ell\}$. Also, an element $(x_{12},x_{3})\in\mathrm{Cover}_{k}(D)$ is said to be covered.
+Again, we can easily verify that $D\subset\mathrm{Cover}_{k}(D)$ by choosing $\ell=0$ in the definition above.
+However, the inclusion relation between the in-domain coverage ($\overline{D}$) and the $k$-coverage ($\mathrm{Cover}_{k}(D)$) cannot be explicitly determined in general.
+From the next sub-section, we aim to analyze when (or, with how large $D$) we do have a relation $\overline{D}\subset\mathrm{Cover}_{k}(D)$”, which we call perfect coverage of in-domain data, with high probability (with respect to the randomness of $D$).
+We additionally introduce some useful symbols for our later proofs. We denote by $B_{D}$ the set of all intermediate states observed in $D$: i.e.,
+$B_{D}:=\left\{b\in B:\exists(x_{12},x_{3})\in D\text{ such that }f_{1}(x_{12})=b\right\}.$
+Moreover, we introduce the $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$, defined as a subgraph of ${\mathcal{G}}_{\bullet}^{(D,k)}$ induced by the set $V_{b}=f_{1}^{-1}(\{b\}):=\left\{x_{12}\in{\mathcal{X}}_{12}:f_{1}(x_{12})=b\right\}\subset V_{\bullet}$ of vertices sharing the same intermediate state $b\in B$.
+Definition F.5 (Evidence graphs).
+For a train dataset $D$, a positive integer $k$, and an intermediate state $b\in B$, we define a $b$-evidence graph (in terms of $I=\{1,2\}$) as ${\mathcal{G}}_{b}^{(D,k)}=(V_{b},E_{b}^{(D,k)})$.
+Here, the vertex set is $V_{b}:=f_{1}^{-1}(\{b\})$ and the edge set is
+$E_{b}^{(D,k)}:=\left\{\left\{x_{12},x^{\prime}_{12}\right\}\subset V_{b}:x_{12}\neq x^{\prime}_{12}\text{ and }\left|S(x_{12},x^{\prime}_{12}\mid D)\right|\geq k\right\},$ (5)
+where the set $S(x_{12},x^{\prime}_{12}\mid D)$ is defined as
+$S(x_{12},x^{\prime}_{12}\mid D):=\left\{\bar{x}_{3}\in{\mathcal{X}}_{3}:(x_{12},\bar{x}_{3})\in D,(x^{\prime}_{12},\bar{x}_{3})\in D\right\}.$ (6)
+Lastly, we introduce an assumption that will be assumed only when we claim the tightness of our sample complexity upper bound results.
+Assumption F.6.
+The learner uses the parameter $k\geq k_{\star}+1$, where $k_{\star}\geq 0$ is defined as
+$k_{\star}=\max_{(b,b^{\prime})\in B^{2}}\left|H(b,b^{\prime})\right|\qquad\text{subject to}\quad b\neq b^{\prime},$
+and $H(b,b^{\prime})=\left\{x_{3}\in{\mathcal{X}}_{3}:f_{2}(b,x_{3})=f_{2}(b^{\prime},x_{3})\right\}$.
+According to the assumption above, $f_{2}(b,x_{3})=f_{2}(b^{\prime},x_{3})$ can hold for more than $k_{\star}$ elements in ${\mathcal{X}}_{3}$ only when $b=b^{\prime}$.
+It is quite a natural assumption for general 2-Hop tasks since it naturally eliminates pathological examples that can barely be regarded as 2-Hop tasks (e.g., a constant function $f$ on $\mathbb{X}$}).
+The assumption above additionally assumes that our learner pessimistically recognizes the functional equivalences between $\{1,2\}$-subsequences, by using $k\geq k_{\star}+1$.
+If $k=k_{\star}+1$, the learner is using the precise value of $k$ to recognize functionally equivalent subsequences sharing the same intermediate state; a higher value of $k$ indicates a more pessimistic learner.
+Because of this pessimism, once the learner recognizes that $x_{12}$ and $x^{\prime}_{12}$ are functionally $k$-equivalent (i.e., $\left\{x_{12},x^{\prime}_{12}\right\}\in E^{(D,k)}_{\bullet}$), it is guaranteed that $x_{12}$ and $x^{\prime}_{12}$ share the same intermediate state (i.e., $f_{1}(x_{12})=f_{1}(x^{\prime}_{12})$).
+More importantly, it implies that there must be no edges (in $E^{(D,k)}_{\bullet}$) connecting two graphs ${\mathcal{G}}^{(D,k)}_{b}$ and ${\mathcal{G}}^{(D,k)}_{b^{\prime}}$ for distinct $b,b^{\prime}\in B$; otherwise, there must be a pair of subsequences $x_{12}\in V_{b}$ and $x^{\prime}_{12}\in V_{b^{\prime}}$ that are functionally $k$-equivalent, contradicting the assumption.
+Even though Assumption˜F.6 appears to pertain to the learning rule of the learner, it can also be interpreted as an assumption about the ground-truth mapping, which necessitates $k_{\star}\leq k-1$ for a given $k\geq 1$.
+F.2 Main Theorem: Sample Complexity Bound for Perfect Coverage With High Probability
+The ultimate goal of this appendix is to characterize a sufficient number of training data $|D|$ (i.e., sample complexity upper bound) so that the $k$-coverage includes all in-domain data, i.e., $\overline{D}\subset\mathrm{Cover}_{k}(D)$, with high probability.
+Here, we will state and prove our main theorems, Theorem˜F.7 (for $k\geq 2$) and Theorem˜F.8 (for $k=1$), which are followed by a simplified combination of them (Corollary˜F.9).
+Soon after, we will also claim the near-tightness of these theorems by showing that, for $k\geq 2$, the training datasets slightly smaller than the obtained sample complexity are sufficient to guarantee $\overline{D}\not\subset\mathrm{Cover}_{k}(D)$ with high probability (Theorems˜F.16 and F.17).
+Although we present the proofs of the theorems here, the complete proofs of most lemmas used in the main proofs are postponed to the later part of this appendix (Sec.˜F.3).
+Here are the statements of our main theorems for the sample complexity upper bounds, which will be proved soon after.
+Theorem F.7 (Sample Complexity Upper Bound, $k\geq 2$, $\downarrow$).
+Under the two-hop task setup described in Sec.˜F.1, suppose that $k\geq 2$ is a universal constant.
+Let $\hat{n}:=\min_{b\in B_{D}}\left|V_{b}\right|$ and $\check{n}:=\max_{b\in B_{D}}\left|V_{b}\right|$, and assume that $\left|{\mathcal{X}}_{3}\right|=\Omega(\check{n})$.
+Fix any $\delta>0$.
+Then, there exists $N_{\delta}>0$ satisfying $N_{\delta}\leq O\left(\max\left\{\left(\frac{k!\,(\ln\hat{n})}{\hat{n}}\right)^{\!\!\frac{1}{2k}}\!\cdot\left|{\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\right|\!\sqrt{\left|{\mathcal{X}}_{3}\right|},\;\ln\left(\frac{1}{\delta}\right)\right\}\right)$ such that, for any large enough $\hat{n}$ and any integer $N>N_{\delta}$, we have ${\mathbb{P}}_{N}\left(\overline{D}\subset\mathrm{Cover}_{k}(D)\right)\geq 1-\delta.$ Here, ${\mathbb{P}}_{N}$ is the probability measure for the train dataset $D$ constructed by uniformly randomly sampling its elements NN times, allowing duplicates, independently from ${\mathbb{X}}={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\times{\mathcal{X}}_{3}$.
+Theorem F.8 (Sample Complexity Upper Bound, $k=1$, $\downarrow$).
 […]
-$\Pr[Y_{ij}\geq k]\leq\Pr[Y_{ij}^{\star}\geq k]=\sum_{r=k}^{\infty}\frac{e^{-\mu}\mu^{r}}{r!}\leq\frac{\mu^{k}}{k!}\bigl{(}1+o(1)\bigr{)}.$
-This upper bound $p_{k}:=\mu^{k}/k!(1+o(1))$ will be the edge probability used in the connectivity threshold of Step 3.
-E.3 Step 3: connectivity inside each equivalence class
-With the edge probability $p_{k}$ from Step 2, we now analyze when the $k$-evidence graphs become connected.
-Recall that under our learner assumption, perfect generalization requires every equivalence class to form a connected component in the $k$-evidence graph.
-We model this as a random graph connectivity problem.
-Random graph construction
-Fix one balanced class $E_{b}$ of size $n:=|{\mathcal{X}}|$ (Assumption (A1)).
-Create a graph $G_{b}$ whose vertices are the first-hop pairs in $E_{b}$, and place an undirected edge $\{i,j\}$ exactly when the pair $(i,j)$ has been observed in at least $k$ distinct shared contexts.
-By Step E.2, each potential edge appears with probability
-$p_{k}=\Pr\bigl{[}Y_{ij}\geq k\bigr{]}=\frac{\mu^{k}}{k!}\,,\qquad\mu=\frac{N^{2}}{|{\mathcal{X}}|^{5}}.$
-Why $G_{b}$ can be approximated as Erdős–Rényi
-As mentioned earlier, the dependence among distinct edges in $G_{b}$ arises from the constraint that the total sample size is $N$, which induces a negative correlation.
-Such negative dependence reduces the likelihood of simultaneously creating many edges.
-Consequently, viewing $G_{b}$ as an independent Erdős–Rényi graph $G(n,p_{k})$ provides a conservative model, and any threshold we derive for connectivity under independence remains valid (or becomes easier to satisfy).
-Classwise connectivity threshold
-For Erdős–Rényi graphs, the classical result of Erdős–Rényi [97] states
-$\Pr\bigl{[}G(n,p)\text{ is connected}\bigr{]}\xrightarrow[n\to\infty]{}1\quad\Longleftrightarrow\quad p\geq\frac{\log n+\omega(1)}{n}.$
-Setting $n=|{\mathcal{X}}|$ and $p=p_{k}$ yields the requirement
-$p_{k}=\frac{\mu^{k}}{k!}\geq\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|}(1+o(1)).$
-Substituting $\mu=N^{2}/|{\mathcal{X}}|^{5}$ and rearranging gives
-$\frac{N^{2k}}{|{\mathcal{X}}|^{5k}\,k!}\gtrsim\frac{\log|{\mathcal{X}}|}{|{\mathcal{X}}|},\qquad\Longrightarrow\qquad N\gtrsim|{\mathcal{X}}|^{\frac{5k-1}{2k}}\,\bigl{(}k!\log|{\mathcal{X}}|\bigr{)}^{1/(2k)}.$
-Because $(k!\log|{\mathcal{X}}|)^{1/(2k)}=(\log|{\mathcal{X}}|)^{O(1/k)}$ grows only poly-logarithmically, we hide it inside a $\tilde{\Omega}(\,\cdot\,)$:
+In particular, these two theorems can be easily combined (hence the detailed proof is skipped) in the regime of almost balanced cardinalities as $\Theta(n)$, which is described in the following corollary.
+Corollary F.9 (Power-Law Sample Complexity Upper Bound, Simple).
+Under the two-hop task setup described in Sec.˜F.1, Assume that ${\mathcal{X}}_{1}$, ${\mathcal{X}}_{2}$, ${\mathcal{X}}_{3}$, and $V_{b}$ ($\forall b\in B$) are all the sets of size $\Theta(n)$.
+Fix any $k\geq 1$ and any $\delta>0$ Then, there exists $N_{\delta}(n)>0$ satisfying $N_{\delta}(n)\leq O\left(\max\left\{n^{2.5-\frac{0.5}{k}}\cdot(k!\cdot\ln n)^{\gamma},\;\ln\frac{2}{\delta}\right\}\right),\qquad\text{with }\gamma=\begin{dcases}1,&\text{for }k=1;\\ \frac{1}{2k},&\text{for }k\geq 2,\end{dcases}$ such that, for any large enough $n$ and any $N>N_{\delta}(n)$, the learner with a uniformly randomly sampled training dataset $D$ (with replacements) achieves a perfect coverage of in-domain data, i.e., $\overline{D}\subset\mathrm{Cover}_{k}(D)$, with probability at least $1-\delta$.
+Now, we prove our main theorem for $k\geq 2$ (Theorem˜F.7).
+After that, we will present the proof of the other case ($k=1$, Theorem˜F.8), which follows a similar plot except for a few steps at the end.
+Proof of Theorem˜F.7.
+We begin the proof with the following observation: a sufficient condition for a perfect coverage is that every $b$-evidence graph (Def.˜F.5) is a connected graph.
+Lemma F.10 ($\downarrow$).
+If all ${\mathcal{G}}_{b}^{(D,k)}$ are connected graphs ($\forall b\in B_{D}$), then we have $\overline{D}\subset\mathrm{Cover}_{k}(D)$.
+Refer to Sec.˜F.3 for the proof. Thanks to the lemma above, it suffices to show the following with sufficiently large $N=\left|D\right|$:
 […]
-Since every class must be connected to achieve the data–sufficiency criterion in the problem setting, we conclude
+To this end, consider the multinomial random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ corresponding to a randomly sampled training dataset $D$, as defined in Proposition˜F.1.
+Define a property $\mathcal{C}$} as
+$(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}\in{\mathcal{C}}\iff{\mathcal{G}}_{b}^{(D,k)}\text{ is connected }\forall b\in B_{D}.$ (7)
+Then, we want to compute an upper bound of a probability ${\mathbb{P}}_{N}\left((Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\notin{\mathcal{C}}\right)$.
+Observe that the property $\mathcal{C}$} is a non-decreasing property (Def.˜F.11): if a vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ satisfies $\mathcal{C}$}, then it still satisfies $\mathcal{C}$} after increasing in its entries by non-negative amounts.
+In other words, the negation of $\mathcal{C}$} (“not satisfying $\mathcal{C}$}”) is a non-increasing property.
+This is because adding a data point to the training dataset $D$ does not remove any edge from all evidence graphs, thereby preserving the connectivity.
+In general, we define the monotone property of vectors as below.
+Definition F.11 (Monotone property).
+We refer to a property $\mathcal{A}$} of $m$-dimensional vectors as a non-decreasing property when, for any vector $(\Delta_{1},\cdots,\Delta_{m})$ with non-negative entries $\Delta_{i}\geq 0$ ($\forall i=1,\cdots,m$),
+$(z_{1},\cdots,z_{m})\in{\mathcal{A}}\implies(z_{1}+\Delta_{1},\cdots,z_{m}+\Delta_{m})\in{\mathcal{A}}.$ (Non-Decreasing Property)
+On the other hand, we define non-increasing property $\mathcal{A}$} as what satisfies
+$(z_{1},\cdots,z_{m})\in{\mathcal{A}}\implies(z_{1}-\Delta_{1},\cdots,z_{m}-\Delta_{m})\in{\mathcal{A}}.$ (Non-Increasing Property)
+Taking advantage of the monotonicity, we can apply the Poissonization technique (Sec.˜F.3.1) to obtain an upper bound on the probability ${\mathbb{P}}_{N}\left((Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\notin{\mathcal{C}}\right)$.
+The Poissonization technique we use is summarized as the lemma below, which will be proved in Sec.˜F.3.1:
+Lemma F.12 (De-Poissonizaiton of Monotone Multinomial Events, $\downarrow$).
+Fix any $n\geq 1$. Define
+$P_{n}:={\mathbb{P}}_{(Z_{1},\cdots,Z_{m})\sim\mathrm{Multinomial}(n;\,p_{1},\cdots,p_{m})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right).$
+Let $\mathcal{A}$} be a non-decreasing property (Def.˜F.11) of $m$-dimensional vector. Then, we have an upper bound and a lower bound for $P_{n}$ as follows:
+$P_{n}$ $\leq{\mathbb{P}}_{\mathrm{Po}(n+{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)+\exp\left(\textstyle-\frac{(3-c){\varepsilon}^{2}}{6n}\right),$ $(\forall c\in(0,3),~\forall{\varepsilon}\in(0,cn))$
+$P_{n}$ $\geq{\mathbb{P}}_{\mathrm{Po}(n-{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)-\exp\left(\textstyle-\frac{{\varepsilon}^{2}}{2n}\right).$ $(\forall{\varepsilon}\in(0,n))$
+If $\mathcal{A}$} is a non-increasing property (Def.˜F.11), then we have similar upper and lower bounds for $P_{n}$:
+$P_{n}$ $\leq{\mathbb{P}}_{\mathrm{Po}(n-{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)+\exp\left(\textstyle-\frac{{\varepsilon}^{2}}{2n}\right),$ $(\forall{\varepsilon}\in(0,n))$
+$P_{n}$ $\geq{\mathbb{P}}_{\mathrm{Po}(n+{\varepsilon})}\left(\left(Z_{1},\cdots,Z_{m}\right)\in{\mathcal{A}}\right)-\exp\left(\textstyle-\frac{(3-c){\varepsilon}^{2}}{6n}\right).$ $(\forall c\in(0,3),~\forall{\varepsilon}\in(0,cn))$
+Here, we denote by ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$ the probability measure under $Z_{i}\overset{\text{\emph{indep.}}}{\sim}\mathrm{Poisson}(\lambda p_{i})$ ($\forall i=1,\cdots,m$).
+In particular, considering the disconnectedness ($\exists b\in B_{D}$) as a non-increasing property, we have
+$\ {\mathbb{P}}_{N}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)$
+$\leq{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$
+$\leq\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$ $(\because\text{union bound})$
+for any positive number ${\varepsilon}<N$, where ${\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}$ is the probability measure under i.i.d.
+Poisson random variables $Z_{\bm{x}}\sim\mathrm{Poisson}\left(\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)$ ($\forall{\bm{x}}\in{\mathbb{X}}$).
+Let us take ${\varepsilon}=\sqrt{2N\ln\left(\frac{2}{\delta}\right)}$, which is smaller than $N$ by our choice of $N_{\delta}$ and ensures that
+$\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)=\exp\left(-\ln\left(\frac{2}{\delta}\right)\right)=\frac{\delta}{2}.$
+Now, we claim that each $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ is a binomial random $k$-intersection graph (Sec.˜F.3.2) for every $b\in B_{D}$, under the Poissonization governed by ${\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}$.
+Lemma F.13 (Evidence Graphs are Binomial $k$-Intersection Graphs under Poissonization, $\downarrow$).
+Let any $b\in B_{D}$.
+Consider a vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ associated with a dataset $D$ (i.e., an input sequence $\bm{x}$} is sampled $Z_{\bm{x}}$ times in $D$).
+Let ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$ be a probability measure for $Z_{\bm{x}}\overset{\text{i.i.d.}}{\sim}\mathrm{Poisson}\left(\lambda/\left|{\mathbb{X}}\right|\right)$.
+Then, under ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$, the $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ (Def.˜F.5) is an instance of binomial random $k$-intersection graph ${\mathcal{G}}^{(k)}(n_{b},m,p)$ with parameters
+$n_{b}=\left|V_{b}\right|,\quad m=\left|{\mathcal{X}}_{3}\right|,\quad p=1-\exp\left(-\frac{\lambda}{\left|{\mathbb{X}}\right|}\right).$
+Based on such an observation, we can use the following seminal result about the connectivity of binomial random $k$-intersection graphs with $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$.
+In particular, we will use only the $k\geq 2$” part (which tightly matches our assumption $m=\Omega(n_{b})$ ($\forall b\in B_{D}$)) below:
+Lemma F.14 (zhao2014k, Theorem 2; zhao2017connectivity, Theorem 1 & Remark 1).
+Fix any $k\geq 1$. Suppose that
+$m=\begin{cases}\Omega\left(\min\left\{n\left(\ln n\right)^{5},n^{\rho}\right\}\right),&\text{if }k=1,\text{ for any }\rho>1;\\ \Omega\left(n\right),&\text{if }k\geq 2,\end{cases}$
+and
+$p=\left(\frac{k!\,\left(\ln n+\alpha_{n}\right)}{n}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{m}}$ (8)
+for any sequence $\{\alpha_{n}\}$ which attains a limit $\alpha_{\infty}\in\left[-\infty,+\infty\right]$ as $n\to\infty$. Then,
+$\lim_{n\to\infty}{\mathbb{P}}\left({\mathcal{G}}^{(k)}(n,m,p)\text{ \emph{is connected}}\right)=\lim_{n\to\infty}{\mathbb{P}}\left(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1\right)=\exp\left(-e^{-\alpha_{\infty}}\right),$
+where we compute $\exp\left(-e^{-(-\infty)}\right)=0$ and $\exp\left(-e^{-(+\infty)}\right)=1$.
+Now, we aim to find a sufficient condition for $N$ to have
+$\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)\leq\frac{\delta}{2}.$ (9)
+Let us choose the sequence $\alpha_{n}=\ln n$ in Lemma˜F.14, which guarantees that
+${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1\quad\text{ as }n\to\infty.$
+From the definition of the limit of a sequence, for any choice of $\delta>0$, let us define $n_{0}(\delta)>e>0$ such that
+${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is disconnected})\leq\delta,\qquad\forall n>n_{0}(\delta).$
+Then, choosing all $n_{b}=\left|V_{b}\right|$ to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the bound (9).
+Note that a choice of $p$ larger than the threshold in Eq.˜8 will never change that ${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$.
+Thus, since we choose $\alpha_{n}=\ln n$, it suffices to have
+$p=1-\exp\left(-\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)\geq\left(\frac{k!\,(2\ln n_{b})}{n_{b}}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{\left|{\mathcal{X}}_{3}\right|}}$
+for each $n_{b}$ ($\forall b\in B_{D}$). Since $1-e^{-u}\geq(1-\frac{1}{e})u$ for $0<u<1$ and $\frac{\ln u}{u}$ is a decreasing function for large enough $u>e$, it suffices to have
+$\left(1-\frac{1}{e}\right)\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\geq\left(\frac{k!\,(2\ln\hat{n})}{\hat{n}}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{\left|{\mathcal{X}}_{3}\right|}}$
+for $\hat{n}:=\min_{b\in B_{D}}n_{b}$.
+Plugging in ${\varepsilon}=\sqrt{2N\ln\left(\frac{2}{\delta}\right)}$ and $\left|{\mathbb{X}}\right|=\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\left|{\mathcal{X}}_{3}\right|$, we have
+$N-\sqrt{2N\ln\left(\frac{2}{\delta}\right)}\geq\left(\frac{e}{e-1}\right)\cdot\left(\frac{k!\,(2\ln\hat{n})}{\hat{n}}\right)^{\!\!\frac{1}{2k}}\!\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\sqrt{\left|{\mathcal{X}}_{3}\right|},$
+which is satisfied by
+$N\geq 4\cdot\max\left\{\left(\frac{e}{e-1}\right)\cdot\left(\frac{k!\,(2\ln\hat{n})}{\hat{n}}\right)^{\!\!\frac{1}{2k}}\!\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\!\sqrt{\left|{\mathcal{X}}_{3}\right|},\;\ln\left(\frac{2}{\delta}\right)\right\}.$
+In conclusion, for large enough $\hat{n}>\max\left\{n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right),e\right\}$, provided that the training dataset size $N$ satisfies the inequality above, we finally have
+$\ {\mathbb{P}}_{N}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)$
+$\leq\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,k)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$
+$\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta.$
+This concludes the proof of the theorem. ∎
+Proof of Theorem˜F.8.
+Even for $k=1$, we follow almost the same proof as that of Theorem˜F.7, except for the last few steps.
+Namely, we again observe that every ${\mathcal{G}}_{b}^{(D,1)}$ is a binomial random $1$-intersection graph ${\mathcal{G}}^{(1)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$ (see the definition of binomial random intersection graph in Sec.˜F.3.2).
+To guarantee the union bound in Eq.˜9, however, we use a different lemma: Lemma˜F.15 (instead of Lemma˜F.14).
+In particular, we will use only the $\rho\leq 1$” part (which tightly matches our assumption $m=O(n_{b})$ ($b\in B_{D}$)) of the lemma below:
+Lemma F.15 (singer1995random, Propositions 3.1–2, Theorem 3.3).
+Let $k=1$. Suppose that $m=n^{\rho}$ for $\rho>0$ and
+$p=\begin{dcases}\frac{\ln n+\alpha_{n}}{m}&\text{for }\rho\leq 1;\\ \sqrt{\frac{\ln n+\alpha_{n}}{mn}}&\text{for }\rho>1,\end{dcases}$
+for any sequence $\{\alpha_{n}\}$ which attains a limit $\alpha_{\infty}\in\left\{-\infty,+\infty\right\}$ as $n\to\infty$. Then,
 […]
+Again, let us choose the sequence $\alpha_{n}=\ln n$ in Lemma˜F.15, which guarantees that ${\mathbb{P}}({\mathcal{G}}^{(1)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$.
+From the definition of the limit of a sequence, for any choice of $\delta>0$, let us define $n_{0}(\delta)>1$ such that
+${\mathbb{P}}({\mathcal{G}}^{(1)}(n,m,p)\text{ is disconnected})\leq\delta,\qquad\forall n>n_{0}(\delta).$
+Then, choosing all $n_{b}=\left|V_{b}\right|$ (and thus $\hat{n}=\min_{b\in B_{D}}n_{b}$) to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the same bound (9).
+Now, by our choice $\alpha_{n}=\ln n$, it suffices to have
+$p=1-\exp\left(-\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)\geq\frac{2\ln n_{b}}{|{\mathcal{X}}_{3}|}$
+for each $n_{b}$ ($\forall b\in B_{D}$). Since $1-e^{-u}\geq(1-\frac{1}{e})u$ for $0<u<1$ and $\ln u$ is an increasing function for $u>1$, it suffices to have
+$\left(1-\frac{1}{e}\right)\frac{N-{\varepsilon}}{\left|{\mathbb{X}}\right|}\geq\frac{2\ln\check{n}}{|{\mathcal{X}}_{3}|}$
+for $\check{n}:=\max_{b\in B_{D}}n_{b}$.
+Plugging in ${\varepsilon}=\sqrt{2N\ln\left(\frac{2}{\delta}\right)}$ and $\left|{\mathbb{X}}\right|=\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\left|{\mathcal{X}}_{3}\right|$, we have
+$N-\sqrt{2N\ln\left(\frac{2}{\delta}\right)}\cdot\geq\left(\frac{e}{e-1}\right)(2\ln\check{n})\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|,$
+which is satisfied by
+$N\geq 4\cdot\max\left\{\left(\frac{e}{e-1}\right)\cdot(2\ln\check{n})\!\cdot\!\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|,\;\ln\left(\frac{2}{\delta}\right)\right\}.$
+In conclusion, for large enough $\check{n}\geq\hat{n}>\max\left\{n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right),1\right\}$, provided that the training dataset size $N$ satisfies the inequality above, we finally have
+$\ {\mathbb{P}}_{N}\left(\exists b\in B_{D}\text{ such that }{\mathcal{G}}_{b}^{(D,1)}\text{ is disconnected}\right)$
+$\leq\sum_{b\in B_{D}}{\mathbb{P}}_{\mathrm{Po}(N-{\varepsilon})}\left({\mathcal{G}}_{b}^{(D,1)}\text{ is disconnected}\right)+\exp\left(-\frac{{\varepsilon}^{2}}{2N}\right)$
+$\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta.$
+This concludes the proof. ∎
+Subsequently, we argue that the sample complexity upper bound obtained in Theorem˜F.7 (for $k\geq 2$) is nearly tight.
+[^7] We often say a complexity upper bound is tight if we can find a worst-case example whose complexity lower bound is almost identical to the upper bound.
+To this end, we show that the learner (we assumed in Sec.˜F.1) cannot avoid an incomplete coverage (i.e., $\overline{D}\not\subset\mathrm{Cover}_{k}(D)$) with high probability, with a dataset slightly smaller than our upper bound for certain instances of the 2-Hop task.
+In particular, it is enough to consider a subclass of the 2-Hop task satisfying a mild condition (Assumption˜F.6) that any two distinct intermediate states in $B$ never share the same output for more than $k_{\star}$(<k<k) tokens in ${\mathcal{X}}_{3}$.
+The formal statement of the tightness result is shown below:
+Theorem F.16 (Near-Tightness of the Sample Complexity Upper Bound for $k\geq 2$, $\downarrow$).
 […]
-Appendix F Additional results for power-law scaling analysis
-F.1 Measurement protocol for $N_{\mathrm{req}}$
+A caveat of Theorem˜F.16 is that it makes sense only when the range of $N$ is nonempty.
+Luckily, it is easy to verify its validity in the regime of balanced cardinalities, similar to what we assumed in Corollary˜F.9.
+As a result, we establish the tightness (for $k\geq 2$, up to a constant factor) of our sample complexity upper bound in Corollary˜F.9, among the data sizes $N\gtrsim n^{2}\ln n$ (Corollary˜F.17).
+Note that, as shown in the middle of the proof of Theorem˜F.16, $O(n^{2}\ln n)$ is indeed the data size which is sufficient to guarantee that the learner observes all possible pairs in ${\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}$ and $B\!\times\!{\mathcal{X}}_{3}$, ensuring that all possible input sequences in $\mathbb{X}$} are in-domain, with high probability.
+Corollary F.17 (Tightness of Sample Complexity Upper Bound for $k\geq 2$, Simple).
+Under the two-hop task setup described in Sec.˜F.1, suppose that $\left|{\mathcal{X}}_{1}\right|=\left|{\mathcal{X}}_{2}\right|=\left|{\mathcal{X}}_{3}\right|=\left|B\right|=\left|V_{b}\right|=n$ ($\forall b\in B$).
+Also, assume that the ground-truth mapping $f$ satisfies Assumption˜F.6.
+Fix any $k\geq 2$ and any $\delta>0$.
+Then, there exists $\widetilde{N}_{\delta}(n)>0$ satisfying $\widetilde{N}_{\delta}(n)\geq n^{2.5-\frac{0.5}{k}}\cdot((k-1)!\cdot\ln n)^{\frac{1}{2k}}$ such that, for any large enough $n$ (e.g., such that $kn>256e(\ln n)^{3}$ and $\sqrt{n}(\ln n)^{\frac{1}{k}}>4(\ln\frac{4}{\delta})^{2}$) and any integer NN satisfying that $n^{2}\left(2\ln n+\ln\frac{4}{\delta}\right)\leq N<\widetilde{N}_{\delta}(n),$ the learner with a uniformly randomly sampled training dataset $D$ (with replacements) cannot achieve a perfect coverage of in-domain data, i.e., $\overline{D}\not\subset\mathrm{Cover}_{k}(D)$, with probability at least $1-\delta$.
+From now on, we prove Theorem˜F.16, the tightness result of the sample complexity upper bound obtained in Theorem˜F.7.
+Proof of Theorem˜F.16.
+We aim to prove that, even when the dataset $D$ is large enough to ensure that $\overline{D}={\mathbb{X}}$, the incomplete coverage may happen with high probability for a certain range of data size $N$.
+To this end, let us first define two sets $D_{12}$ and $D_{B3}$ as follows:
+$D_{12}$ $:=\left\{x_{12}\in{\mathcal{X}}_{12}:\exists x_{3}\in{\mathcal{X}}_{3}\text{ such that }(x_{12},x_{3})\in D\right\};$
+$D_{B3}$ $:=\left\{(b,x_{3})\in B\!\times\!{\mathcal{X}}_{3}:\exists x_{12}\in V_{b}\text{ such that }(x_{12},x_{3})\in D\right\}.$
+Also, we denote by $\min\deg{\mathcal{G}}:=\min_{v\in V}\deg_{\mathcal{G}}(v)$ the minimum degree among all vertices of a graph ${\mathcal{G}}=(V,E)$.
+With these definitions in place, we now apply the following lemma, which describes a necessary condition for perfect coverage.
+Lemma F.18 ($\downarrow$).
+Assume that $D_{12}={\mathcal{X}}_{12}$ and $D_{B3}=B\times{\mathcal{X}}_{3}$ hold.
+Then, these imply that $\overline{D}={\mathbb{X}}$.
+Furthermore, suppose that Assumption˜F.6 holds for a given $1\leq k\leq\left|{\mathcal{X}}_{3}\right|$.
+Then, $\overline{D}\subset\mathrm{Cover}_{k}(D)$ implies that $\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1$ for all $b\in B$.
+Refer to Sec.˜F.3 for its proof. Now, it suffices to show the inequality below:
+$\ {\mathbb{P}}_{N}\left(\overline{D}\subset\mathrm{Cover}_{k}(D)\right)$
+$\leq{\mathbb{P}}_{N}\left(\overline{D}\subset\mathrm{Cover}_{k}(D)\ \text{ or }\ D_{12}\neq{\mathcal{X}}_{12}\ \text{ or }\ D_{B3}\neq B\times{\mathcal{X}}_{3}\right)$
+[…]
+$\leq{\mathbb{P}}_{N}\left(\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1\ (\forall b\in B)\right)+{\mathbb{P}}_{N}\left(D_{12}\neq{\mathcal{X}}_{12}\right)+{\mathbb{P}}_{N}\left(D_{B3}\neq B\times{\mathcal{X}}_{3}\right)\leq\delta.$ (10)
+To this end, consider the multinomial random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}\sim\mathrm{Multinomial}(N;(p_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}})$ corresponding to a randomly sampled training dataset $D$, as defined in Proposition˜F.1.
+Define a property ${\mathcal{M}}_{1}$ as
+$(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}\in{\mathcal{M}}_{1}\iff\min\deg{\mathcal{G}}_{b}^{(D,k)}\geq 1\;\;(\forall b\in B).$ (11)
+Also, observe that
+$(T_{x_{12}})_{x_{12}\in{\mathcal{X}}_{12}}$ $\sim\mathrm{Multinomial}(N;(q_{x_{12}})_{x_{12}\in{\mathcal{X}}_{12}}),$ $\text{if }\begin{dcases}T_{x_{12}}=\sum_{x_{3}\in X_{3}}Z_{(x_{12},x_{3})},\\ q_{x_{12}}=\sum_{x_{3}\in X_{3}}p_{(x_{12},x_{3})},\end{dcases}$
+$(U_{w})_{w\in B\times{\mathcal{X}}_{3}}$ $\sim\mathrm{Multinomial}(N;(r_{w})_{w\in B\times{\mathcal{X}}_{12}}),$ $\text{if }\begin{dcases}U_{(b,x_{3})}=\sum_{x_{12}\in V_{b}}Z_{(x_{12},x_{3})},\\ r_{(b,x_{3})}=\sum_{x_{12}\in V_{b}}p_{(x_{12},x_{3})},\end{dcases}$
+where we can actually put $p_{(x_{12},x_{3})}=\frac{1}{\left|{\mathbb{X}}\right|}$, $q_{x_{12}}=\frac{\left|{\mathcal{X}}_{3}\right|}{\left|{\mathbb{X}}\right|}=\frac{1}{\left|{\mathcal{X}}_{12}\right|}$, and $r_{(b,x_{3})}=\frac{\left|V_{b}\right|}{\left|{\mathbb{X}}\right|}$ above.
+Using this notation, we know that
+$D_{12}={\mathcal{X}}_{12}$ $\iff T_{x_{12}}\geq 1\;\;(\forall x_{12}\in{\mathcal{X}}_{12});$
+$D_{B3}=B\times{\mathcal{X}}_{3}$ $\iff U_{w}\geq 1\;\;(\forall w\in B\times{\mathcal{X}}_{3}).$
+Hence, to show Eq.˜10, it suffices to prove the following three inequalities:
+${\mathbb{P}}_{(Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\sim\mathrm{Multinomial}(N;(p_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}})}\left((Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\in{\mathcal{M}}_{1}\right)$ $\leq\frac{\delta}{2};$ (12)
+${\mathbb{P}}_{(T_{x_{12}})_{x_{12}\in{\mathcal{X}}_{12}}\sim\mathrm{Multinomial}(N;(q_{x_{12}})_{x_{12}\in{\mathcal{X}}_{12}})}\left(\exists x_{12}\in{\mathcal{X}}_{12}\ \text{ such that }\ T_{x_{12}}=0\right)$ $\leq\frac{\delta}{4};$ (13)
+${\mathbb{P}}_{(U_{w})_{w\in B\times{\mathcal{X}}_{3}}\sim\mathrm{Multinomial}(N;(r_{w})_{w\in B\times{\mathcal{X}}_{12}})}\left(\exists w\in B\times{\mathcal{X}}_{3}\ \text{ such that }\ U_{w}=0\right)$ $\leq\frac{\delta}{4}.$ (14)
+A size $N$ of training dataset $D$ that is sufficient to ensure Eqs.˜13 and 14 can be characterized with the following lemma.
+Lemma F.19 (A Tail Bound for Coupon Collector’s Problem, $\downarrow$).
+Consider a multinomial random vector $(Z_{1},\cdots,Z_{m})\sim\mathrm{Multinomial}(n;p_{1},\cdots,p_{m})$. Then,
+${\mathbb{P}}(\exists i\ \text{ \emph{such that} }\ Z_{i}=0)\leq\sum_{i=1}^{m}(1-p_{i})^{n}.$
+In particular, if $\hat{p}=\min_{1\leq i\leq m}p_{i}$, then for any $\delta>0$,
+$n\geq\frac{1}{\hat{p}}\ln\frac{m}{\delta}\implies{\mathbb{P}}(\exists i\ \text{ \emph{such that} }\ Z_{i}=0)\leq\delta.$
+See Sec.˜F.3 for the proof. Applying the lemma and $\hat{n}=\min_{b\in B}\left|V_{b}\right|$, we can figure out that the data size
+$N\geq\max\left\{\left|{\mathcal{X}}_{12}\right|\ln\left(\frac{4\left|{\mathcal{X}}_{12}\right|}{\delta}\right),\;\frac{\left|{\mathbb{X}}\right|}{\hat{n}}\ln\left(\frac{4\left|B\!\times\!{\mathcal{X}}_{3}\right|}{\delta}\right)\right\}$ (15)
+is enough to ensure both Eqs.˜13 and 14. The remaining task now is to determine a condition for $N$ that guarantees Eq.˜12.
+Observe that ${\mathcal{M}}_{1}$ (defined as Eq.˜11) is a non-decreasing property (Def.˜F.11) because of a similar reason for the monotonicity of the property $\mathcal{C}$} (defined as Eq.˜7; see the proof of Theorem˜F.7).
+Thanks to the monotonicity of ${\mathcal{M}}_{1}$, we can apply the Poissonization technique (Lemma˜F.12) once again.
+That is, taking any fixed $c\in(0,2)$, we have
+$\ {\mathbb{P}}_{N}\left(\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1\;\;(\forall b\in B)\right)$
+$\leq{\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}\left(\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1\;\;(\forall b\in B)\right)+\exp\left(-\frac{(3-c){\varepsilon}^{2}}{6N}\right)$
+$\leq{\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}\left(\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1\right)+\exp\left(-\frac{(3-c){\varepsilon}^{2}}{6N}\right)\qquad(\forall b\in B),$
+for any ${\varepsilon}\in(0,cN)$, where ${\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}$ is the probability measure under i.i.d.
+Poisson random variables $Z_{\bm{x}}\sim\mathrm{Poisson}\left(\frac{N+{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)$ ($\forall{\bm{x}}\in{\mathbb{X}}$).
+Let us take ${\varepsilon}=\sqrt{\frac{6N}{3-c}\ln\left(\frac{4}{\delta}\right)}$, which is smaller than $N$ by Eq.˜15 (since $\left|{\mathcal{X}}_{12}\right|\geq 6>\frac{6}{3-c}$) and ensures that
+$\exp\left(-\frac{(3-c){\varepsilon}^{2}}{6N}\right)=\exp\left(-\ln\left(\frac{4}{\delta}\right)\right)=\frac{\delta}{4}.$
+[…]
+${\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}\left(\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1\right)\leq\frac{\delta}{4}.$ (16)
+Let us choose the sequence $\alpha_{n}=-(1-\frac{1}{k})\ln n$ in Lemma˜F.14, which guarantees that
+${\mathbb{P}}(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1)\to 0\quad\text{as }n\to\infty.$
+From the definition of the limit of a sequence, for any choice of $\delta>0$, let us consider $n_{0}(\delta)>0$ satisfying that
+${\mathbb{P}}\left(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1\right)\leq\delta,\qquad n>n_{0}(\delta).$
+Then, choosing $\hat{n}$ greater than $n_{0}\left(\frac{\delta}{4}\right)$, we yield the bound (16).
+Observe that a choice of $p$ smaller than the threshold in Eq.˜8 will never change the limit ${\mathbb{P}}(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1)\to 0$.
+Thus, by our choice $\alpha_{n}=-(1-\frac{1}{k})\ln n$, it is enough to have
+$p=1-\exp\left(-\frac{N+{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)\leq\left(\frac{(k-1)!\,(\ln\hat{n})}{\hat{n}}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{\left|{\mathcal{X}}_{3}\right|}}.$
+Applying $1-e^{-u}\leq u$ for $u\in{\mathbb{R}}$, it suffices to have
+$\frac{N+{\varepsilon}}{\left|{\mathbb{X}}\right|}\leq\left(\frac{(k-1)!\,(\ln\hat{n})}{\hat{n}}\right)^{\frac{1}{2k}}\cdot\frac{1}{\sqrt{\left|{\mathcal{X}}_{3}\right|}}.$
+Plugging ${\varepsilon}<cN$ and $\left|{\mathbb{X}}\right|=\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\left|{\mathcal{X}}_{3}\right|$ in, we eventually have a sufficient condition
+$N\leq\frac{1}{1+c}\cdot\left(\frac{(k-1)!\,(\ln\hat{n})}{\hat{n}}\right)^{\frac{1}{2k}}\cdot\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\sqrt{\left|{\mathcal{X}}_{3}\right|}.$
+To summarize, for large enough $\hat{n}>n_{0}\left(\frac{\delta}{4}\right)$, provided that the data size $N$ satisfies
+$\max\left\{\left|{\mathcal{X}}_{12}\right|\ln\frac{4\left|{\mathcal{X}}_{12}\right|}{\delta}\!,\frac{\left|{\mathbb{X}}\right|}{\hat{n}}\ln\frac{4\left|B\!\times\!{\mathcal{X}}_{3}\right|}{\delta}\right\}\leq N\leq\frac{1}{1\!+\!c}\!\cdot\left(\!\frac{(k\!-\!1)!(\ln\hat{n})}{\hat{n}}\!\right)^{\!\frac{1}{2k}}\!\cdot\!\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\!\sqrt{\left|{\mathcal{X}}_{3}\right|}$
+and we take ${\varepsilon}=\sqrt{\frac{6N}{3-c}\ln\left(\frac{4}{\delta}\right)}$, we finally have
+$\ {\mathbb{P}}_{N}\left(\overline{D}\subset\mathrm{Cover}_{k}(D)\right)$
+$\leq{\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}\left(\min\deg{\mathcal{G}}^{(D,k)}_{b}\geq 1\right)+\exp\left(-\frac{(3-c){\varepsilon}^{2}}{6N}\right)+{\mathbb{P}}_{N}\left(D_{12}\neq{\mathcal{X}}_{12}\right)+{\mathbb{P}}_{N}\left(D_{B3}\neq B\!\times\!{\mathcal{X}}_{3}\right)$
+$\leq\frac{\delta}{4}+\frac{\delta}{4}+\frac{\delta}{4}+\frac{\delta}{4}=\delta.$
+Since the choice of $c\in(0,2)$ is arbitrary, we obtain the same result (${\mathbb{P}}_{N}(\cdots)\leq\delta$) by letting $c\searrow 0$ and choosing $N$ which satisfies that
+$\max\left\{\left|{\mathcal{X}}_{12}\right|\ln\frac{4\left|{\mathcal{X}}_{12}\right|}{\delta},\;\frac{\left|{\mathbb{X}}\right|}{\hat{n}}\ln\frac{4\left|B\!\times\!{\mathcal{X}}_{3}\right|}{\delta}\right\}\leq N<\left(\!\frac{(k\!-\!1)!(\ln\hat{n})}{\hat{n}}\!\right)^{\!\frac{1}{2k}}\!\cdot\!\left|{\mathcal{X}}_{1}\!\times\!{\mathcal{X}}_{2}\right|\!\sqrt{\left|{\mathcal{X}}_{3}\right|}.$
+Before moving on to the postponed proofs of lemmas, we lastly remark that the same proof (of Theorem˜F.16) can hardly apply to the case of $k=1$ in general.
+This is because, for the sake of simplicity in applying the necessary condition for a perfect coverage (Lemma˜F.18), we first characterize a minimal data size to guarantee $\overline{D}={\mathbb{X}}$ with high probability as Eq.˜15, using the tail bound of the coupon collector’s problem (Lemma˜F.19).
+Unfortunately, it already exceeds the sample complexity upper bound to ensure $\overline{D}\subset\mathrm{Cover}_{1}(D)$ (obtained as Theorem˜F.8) for $k=1$ with high probability, especially in the regime of balanced cardinalities assumed in Corollary˜F.17.
+F.3 Backgrounds, Useful Facts, and Lemmas
+Now, we delve into the detailed backgrounds to provide a comprehensive understanding of the main theorems’ proof, along with the deferred proofs of the lemmas used earlier.
+We begin with the proof of a sufficient condition for a perfect coverage (Lemma˜F.10). For readability, we restate the lemma here.
+See F.10
+Proof of Lemma˜F.10.
+Take any $(x_{12},x_{3})\in\overline{D}$.
+By definition (Def.˜F.2), there exists $x^{\prime}_{12}\in{\mathcal{X}}_{12}$ such that $f_{1}(x_{12})=f_{1}(x^{\prime}_{12})$ and $(x^{\prime}_{12},x_{3})\in D$.
+Let $b=f_{1}(x^{\prime}_{12})\in B_{D}$.
+By assumption, there exists a path in ${\mathcal{G}}_{b}^{(D,k)}$ connecting $x_{12}$ and $x^{\prime}_{12}$, which we write as $u_{0}(=x_{12}),u_{1},\ldots,u_{\ell}(=x^{\prime}_{12})$ for some integer $\ell\geq 1$.
+Then, for each $i\in[\ell]$, we have a set $S(u_{i-1},u_{i}\mid D)$ (defined in Def.˜F.5) of size at least $k$.
+For each $\bar{x}_{3}\in S(u_{i-1},u_{i}\mid D)$, both $(u_{i-1},\bar{x}_{3})$ and $(u_{i},\bar{x}_{3})$ are both in $D$.
+Also,
+$f(u_{i-1},\bar{x}_{3})=f_{2}(b,\bar{x}_{3})=f(u_{i},\bar{x}_{3}),$
+meaning that the set $S_{f}(u_{i-1},u_{i}\mid D)$ (defined in Def.˜F.3) is the same as $S(u_{i-1},u_{i}\mid D)$, a set of size at least $k$, for each $i\in[\ell]$.
+Hence, $(x_{12},x_{3})$ satisfies the definition of $k$-coverage with a path $u_{0},\ldots,u_{\ell}$, i.e., $(x_{12},x_{3})\in\mathrm{Cover}_{k}(D)$.
+∎
+Next, we show a necessary condition for a perfect coverage (Lemma˜F.18). Before proving the lemma, we recall that we have defined two sets:
+$D_{12}$ $:=\left\{x_{12}\in{\mathcal{X}}_{12}:\exists x_{3}\in{\mathcal{X}}_{3}\text{ such that }(x_{12},x_{3})\in D\right\};$
+$D_{B3}$ $:=\left\{(b,x_{3})\in B\!\times\!{\mathcal{X}}_{3}:\exists x_{12}\in V_{b}\text{ such that }(x_{12},x_{3})\in D\right\}.$
+Also, for a graph ${\mathcal{G}}=(V,E)$, we define $\min\deg{\mathcal{G}}:=\min_{v\in V}\deg_{\mathcal{G}}(v)$ as the minimum degree among all vertices of $\mathcal{G}$}.
+See F.18
+Proof of Lemma˜F.18.
+We first show that $D_{12}={\mathcal{X}}_{12}$ and $D_{B3}=B\times{\mathcal{X}}_{3}$ imply that $\overline{D}={\mathbb{X}}$.
+Since we already have $\overline{D}\subset{\mathbb{X}}$ by definition, it suffices to prove ${\mathbb{X}}\subset\overline{D}$.
+Let any $x_{12}\in{\mathcal{X}}_{12}$ and $x_{3}\in{\mathcal{X}}_{3}$, and take $b=f_{1}(x_{12})\in B$.
+Let us define:
+$W_{x_{12}}$ $=\left\{x^{\prime}_{3}\in{\mathcal{X}}_{3}:(x_{12},x^{\prime}_{3})\in D\right\};$
+$\widetilde{W}^{(b)}_{x_{3}}$ $=\left\{x^{\prime}_{12}\in V_{b}:(x^{\prime}_{12},x_{3})\in D\right\}.$
+The conditions $D_{12}={\mathcal{X}}_{12}$ and $D_{B3}=B\times{\mathcal{X}}_{3}$ imply that $W_{x_{12}}\neq\varnothing$ and $\widetilde{W}^{(b)}_{x_{3}}\neq\varnothing$, respectively.
+If we take any $x^{\prime}_{3}\in W_{x_{12}}$ and $x^{\prime}_{12}\in\widetilde{W}^{(b)}_{x_{3}}$, they hold that $(x_{12},x^{\prime}_{3})\in D$, $(x^{\prime}_{12},x_{3})\in D$, and $f_{1}(x_{12})=b=f_{1}(x^{\prime}_{12})$.
+This set of conditions is equivalent to $(x_{12},x_{3})\in\overline{D}$; hence, we have just shown that
+$D_{12}={\mathcal{X}}_{12}\ \text{ and }\ D_{B3}=B\times{\mathcal{X}}_{3}\implies\overline{D}={\mathbb{X}}.$
+Now, we prove the contrapositive of the lemma.
+Suppose that $\min\deg{\mathcal{G}}^{(D,k)}_{b}=0$ for some $b\in B$; we aim to show that $\mathrm{Cover}_{k}(D)\neq{\mathbb{X}}=\overline{D}$, or, ${\mathbb{X}}\setminus\mathrm{Cover}_{k}(D)\neq\varnothing$.
+Take a $b\in B$ such that $\min\deg{\mathcal{G}}^{(D,k)}_{b}=0$, which implies that such a graph ${\mathcal{G}}^{(D,k)}_{b}$ has an isolated vertex $x_{12}\in V_{b}$ (i.e., $\deg_{{\mathcal{G}}^{(D,k)}_{b}}(x_{12})=0$).
+Let us fix such an $x_{12}$.
+Since we assume Assumption˜F.6, there must not be any edge connection between ${\mathcal{G}}^{(D,k)}_{b}$ and ${\mathcal{G}}^{(D,k)}_{b^{\prime}}$ for distinct $b,b^{\prime}\in B$.
+Thus, $x_{12}$ must be an isolated vertex in the whole substitution graph ${\mathcal{G}}^{(D,k)}_{\bullet}$.
+The isolation implies that, for any $x_{3}\in{\mathcal{X}}_{3}$, the input sequence $(x_{12},x_{3})$ cannot be in the $k$-coverage of $D$ unless it is already in $D$.
+Observe that $W_{x^{\prime}_{12}}\neq\varnothing$ for any $x^{\prime}_{12}\in{\mathcal{X}}_{12}$; otherwise, it inevitably holds that $(x^{\prime}_{12},\tilde{x}_{3})\neq\overline{D}$ for every $\tilde{x}_{3}\in{\mathcal{X}}_{3}$, which contradicts to the fact $\overline{D}={\mathbb{X}}$ (under the assumption of the lemma).
+Thus, we can think of the following two cases:
+  • Case 1 ($\varnothing\neq W_{x_{12}}\subsetneq{\mathcal{X}}_{3}$).
+Take any $x_{3}\notin W_{x_{12}}$, which implies $(x_{12},x_{3})\notin D$.
+Thus, we have an element $(x_{12},x_{3})\in{\mathbb{X}}\setminus\mathrm{Cover}_{k}(D)$ because $x_{12}$ is an isolated vertex.
+  • Case 2 ($W_{x_{12}}={\mathcal{X}}_{3}$).
+Since $x_{12}$ is an isolated vertex, it is not adjacent to any other vertex $x^{\prime}_{12}$.
+Fix any such an $x^{\prime}_{12}\in{\mathcal{X}}_{12}\setminus\left\{x_{12}\right\}$.
+By definition of the edges in $E^{(D,k)}_{\bullet}$, we know that $\left|W_{x_{12}}\cap W_{x^{\prime}_{12}}\right|<k$.
+Since $W_{x_{12}}={\mathcal{X}}_{3}$, we also know that $\left|W_{x^{\prime}_{12}}\right|<k\leq\left|{\mathcal{X}}_{3}\right|$.
+This implies that $x^{\prime}_{12}$ cannot be adjacent to any other vertices in ${\mathcal{X}}_{12}$, meaning that $x^{\prime}_{12}$ is an isolated vertex.
+Moreover, since $W_{x^{\prime}_{12}}\subsetneq{\mathcal{X}}_{3}$ ($\because\left|W_{x^{\prime}_{12}}\right|<\left|{\mathcal{X}}_{3}\right|$), we take any $x_{3}\notin W_{x^{\prime}_{12}}$.
+Then, since $(x^{\prime}_{12},x_{3})\notin D$ and $x^{\prime}_{12}$ is isolated, we have an element $(x^{\prime}_{12},x_{3})\in{\mathbb{X}}\setminus\mathrm{Cover}_{k}(D)$.
+In both cases above, we obtain the same result that ${\mathbb{X}}\setminus\mathrm{Cover}_{k}(D)\neq\varnothing$. It concludes the proof of the lemma. ∎
+We also prove the tail probability bound for the coupon collector’s problem (newman1960double; erdHos1961classical) here. We again restate the lemma here for the sake of readability.
+See F.19
+Proof of Lemma˜F.19.
+Observe that $Z_{i}\sim\mathrm{Bin}(n,p_{i})$ for each $i\in[m]$.
+It implies that ${\mathbb{P}}(Z_{i}=0)=(1-p_{i})^{n}$.
+One can yield the same result by directly summing up the multinomial probability masses and applying the multinomial theorem: for instance,
+${\mathbb{P}}(Z_{1}=0)=\sum_{\begin{subarray}{c}z_{2}+\cdots+z_{m}=n\\ z_{2},\cdots,z_{m}\geq 0\end{subarray}}\frac{n!}{z_{2}!\cdots z_{m}!}\cdot p_{2}^{z_{2}}\cdots p_{m}^{z_{m}}=(p_{2}+\cdots+p_{m})^{n}=(1-p_{1})^{n}.$
+Thus, by applying a union bound,
+${\mathbb{P}}(\exists i\ \text{ such that }\ Z_{i}=0)\leq\sum_{i=1}^{m}{\mathbb{P}}(Z_{i}=0)=\sum_{i=1}^{m}(1-p_{i})^{n}\leq m(1-\hat{p})^{n}\leq m\,e^{-n\hat{p}}$,
+where we apply $\hat{p}=\min_{1\leq i\leq m}p_{i}$ and $1+u\leq e^{u}$ ($\forall u\in{\mathbb{R}}$) in the last two inequalities above.
+Solving $m\,e^{-n\hat{p}}\leq\delta$, we conclude that $n\geq\frac{1}{\hat{p}}\ln\frac{m}{\delta}$ implies ${\mathbb{P}}(\exists i\ \text{ such that }\ Z_{i}=0)\leq\delta$.
+∎
+F.3.1 (De-)Poissonization of Monotone Multinomial Events
+Randomization is a probabilistic technique that provides a convenient way to analyze a sequence by treating the sequence index as a random variable/process.
+When the sequence $\{a_{n}\}_{n\geq 0}$ is monotone and bounded, in particular, it provides us upper/lower bounds of the difference between $a_{n}$ and the expectation $\mathbb{E}[a_{N}]$ in terms of a non-negative integral random variable $N$.
+The following lemma provides a simple conversion between a deterministic monotone bounded sequence and the expectation of the randomized sequence, which we will prove later.
+Lemma F.20 (De-Randomization Lemma, General, $\downarrow$).
+Consider a non-decreasing real-valued sequence $\{a_{j}\}_{j\geq 0}$ which lies in a closed interval $[m,M]$, i.e., $m\leq a_{0}\leq a_{1}\leq a_{2}\leq\cdots\leq M$.
+Let $N$ be a non-negative integer-valued random variable with probability mass ${\mathbb{P}}(N=j)=p_{j}$ ($j=0,1,2,\cdots$).
+Then, for any $n\geq 0$, it holds that
+$\mathbb{E}[a_{N}]-(M-m)\cdot{\mathbb{P}}(N\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}>}\,n)\leq a_{n}\leq\mathbb{E}[a_{N}]+(M-m)\cdot{\mathbb{P}}(N\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}<}\,n).$ (Non-Decreasing Case)
+On the other hand, if $\{a_{j}\}_{j\geq 0}$ is non-increasing, i.e., $M\geq a_{0}\geq a_{1}\geq a_{2}\geq\cdots\geq m$, a similar result holds that
+$\mathbb{E}[a_{N}]-(M-m)\cdot{\mathbb{P}}(N\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}<}\,n)\leq a_{n}\leq\mathbb{E}[a_{N}]+(M-m)\cdot{\mathbb{P}}(N\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}>}\,n).$ (Non-Increasing Case)
+Here, all expectations are taken with respect to $N$.
+Remark F.21 (On the choice of random index $N$).
+When obtaining an estimate of a sequence value using the lemma above, it is unnecessary to utilize the same random variable $N$ for both upper and lower bounds.
+Namely, when the sequence is non-decreasing, a choice of $N$ with $\mathbb{E}[N]>n$ gives a small lower tail ${\mathbb{P}}(N\,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}<}\,n)$ for the upper bound; whereas an $N$ with $0<\mathbb{E}[N]<n$ gives a small upper tail ${\mathbb{P}}(N\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}>}\,n)$ for the lower bound.
+We can think of the opposite relationship for a non-increasing sequence as well.
+Moreover, to guarantee the tail bounds to be small, we often choose $N$ having a lot of mass concentrated around its mean (e.g., sub-Gaussian); then, we can take advantage of concentration inequalities associated with $N$ to have a better estimate of $a_{n}\approx\mathbb{E}[a_{N}]$.
+Poissonization (kac1949deviations; holst1986birthday; aldous1989probability; jacquet1998analytical; johansson1998longest; borodin2000asymptotics) can be thought of as a special case of randomization using Poisson random variables.
+A Poisson random variable $N\sim\mathrm{Poisson}(\lambda)$ with a parameter $\lambda>0$ is equipped with a probability mass function
+${\mathbb{P}}(N=n)=\frac{e^{-\lambda}\lambda^{n}}{n!}.\qquad(n=0,1,2,\cdots)$ (Poisson)
+In this section, our primary goal is to complete the proof of Lemma˜F.12, asserting that the Poissonization technique is particularly advantageous for analyzing a monotone property (Def.˜F.11) of multinomial random vectors (e.g., the connectivity of a random graph whose edge connections are sampled with replacements).
+To see why, let us first review an elementary property of Poisson random variables: the sum of independent Poisson random variables is again a Poisson random variable, whose parameter is the sum of the parameters of individual Poisson variables.
+We defer the proof for brevity.
+Lemma F.22 ($\downarrow$).
+Consider mutually independent Poisson random variables $Z_{i}\sim\mathrm{Poisson}(\lambda_{i})$ ($i=1,\cdots,m$). Then, $\sum_{i=1}^{m}Z_{i}\sim\mathrm{Poisson}\left(\sum_{i=1}^{m}\lambda_{i}\right)$.
+Next, we explore the relationship between multinomial and Poisson distributions by establishing an equivalence between the probability associated with multinomial random vectors and the conditional probability of mutually independent Poisson random variables whose sum is fixed.
+As a result, we can construct a Poissonization of a sequence of probabilities by regarding the multinomial random vector’s parameter $n$ as a Poisson random variable.
+The detailed formal statement is below, although we again defer its proof.
+Lemma F.23 ($\downarrow$).
+Let $\mathcal{A}$} be any property of a (random) $m$-dimensional vector: we write $(z_{1},\cdots,z_{m})\in{\mathcal{A}}$ if $(z_{1},\cdots,z_{m})$ satisfies the property $\mathcal{A}$}.
+Then, for any $\lambda>0$, and $p_{1},\cdots,p_{m}>0$ that sums to 1 (i.e., $\sum_{i=1}^{m}p_{i}=1$),
+${\mathbb{P}}_{\mathrm{Multi}(n)}\Bigl((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\Bigr)={\mathbb{P}}_{\mathrm{Po}(\lambda)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\;\middle|\;\sum_{i=1}^{m}Z_{i}=n\right).$
+Here, ${\mathbb{P}}_{\mathrm{Multi}(n)}$ is the probability measure under $(Z_{1},\ldots,Z_{m})\sim\mathrm{Multinomial}\left(n;p_{1},\ldots,p_{m}\right)$, whereas ${\mathbb{P}}_{\mathrm{Po}(\lambda)}$ is the probability measure under $Z_{i}\sim\mathrm{Poisson}(\lambda p_{i})$ ($i=1,\cdots,m$) which are mutually independent.
+As a result, it holds that
+$\mathbb{E}_{N\sim\mathrm{Poisson}(\lambda)}\left[{\mathbb{P}}_{\mathrm{Multi}(N)}\Bigl((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\Bigr)\right]={\mathbb{P}}_{\mathrm{Po}(\lambda)}\Bigl((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\Bigr).$
+We now move our attention to a monotone property of (finite-dimensional) vectors, of which we recall the definition again.
+See F.11
+A monotone property of multinomial random vectors has an interesting feature: the probability of satisfying the property is also monotone in the parameter $n$ of the multinomial distribution (Lemma˜F.25).
+This is roughly because the multinomial distribution is inspired by multiple independent trials of with-replacement sampling: an additional trial corresponds to the increase of the parameter $n$ by 1.
+This intuitive explanation can be formalized into the following lemma.
+Lemma F.24 ($\downarrow$).
+Consider a random vector $(Z_{1},\cdots,Z_{m})\!\sim\!\mathrm{Multinomial}(n;p_{1},\cdots,p_{m})$ for $n\geq 1$ and $p_{1},\cdots,p_{m}>0$ such that $\sum_{i=1}^{m}p_{i}=1$.
+That is, for any non-negative integers $z_{1},\cdots,z_{m}$ whose sum is $n$,
+${\mathbb{P}}(Z_{1}=z_{1},\cdots,Z_{m}=z_{m})=\frac{n!}{z_{1}!\cdots z_{m}!}\cdot p_{1}^{z_{1}}\cdots p_{m}^{z_{m}}.$ (Multinomial)
+Also, consider a categorical (so-called multinoulli) random variable $j\sim\mathrm{Categorical}(p_{1},...,p_{m})$ which can have a value among the integers $\{1,2,\cdots,m\}$, i.e., for each $i=1,\cdots,m$,
+${\mathbb{P}}(j=i)=p_{i}.$ (Categorical)
+Then, if we let $\widetilde{Z}_{i}=Z_{i}+\mathbbm{1}_{\{j=i\}}$ for $1\leq i\leq m$, we have a new multinomial random vector as follows:
+$(\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})\sim\mathrm{Multinomial}(n+1;p_{1},\cdots,p_{m}).$
+Using the above as a key lemma, we can prove the following lemma about the monotone property of a multinomial random vector.
+Lemma F.25 ($\downarrow$).
+Recall the definition of ${\mathbb{P}}_{\mathrm{Multi}(n)}$ from Lemma˜F.23. Fix any integers $1\leq n\leq n^{\prime}$. If a property $\mathcal{A}$} is non-decreasing, then
+${\mathbb{P}}_{\mathrm{Multi}(n)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right)\leq{\mathbb{P}}_{\mathrm{Multi}(n^{\prime})}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right).$
+The direction of the inequality should be opposite ($\geq$”) if the property is non-increasing.
+Most importantly, thanks to Lemmas˜F.23 and F.25, we can apply the de-randomization lemma (Lemma˜F.20) for the sequence $\{a_{n}\}$ of probabilities defined with a monotone property $\mathcal{A}$} as
+$a_{n}:={\mathbb{P}}_{\mathrm{Multi}(n)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right).$
+The only things left to obtain a complete (de)-Poissonization lemma for upper/lower-bounding the sequence $a_{n}$ are the tail probability bounds for a Poisson distribution.
+The concentration inequalities for the Poisson distribution are already well-known, [^8] For example, see a document online: “A short note on Poisson tail bounds.” although we derive slightly different forms of them for our own purposes.
+Lemma F.26 ($\downarrow$).
+If $N\sim\mathrm{Poisson}(n-{\varepsilon})$ for any $n>0$ and $0<{\varepsilon}<n$, it holds that
+$\textstyle{\mathbb{P}}(N\geq n)\leq\exp\left(-\frac{{\varepsilon}^{2}}{2n}\right).$ (Upper Tail Bound)
+On the other hand, if $N\sim\mathrm{Poisson}(n+{\varepsilon})$ for any $n>0$ and $0<{\varepsilon}<cn$ for some $c\in(0,3)$,
+$\textstyle{\mathbb{P}}(N\leq n)\leq\exp\left(-\frac{(3-c){\varepsilon}^{2}}{6n}\right).$ (Lower Tail Bound)
+As a result, we finally summarize the arguments in this section as the following de-Poissonization lemma for a monotone property of multinomial random vectors:
+See F.12
+Proof of Lemma˜F.12.
+This is the combination of Lemmas˜F.20, F.23, F.25 and F.26. ∎
+We use this de-Poissonization lemma in the proof of our main theorem (Theorems˜F.7 and F.8).
+Lastly, we conclude this section by presenting all the deferred proofs.
+Proof of Lemma˜F.20.
+In the first part of the proof, we deal with the non-decreasing case.
+Observe that, since $a_{N}$ has a finite expectation ($m\leq\mathbb{E}[a_{N}]\leq M$), the infinite sum $\sum_{j\geq 0}p_{j}a_{j}$ converges.
+Now, let us first obtain the upper bound.
+Since $\sum_{j\geq 0}p_{j}=1$,
+$a_{n}$ $=\sum_{0\leq j<n}p_{j}a_{n}+\sum_{j\geq n}p_{j}a_{n}$
+$\leq\sum_{0\leq j<n}p_{j}a_{n}+\sum_{j\geq n}p_{j}a_{j}$ $(\because a_{j}\geq a_{n}\text{ if }j\geq n)$
+$=\sum_{0\leq j<n}p_{j}(a_{n}-a_{j})+\sum_{j\geq 0}p_{j}a_{j}$
+$\leq(M-m)\cdot\sum_{0\leq j<n}p_{j}+\sum_{j>0}p_{j}a_{j}$ $(\because a_{n}-a_{j}\leq M-m)$
+$=(M-m)\cdot{\mathbb{P}}(N<n)+\mathbb{E}[a_{N}].$
+Likewise, we obtain the lower bound.
+$a_{n}$ $=\sum_{0\leq j\leq n}p_{j}a_{n}+\sum_{j>n}p_{j}a_{n}$
+$\geq\sum_{0\leq j\leq n}p_{j}a_{j}+\sum_{j>n}p_{j}a_{n}$ $(\because a_{j}\leq a_{n}\text{ if }j\leq n)$
+$=\sum_{j\geq 0}p_{j}a_{j}-\sum_{j>n}p_{j}(a_{j}-a_{n})$
+$\geq\sum_{j\geq 0}p_{j}a_{j}-(M-m)\cdot\sum_{j>n}p_{j}$ $(\because a_{n}-a_{j}\geq m-M)$
+$=\mathbb{E}[a_{N}]-(M-m)\cdot{\mathbb{P}}(N>n).$
+These imply the inequality (Non-Decreasing Case) and prove the first part of the lemma.
+The second part for the non-increasing sequence $\{a_{n}\}$ directly follows by applying the first part to the non-decreasing sequence $\{b_{n}\}$ defined as $b_{n}:=M+m-a_{n}$. Indeed, we have
+$\mathbb{E}[M+m-a_{N}]-(M-m)\cdot{\mathbb{P}}(N>n)\leq M+m-a_{n}\leq\mathbb{E}[M+m-a_{N}]+(M-m)\cdot{\mathbb{P}}(N<n),$
+which implies the inequality (Non-Increasing Case) and concludes the proof of the lemma. ∎
+Proof of Lemma˜F.22.
+Denote the sum of random variables by $\bar{Z}_{m}:=\sum_{i=1}^{m}Z_{i}$.
+Likewise, let us write $\bar{\lambda}_{m}:=\sum_{i=1}^{m}\lambda_{i}$.
+To prove the lemma, we proceed with the induction on $m\geq 1$.
+Since the base case is obvious ($Z_{1}\sim\mathrm{Poisson}(\lambda_{1})$), let us assume $m\geq 2$ and compute the probability mass.
+The inductive assumption says $\bar{Z}_{m-1}\sim\mathrm{Poisson}\left(\bar{\lambda}_{m-1}\right)$, which is independent of $Z_{m}$ due to the mutual independence of $Z_{1},\cdots,Z_{m}$.
+Then, for any non-negative integer $\bar{z}$,
+${\mathbb{P}}\left(\bar{Z}_{m}=\bar{z}\right)$ $=\sum_{z_{m}=0}^{\bar{z}}{\mathbb{P}}(\bar{Z}_{m-1}=\bar{z}-z_{m},Z_{m}=z_{m})$
+$=1\cdot\sum_{z_{m}=0}^{\bar{z}}{\mathbb{P}}(\bar{Z}_{m-1}=\bar{z}-z_{m})\cdot{\mathbb{P}}(Z_{m}=z_{m})\qquad(\because\bar{Z}_{m-1}\perp Z_{m})$
+$=\frac{\bar{z}!}{\bar{z}!}\cdot\sum_{z_{m}=0}^{\bar{z}}\frac{\exp\left(-\bar{\lambda}_{m-1}\right)\cdot(\bar{\lambda}_{m-1})^{\bar{z}-z_{m}}}{(\bar{z}-z_{m})!}\cdot\frac{\exp(-\lambda_{m})\cdot(\lambda_{m})^{z_{m}}}{z_{m}!}$
+$=\frac{\exp\left(-\sum_{i=1}^{m}\lambda_{i}\right)}{\bar{z}!}\cdot\sum_{z_{m}=0}^{\bar{z}}\frac{\bar{z}!}{(\bar{z}-z_{m})!\cdot z_{m}!}(\bar{\lambda}_{m-1})^{\bar{z}-z_{m}}\lambda_{m}^{z_{m}}$
+$=\frac{\exp\left(-\sum_{i=1}^{m}\lambda_{i}\right)}{\bar{z}!}\cdot\left(\bar{\lambda}_{m-1}+\lambda_{m}\right)^{\bar{z}}\qquad(\because\text{ binomial theorem})$
+$=\frac{\exp\left(-\sum_{i=1}^{m}\lambda_{i}\right)}{\bar{z}!}\cdot\left(\bar{\lambda}_{m}\right)^{\bar{z}}.$
+Since the choice of $\bar{z}$ is arbitrary, by induction, this ends up proving that $\bar{Z}_{m}\sim\mathrm{Poisson}\left(\bar{\lambda}_{m}\right)$.
+Lastly, we remark that one may prove the same result without induction, by directly applying the multinomial theorem (which is essentially a recurrent application of the binomial theorem). ∎
+Proof of Lemma˜F.23.
+By the law of total probability, for any probability measure $\mathbb{P}$} under any distribution of a random vector $(Z_{1},\cdots,Z_{m})$,
+${\mathbb{P}}((Z_{1},\cdots,Z_{m})\in{\mathcal{A}})=\sum_{(z_{1},\cdots,z_{m})\in{\mathcal{A}}}{\mathbb{P}}(Z_{1}=z_{1},\cdots,Z_{m}=z_{m}).$
+Thus, it suffices to compare the multinomial distribution and the conditional distribution of a Poisson random vector given a fixed sum: namely, we aim to show here that
+${\mathbb{P}}_{\mathrm{Multi}(n)}(Z_{1}=z_{1},\cdots,Z_{m}=z_{m})={\mathbb{P}}_{\mathrm{Po}(\lambda)}\left(Z_{1}=z_{1},\cdots,Z_{m}=z_{m}\;\middle|\;\sum\nolimits_{i=1}^{m}Z_{i}=n\right).$ (17)
+Moreover, it is sufficient for proving the first equation to study the case with non-negative integers $z_{1},\cdots,z_{m}$ such that $\sum_{i=1}^{m}z_{i}=n$; otherwise, both sides of Eq.˜17 are zero.
+In this case, observe an inclusion between events
+$\left\{Z_{1}=z_{1},\cdots,Z_{m}=z_{m}\right\}\subseteq\left\{\sum\nolimits_{i=1}^{m}Z_{i}=n\right\}.$ (18)
+Because of this, we can compute a conditional probability mass as:
+${\mathbb{P}}_{\mathrm{Po}(\lambda)}\left(Z_{1}=z_{1},\cdots,Z_{m}=z_{m}\;\middle|\;\sum\nolimits_{i=1}^{m}Z_{i}=n\right)$
+[…]
+$={\mathbb{P}}_{\mathrm{Multi}(n)}(Z_{1}=z_{1},\cdots,Z_{m}=z_{m}).$
+Therefore, we have just proved Eq.˜17.
+[…]
+$\mathbb{E}_{N\sim\mathrm{Poisson}(\lambda)}\left[{\mathbb{P}}_{\mathrm{Multi}(N)}\Bigl((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\Bigr)\right]$
+$=\sum_{j\geq 0}{\mathbb{P}}_{\mathrm{Multi}(j)}\Bigl((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\Bigr)\cdot{\mathbb{P}}(N=j)$
+$=\sum_{j\geq 0}{\mathbb{P}}_{\mathrm{Po}(\lambda)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\;\middle|\;\sum\nolimits_{i=1}^{m}Z_{i}=j\right)\cdot{\mathbb{P}}(N=j)$
+$=\sum_{j\geq 0}{\mathbb{P}}_{\mathrm{Po}(\lambda)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\text{ and }\sum\nolimits_{i=1}^{m}Z_{i}=j\right)$
+$={\mathbb{P}}_{\mathrm{Po}(\lambda)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right).$
+Proof of Lemma˜F.24.
+Take any integers $z_{1},\cdots,z_{m}$ such that $z_{1}+\cdots+z_{m}=n+1$. Let us calculate the probability mass. Applying the law of total probability,
+${\mathbb{P}}\left(\widetilde{Z}_{1}=z_{1},\cdots,\widetilde{Z}_{m}=z_{m}\right)$
+$=\sum_{i=1}^{m}{\mathbb{P}}\left(Z_{1}+\mathbbm{1}_{\{j=1\}}=z_{1},\cdots,Z_{m}+\mathbbm{1}_{\{j=m\}}=z_{m}\mid j=i\right)\cdot{\mathbb{P}}(j=i)$
+$=\sum_{i=1}^{m}{\mathbb{P}}\left(Z_{i}=z_{i}-1,Z_{r}=z_{r}\;\;(\forall r\neq i)\right)\cdot{\mathbb{P}}(j=i)\qquad(\because j\perp(Z_{1},\cdots,Z_{m}))$
+$=\sum_{i=1}^{m}\left(\frac{n!\cdot z_{i}}{z_{1}!\cdots z_{m}!}\cdot\frac{p_{1}^{z_{1}}\cdots p_{m}^{z_{m}}}{p_{i}}\right)\cdot p_{i}$
+$=\frac{n!}{z_{1}!\cdots z_{m}!}\cdot p_{1}^{z_{1}}\cdots p_{m}^{z_{m}}\cdot\sum_{i=1}^{m}z_{i}$
+$=\frac{(n+1)!}{z_{1}!\cdots z_{m}!}\cdot p_{1}^{z_{1}}\cdots p_{m}^{z_{m}}.$
+Since the choice of $z_{1},\cdots,z_{m}$ is arbitrary (given their fixed sum), it proves that $(\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})$ follows the distribution $\mathrm{Multinomial}(n+1;p_{1},\cdots,p_{m})$, as desired.
+∎
+Proof of Lemma˜F.25.
+We only consider the non-decreasing property $\mathcal{A}$} since the proof of the non-increasing case can be done symmetrically. Then, by induction, it suffices to prove
+${\mathbb{P}}_{\mathrm{Multi}(n)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right)\leq{\mathbb{P}}_{\mathrm{Multi}(n+1)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right).$ (19)
+Fix any $n\geq 1$ and consider $(Z_{1},\cdots,Z_{m})\sim\mathrm{Multinomial}(n;p_{1},\cdots,p_{m})$.
+Then, by a property of multinomial random variables (Lemma˜F.24), we have another multinomial vector $(\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})\sim\mathrm{Multinomial}(n+1;p_{1},\cdots,p_{m})$ by $\widetilde{Z}_{i}=Z_{i}+\mathbbm{1}_{\left\{j=i\right\}}$, where $j\sim\mathrm{Categorical}(p_{1},\cdots,p_{m})$.
+Observe that
+${\mathbb{P}}\left((\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})\in{\mathcal{A}}\;\middle|\;(Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right)=1,$
+since $Z_{i}\leq\widetilde{Z}_{i}$ ($\forall i$) and $\mathcal{A}$} is a non-decreasing property.
+Hence, we can derive the following by applying the law of total probability:
+${\mathbb{P}}\left((\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})\in{\mathcal{A}}\right)$
+$={\mathbb{P}}\left((\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})\in{\mathcal{A}}\;\middle|\;(Z_{1},\cdots,Z_{m})\notin{\mathcal{A}}\right)\cdot{\mathbb{P}}\left((Z_{1},\cdots,Z_{m})\notin{\mathcal{A}}\right)$
+$\phantom{=,}+{\mathbb{P}}\left((\widetilde{Z}_{1},\cdots,\widetilde{Z}_{m})\in{\mathcal{A}}\;\middle|\;(Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right)\cdot{\mathbb{P}}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right)$
+$\geq 0+1\cdot{\mathbb{P}}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right)$
+$={\mathbb{P}}_{\mathrm{Multi}(n)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right).$
+This proves Eq.˜19, concluding the proof. ∎
+Proof of Lemma˜F.26.
+Let us recall the moment generating function (MGF) of a Poisson random variable $N\sim\mathrm{Poisson}(\lambda)$ is $\mathbb{E}[e^{tN}]=\exp\left(\lambda(e^{t}-1)\right)$ ($\forall t\in{\mathbb{R}}$):
+$\mathbb{E}[e^{tN}]=\sum_{j\geq 0}e^{tj}\cdot\frac{e^{-\lambda}\lambda^{j}}{j!}=e^{-\lambda}\cdot\sum_{j\geq 0}\frac{{(\lambda e^{t})}^{j}}{j!}=\exp\left(-\lambda\right)\exp\left(\lambda e^{t}\right)=\exp\left(\lambda(e^{t}-1)\right).$
+For the upper tail (Chernoff) bound, we use $\lambda=n-{\varepsilon}$ and apply Markov inequality on $e^{tN}$: for any $n>{\varepsilon}>0$ and $t>0$,
+${\mathbb{P}}(N\geq n)={\mathbb{P}}\left(e^{tN}\geq e^{tn}\right)\leq\mathbb{E}\left[e^{tN}\right]\cdot e^{-tn}\leq\exp\left((n-{\varepsilon})(e^{t}-1)-tn\right).$
+Since $t$ is arbitrary, we take the infimum of both sides over $t>0$, obtaining
+${\mathbb{P}}(N>n)$ $\leq\inf_{t>0}\exp\left((n-{\varepsilon})(e^{t}-1)-tn\right)$
+$=\exp\left(n\left(\ln\left(1-\frac{{\varepsilon}}{n}\right)+\frac{{\varepsilon}}{n}\right)\right)$
+$\leq\exp\left(-\frac{{\varepsilon}^{2}}{2n}\right).$ $(\because\,\ln(1-u)+u\leq-\frac{u^{2}}{2}\;\;\text{ if }u>0)$
+This proves Eq.˜Upper Tail Bound.
+On the other hand, for the lower tail (Chernoff) bound, we use $\lambda=n+{\varepsilon}$ and apply Markov inequality on $e^{-tN}$: for any $n>0$, ${\varepsilon}>0$, and $t>0$,
+${\mathbb{P}}(N\leq n)={\mathbb{P}}\left(e^{-tN}\geq e^{-tn}\right)\leq\mathbb{E}\left[e^{-tN}\right]\cdot e^{tn}\leq\exp\left((n+{\varepsilon})(e^{-t}-1)+tn\right).$
+Similarly as before, taking the infimum over $t>0$,
+${\mathbb{P}}(N<n)$ $\leq\inf_{t>0}\exp\left((n+{\varepsilon})(e^{-t}-1)+tn\right)$
+$=\exp\left(n\left(\ln\left(1+\frac{{\varepsilon}}{n}\right)-\frac{{\varepsilon}}{n}\right)\right)$
+$\leq\exp\left(-\frac{{\varepsilon}^{2}}{2n}+\frac{{\varepsilon}^{3}}{6n^{2}}\right).$ $(\because\,\ln(1+u)-u\leq-\frac{u^{2}}{2}+\frac{u^{3}}{6}\;\;\text{ if }u>0)$
+Observe that, if $0<{\varepsilon}<cn$ for some $c\in(0,3)$,
+$-\frac{{\varepsilon}^{2}}{2n}+\frac{{\varepsilon}^{3}}{6n^{2}}\leq-\frac{{\varepsilon}^{2}}{2n}+\frac{{\varepsilon}^{2}}{6n}\cdot c=-\frac{(3-c){\varepsilon}^{2}}{6n}.$
+This proves Eq.˜Lower Tail Bound. ∎
+F.3.2 Binomial Random Intersection Graphs
+Let us explain the random $k$-intersection graph (singer1995random; karonski1999random; godehardt2003two).
+Consider a set $V$ of vertices and another set $W$ of items.
+Each vertex $v\in V$ is randomly assigned a subset of items $W_{v}\subset W$.
+Then, a pair of vertices $u$ and $v$ are adjacent (i.e., connected with an edge) in a random $k$-intersection graph if and only if at least $k$ items are shared between $u$ and $v$, i.e., $\left|W_{u}\cap W_{v}\right|\geq k$.
+We remark that $k$ is a universal constant throughout our paper.
+One of the most well-studied models of random intersection graphs is binomial random $k$-intersection graphs ${\mathcal{G}}^{(k)}(n,m,p)$, with $n=\left|V\right|$ and $m=\left|W\right|$.
+In the ${\mathcal{G}}^{(k)}(n,m,p)$ model, for every vertex $v\in V$, each item $w\in W$ is assigned to $v$ independently with the same probability $p\in(0,1)$.
+That is, the indicator variables $\mathbbm{1}_{\left\{\text{$w$ is assigned to $v$}\right\}}$ are i.i.d.
+Bernoulli random variables with the same parameter $p$, for all vertices $v\in V$ and items $w\in W$.
+The word binomial is attached to the name because the number of items assigned to each vertex is a random variable following $\mathrm{Bin}(m,p)$.
+We are particularly interested in the ${\mathcal{G}}^{(k)}(n,m,p)$ model because of the following observation.
+See F.13
+Proof of Lemma˜F.13.
+Recall that the vertex set of ${\mathcal{G}}_{b}^{(D,k)}$ is $V_{b}=\left\{x_{12}\in{\mathcal{X}}_{12}:f_{1}(x_{12})=b\right\}\subset{\mathcal{X}}_{12}$.
+Also, consider $W:={\mathcal{X}}_{3}$ as the set of all ‘items.’ Define a set of items assigned to each vertex $x_{12}\in V_{b}$ as
+$W_{x_{12}}=\left\{x_{3}\in W:(x_{12},x_{3})\in D\right\}=\left\{x_{3}\in W:Z_{(x_{12},x_{3})}\geq 1\right\}.$
+Observe that all vertex-item assignments are i.i.d. Bernoulli random variables with the same probability parameter $p$: for any $x_{12}\in V_{b}$ and $x_{3}\in{\mathcal{X}}_{3}$,
+$p={\mathbb{P}}_{\mathrm{Po}(\lambda)}(Z_{(x_{12},x_{3})}\geq 1)=1-{\mathbb{P}}_{\mathrm{Po}(\lambda)}(Z_{(x_{12},x_{3})}=0)=1-\exp\left(-\frac{\lambda}{\left|{\mathbb{X}}\right|}\right).$
+In addition, observe that the set $S(x_{12},x^{\prime}_{12}\mid D)$ defined in Eq.˜6 is identical to the intersection $W_{x_{12}}\cap W_{x^{\prime}_{12}}$.
+Hence, by the definition of the edge set $E_{b}^{(D,k)}$ in Eq.˜5, two distinct vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if and only if $\left|W_{x_{12}}\cap W_{x^{\prime}_{12}}\right|\geq k$.
+In summary, every $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ is a binomial random $k$-intersection graph ${\mathcal{G}}^{(k)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-\lambda/\left|{\mathbb{X}}\right|)$.
+∎
+Note that $n$, $m$, and $p$ are not necessarily independent of each other; the parameters $m=m_{n}$ and $p=p_{n,m}$ are often regarded as functions of $n$.
+Indeed, it is often studied in the literature on random graphs that a sufficient condition for the parameters (in terms of $n\to\infty$) to guarantee certain graph properties, at least asymptotically.
+Here, we review a few of the seminal results on the connectivity (as well as the disappearance of isolated vertices) of binomial random $k$-intersection graphs (singer1995random; zhao2014k; zhao2017connectivity; rybarczyk2011sharp; rybarczyk2017coupling).
+[^9] Caveat: in the literature of (random) graph theory, the letter ‘$k$’ is often used for $k$-connectivity (i.e., being connected after removing less than $k$ vertices/edges).
+On the other hand, they study ‘random $s$-intersection graphs,’ using $s$ as the intersection constraint parameter.
+The proofs are involved; thus, we omit them.
+See F.14
+See F.15
+Appendix G Additional Results for Power-Law Scaling Analysis
+G.1 Measurement Protocol for $N_{\mathrm{req}}$
 To empirically determine the minimum dataset size required for reliable compositional generalization ($N_{\mathrm{req}}$), we develop a measurement protocol that accounts for practical computational constraints while ensuring robustness.
-For each token set size $|\mathcal{X}|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
+For each token set size $\left|{\mathcal{X}}\right|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
 Specifically, our criterion for “reliable generalization” on ID is defined as:
@@
     Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion.
-This establishes that our measurement error is at most $-\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
-F.2 Measured power-law scaling constants across task structures and model sizes
+This establishes that our measurement error is at most $\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
+G.2 Measured Power-Law Scaling Exponents Across Task Structures and Model Sizes
 Using our measurement protocol, we measure the required dataset size $N_{\mathrm{req}}$ across three different compositional structures (2-Hop, Parallel-2-Hop, and 3-Hop) and three model scales (68M, 96M, and 1.5B parameters).
-For each task structure, we vary the token set size $|\mathcal{X}|$ from 50 to 200, allowing us to observe the scaling relationship.
-Table 2 presents the power-law exponents obtained by linear fitting $\log(|\mathcal{X}|)$ vs.
+For each task structure, we vary the token set size $\left|{\mathcal{X}}\right|$ from 50 to 200, allowing us to observe the scaling relationship.
+Table 2 presents the power-law exponents obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs.
 $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$.
@@
 This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
-Table 2: Power law exponents for different tasks and GPT-2 sizes, obtained by linear fitting $\log(|\mathcal{X}|)$ vs. $\log(N_{\mathrm{req}})$ plots. $R^{2}>0.99$ for all linear fitting.
+Table 2: The power-law exponents for different tasks and GPT-2 sizes, obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs.
+$\log(N_{\mathrm{req}})$ plots.
+$R^{2}>0.99$ for all linear fitting.
 Model Size 2-hop Parallel-2-hop 3-hop
@@
 1.5B 2.28 2.17 2.60
-F.3 Robustness to hyperparameter variations
-Figure 15: Robustness of power-law scaling relationship to hyperparameter variations in the 2-Hop task with $|\mathcal{X}|=50$.
+G.3 Robustness to Hyperparameter Variations
+Figure 16: Robustness of power-law scaling relationship to hyperparameter variations in the 2-Hop task with $\left|{\mathcal{X}}\right|=50$.
 Each line shows the training and test accuracy curves for a different configuration: (1) baseline, (2) reduced learning rate (4e-4, half of baseline), (3) reduced weight decay (0.01, one-tenth of baseline), and (4) changed generalization criteria (test accuracy > 0.95 within 10 epochs after training accuracy > 0.95).
@@
 To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations.
-Figure 15 demonstrates that for the 2-Hop task with $|\mathcal{X}|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
+Figure 16 demonstrates that for the 2-Hop task with $\left|{\mathcal{X}}\right|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
   1. 1.
@@
 This robustness to hyperparameter variations suggests that the power-law relationship between token set size and required dataset size is primarily a property of the compositional generalization process, rather than an artifact of specific optimization settings.
-Appendix G Detailed analysis for Non-tree task
-This section provides additional analyses that support our findings in Sec. 5.4 regarding the challenges of path ambiguity in the Non-tree task.
-G.1 Coverage analysis
-Figure 16: Coverage analysis for Non-tree task with $|\mathcal{X}|=50$.
-The graph shows the percentage of ID test data covered at different $k$ values across various dataset sizes ($N$).
-Compared to the 2-Hop task (Fig.
-3, left), Non-tree has significantly lower coverage at equivalent dataset sizes, indicating that path ambiguity impedes the formation of functional equivalence relationships.
-Fig.
-16 demonstrates that with equivalent training dataset sizes, a smaller percentage of ID test examples fall inside $k$-coverage for the Non-tree task compared to the 2-Hop task shown in Fig.
-3 (Left).
-This aligns with our theoretical analysis in Sec.
-5.4, which predicts that path ambiguity limits the establishment of functional equivalence relationships between input subsequences, as the model cannot generalize across different $x_{2}$ values in the Non-tree structure even when they produce the same intermediate state $b=f_{1}(x_{1},x_{2})$.
-G.2 Effect of model scaling
-Figure 17: ID test accuracy comparison between GPT-2 (96M parameters) and GPT-2-XL (1.5B parameters) on the Non-tree task with $|\mathcal{X}|=50$, measured 100 epochs after training accuracy exceeds 0.99.
+Appendix H Detailed Analysis for Non-tree Task
+This section provides additional analyses that support our findings in Sec.˜7 regarding the challenges of path ambiguity in the Non-tree task.
+H.1 Coverage Analysis
+Figure 17: Coverage analysis for Non-tree task with $\left|{\mathcal{X}}\right|=50$.
+The graph shows the percentage of ID test data covered at different $k$ values across various dataset sizes (NN).
+Compared to the 2-Hop task (Fig.˜3, left), Non-tree has significantly lower coverage at equivalent dataset sizes, indicating that path ambiguity impedes the formation of functional equivalence relationships.
+Fig.˜17 demonstrates that with equivalent training dataset sizes, a smaller percentage of ID test examples fall inside $k$-coverage for the Non-tree task compared to the 2-Hop task shown in Fig.˜3 (Left).
+This aligns with our theoretical analysis in Sec.˜7, which predicts that path ambiguity limits the establishment of functional equivalence relationships between input subsequences, as the model cannot generalize across different $x_{2}$ values in the Non-tree structure even when they produce the same intermediate state $b=f_{1}(x_{1},x_{2})$.
+H.2 Effect of Model Scaling
+Figure 18: ID test accuracy comparison between GPT-2 (96M parameters) and GPT-2-XL (1.5B parameters) on the Non-tree task with $\left|{\mathcal{X}}\right|=50$, measured 100 epochs after training accuracy exceeds 0.99.
 Despite the 15x increase in parameter count, the accuracy does not increase.
-Fig.
-17 shows that scaling up the model size to GPT-2-XL (1.5B parameters) does not significantly improve generalization performance on the Non-tree task, even when measured 100 epochs after reaching training accuracy > 0.99.
+Fig.˜18 shows that scaling up the model size to GPT-2-XL (1.5B parameters) does not significantly improve generalization performance on the Non-tree task, even when measured 100 epochs after reaching training accuracy > 0.99.
 This suggests that the challenges posed by path ambiguity cannot be overcome simply by increasing model capacity, supporting our claim that the limitation is structural rather than related to model capacity.
-G.3 Representation analysis in successful generalization
-For a model that eventually achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $|\mathcal{X}|=50$, $N=50\text{k}$), we conduct causal tracing analysis to understand how it achieves generalization despite path ambiguity.
+H.3 Comparison between Mamba and GPT-2 on Non-tree Task
+Figure 19: ID test accuracy comparison between GPT-2 and Mamba on the Non-tree task with $\left|{\mathcal{X}}\right|=50$, measured 100 epochs after training accuracy exceeds 0.99.
+Fig.˜19 shows that the Mamba model (4 layers, hidden dimension of 256, trained with learning rate of 0.008) shows a similar trend of ID test accuracy on Non-tree task compared to GPT-2, suggesting that the generalization failure is more likely due to the task structure itself, rather than a specific model architecture.
+H.4 Representation Analysis in Successful Generalization
+For a model that eventually achieved near-perfect ID accuracy (0.96) after extended training (36k epochs, $\left|{\mathcal{X}}\right|=50$, $N=50\text{k}$), we conduct causal tracing analysis to understand how it achieves generalization despite path ambiguity.
 Figure 20: Causal tracing analysis for the Non-tree model after extended training.
@@
 Right: different $b$ and $x_{2}$.
-The causal tracing results in Fig.
-18 reveal how the model achieves generalization in the presence of path ambiguity.
+The causal tracing results in Fig.˜20 reveal how the model achieves generalization in the presence of path ambiguity.
 Across all perturbation strategies, the model’s predictions show strong causal dependence on representations at both the $x_{1}$ and $x_{2}$ positions, indicating reliance on direct access to both input tokens rather than an abstracted intermediate computation.
@@
 This represents a fundamentally different solution strategy compared to the 2-Hop task, with implications for both generalization capability and interpretability.
-Appendix H Detailed discussion on the taxonomy of generalization
-In Section 7, we introduce a taxonomy of generalization that extends beyond the coverage principle.
-We propose this taxonomy as an initial conceptual framework to disentangle different mechanisms of generalization that are often conflated in the literature.
-This categorization offers hypotheses for why neural networks succeed or fail at compositional tasks, connecting empirical observations to theoretical principles.
-We emphasize that the taxonomy is preliminary and that Sec.
-H.5 outlines its current limits, open questions, and future directions.
-H.1 Type-I: Structure-based generalization
-Type-I generalization occurs when a model succeeds by identifying and utilizing functionally equivalent components based on how primitive functions are composed.
-This form of generalization is precisely what we have formalized and studied via the coverage principle.
-Crucially, this kind of generalization remains strictly bounded by coverage as defined in Section 3.
-In Type-I generalization, a model learns to recognize when different input fragments yield identical results in shared contexts, allowing it to generalize to new combinations of these fragments.
-However, without exposure to sufficient functional equivalence evidence during training, the model cannot reliably generalize to novel compositions.
-Type-I generalization effectively describes what a pattern matching learner can achieve through the functional equivalence observed by input-output relationships, without requiring abstract variable binding mechanisms.
-This limitation aligns with the classic argument by Fodor and Pylyshyn [9] and Marcus [37] that connectionist networks lacking explicit variable binding capabilities [^13] We expand on the variable binding concept defined in Sec.
-2 beyond flexibly associating variables with values, and we distinguish between self-replicative expansion (e.g., 2-hop $\to$ 3-hop via Type-III) and complex compositional binding (e.g., substituting noun phrases for nouns), where the latter may require explicit variable binding mechanisms beyond parameter sharing.
-cannot exhibit fully systematic generalization.
-We introduce this perspective as a conceptual framing of the coverage principle: it describes the ceiling of pattern-based generalization in neural networks that rely on pattern matching rather than symbolic variable binding.
-H.2 Type-II: Function property-based generalization
-Type-II generalization occurs when a model exploits specific properties intrinsic to individual primitive functions, enabling generalization beyond patterns observed during training.
-Unlike Type-I generalization which requires explicit evidence of functional equivalence, Type-II generalization leverages invariant properties that hold for all inputs of the primitive, making the function’s output fully determined by those invariances even in unobserved contexts.
-We identify several sub-families of property-based generalization:
-  • Algebraic invariances: commutativity, associativity, and similar properties
-  • Group-theoretic structure: as in Fourier basis representations for modular arithmetic [63]
-  • Input irrelevance: attention sparsity that ignores distractors [81]
-The reversal curse [16] exemplifies the layered nature of compositional challenges across multiple generalization types.
-The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”.
-Property-based approaches like bidirectional training [83] partially succeed by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data.
-However, recent work [42] reveals deeper representational binding constraints: even though models can exploit inverse mapping properties in idealized conditions (abstract concept-level tasks), they struggle to maintain consistent concept representations when moving to realistic surface-level predictions, where entities switch roles between subject and object positions.
-This demonstrates how real compositional tasks involve multiple mechanism layers, where achieving one capability may be insufficient due to representational binding limitations.
-Type-II generalization fundamentally differs from Type-I in that it can, in principle, transcend the limitations imposed by the coverage principle, allowing models to generalize correctly to inputs that lie outside the coverage of their training data by leveraging invariant properties of the underlying functions.
-Systematic characterization of pure Type-II extrapolation remains limited and is an important direction for future work.
-H.3 Type-III: Shared-operator generalization
-Type-III generalization emerges through reuse of identical primitive functions across different computational positions, e.g., when $f_{1}=f_{2}$ in a multi-hop reasoning task.
-Unlike Type-I (bounded by coverage) and Type-II (exploiting algebraic properties), Type-III generalization leverages the structural repetition of identical operations.
-This mechanism is not unique to Transformers.
-Recurrent architectures [84] exemplify Type-III generalization through their recurrent weight sharing: the same transformation is applied at each time step, enabling the network to process sequences of varying lengths despite training on fixed-length examples.
-This architectural inductive bias directly encodes the assumption that the same operation should apply across temporal positions, allowing recurrent architectures to extrapolate to longer sequences [85].
-In the Transformer context, Wang et al.
-[19] demonstrate that Universal Transformer layers [86] enable generalization beyond in-domain data in two-hop tasks, while [87] show improved compositional reasoning with weight-shared layers.
-These findings suggest parameter sharing as a key enabling mechanism, where a single learned implementation of a function can be applied across multiple contexts.
-A prominent manifestation of Type-III generalization appears in depth extrapolation and length generalization tasks, where models generalize to deeper or longer compositional structures than seen during training [99, 100, 101, 102].
-While length generalization tasks fundamentally involve applying the same computation to longer inputs (Type-III), successful generalization often requires additional mechanisms like specialized positional encodings that facilitate computation reuse across positions.
-The observed brittleness of current approaches [103] suggests that such inductive biases are necessary but may not be sufficient for robust systematicity.
-In summary, Type-III generalization extends beyond the coverage-bound limitations of Type-I by exploiting identical operations, without requiring the algebraic property understanding characteristic of Type-II.
-Type-III therefore complements Types I and II, but may still operate without explicit variable binding.
-H.4 Relationship to prior taxonomies of generalization
-Prior taxonomies of generalization in neural networks have primarily categorized the phenomena of compositional generalization rather than the underlying mechanisms.
-Lake and Baroni [12] distinguish between “mix-and-match” generalization (for small training-test differences) and systematic compositional generalization (requiring abstract rule application).
-Other frameworks focus on empirical evaluations through benchmarks [23] or decomposing compositional abilities into constituent skills [38].
-Our taxonomy differs by categorizing the mechanisms that enable generalization rather than categorizing based on the phenomena.
-By distinguishing between structure-based pattern matching (Type-I), algebraic property exploitation (Type-II), and shared-operator transfer (Type-III), our framework provides a lens for interpreting why models succeed or fail at certain compositional tasks.
-This mechanistic approach not only explains observed behaviors but also offers hypotheses about when Types II or III become necessary for certain generalization challenges.
-H.5 Open questions and future directions
-To operationalize this taxonomy, future work could develop diagnostics: coverage ablation tests for Type-I (systematically removing training examples to test coverage boundaries), algebraic identity probes for Type-II (testing whether models recognize properties like commutativity without explicit training), and parameter-sharing ablations for Type-III (comparing shared vs.
-independent weights across positions).
-Current examples in this paper are illustrative rather than definitive classifications; real tasks often involve combinations of mechanisms, and our taxonomy provides a scaffold for disentangling them rather than enforcing mutually exclusive categories.
-Several fundamental questions remain.
-First, practical computations often involve mixed ingredients of generalization, spanning multiple types.
-When classification is ambiguous, we suggest identifying the dominant mix of types that contribute, as most realistic tasks involve combinations rather than pure instances of any single type.
-Second, identifying the true compositional structure underlying a task can be ambiguous, as Locatello et al.
-[52] demonstrate regarding multiple valid decompositions.
-Third, real-world datasets contain heterogeneous computational structures, and it remains unclear how neural networks accommodate this diversity.
-Fourth, none of the three mechanisms solves the variable binding problem, and devising diagnostics and architectures for binding remains crucial.
-Future work could explore connections to complementary frameworks:
-  • Minimum Description Length [73]: Type-II may compress regularities that lower description length. Similarly, Type-III can be viewed as compression through parameter reuse.
-  • Formal language learning [104, 105]: Connections to grammatical inference and automata learning
-  • Cognitive science [106]: Human compositional learning strategies that may inspire architectural innovations
-This taxonomy is not intended as a complete classification but rather as a conceptual framework to organize our understanding of generalization mechanisms and stimulate further research into the fundamental capabilities and limitations of neural networks on compositional tasks.
-Ultimately, developing architectural and optimization strategies that enable neural networks to leverage all three types more robustly and ultimately addressing the variable binding problem represents an important frontier for addressing the longstanding challenge of systematicity in connectionist systems.
-Appendix I Partial computation observation drives the alignment of functional equivalence representation and vocabulary space
-In this section, we investigate how exposure to partial computations affects the interpretability of intermediate state representations through vocabulary space alignment.
-We compare two training conditions on a modified 2-Hop task with $|\mathcal{X}|=50$ and $N=10\text{k}$, after 40k epochs of training:
-  1. 1.
-    Standard Training: $f_{1}\neq f_{2}$, model only sees complete two-hop examples $(x_{1},x_{2},x_{3})\mapsto t$.
-  2. 2.
-    With Partial Computation: $f_{1}=f_{2}$, model additionally sees all possible partial computations $(x_{1},x_{2})\mapsto b$ where $b=f_{1}(x_{1},x_{2})$ (2,500 partial examples, not counted toward the $N=10\text{k}$ two-hop training data).
-To assess interpretability, we measure the Mean Reciprocal Rank (MRR) of intermediate state representations when projected to vocabulary space using the unembedding matrix.
-Low MRR indicates that the model’s internal representation of intermediate state $b$ aligns with the corresponding vocabulary token.
-Fig.
-19 shows a striking contrast between the two conditions.
-Under standard training, the MRR score remains very high throughout training, indicating that intermediate representations are not aligned with vocabulary space despite the model successfully learning the compositional task.
-However, when partial computations are included, the MRR score becomes very high, demonstrating clear vocabulary alignment.
-Figure 19: MRR scores for intermediate state representations projected to vocabulary space.
-Left: Standard training ($f_{1}\neq f_{2}$, no partial computation) shows very high MRR regardless of position and layer.
-Right: Training with partial computation ($f_{1}=f_{2}$, with partial examples) shows MRR of 0 in layers 3 to 8 at position $x_{2}$, indicating strong vocabulary alignment.
-This experiment suggests that logit lens interpretability is orthogonal to functional equivalence representation formation.
-A model can develop functionally correct intermediate representations that enable compositional generalization while remaining completely uninterpretable through standard vocabulary projection techniques.
-Interpretability via logit lens requires explicit vocabulary anchoring through exposure to partial computations that map intermediate states to vocabulary tokens.
-This finding has important implications for mechanistic interpretability research: the absence of interpretable representations through logit lens does not indicate the absence of structured internal computation.
-Furthermore, it suggests that interpretability techniques may need to account for how training data shapes the alignment between internal representations and vocabulary space, rather than assuming such alignment emerges naturally from task performance.
+Appendix I Detailed Discussion on the Taxonomy for Understanding Generalization Mechanisms
+In this section, we initiate a discussion to disambiguate the mixed mechanisms of generalization into isolated testable parts by sketching a preliminary taxonomy that distinguishes three complementary mechanisms of generalization.
+We note that we do not view our categorization as a complete one.
+Type-I: Functional equivalence-based generalization (pattern matching).
+This is precisely what we formalized through this work: models learn that different input fragments yield identical results in shared contexts, enabling generalization to new fragment combinations.
+Crucially, this generalization remains bounded by coverage, and reliable generalization fails without sufficient functional equivalence evidence.
+In other words, it describes the ceiling of pattern matching.
+Type-II: Function property-based generalization.
+This mechanism exploits intrinsic properties of individual primitive functions, e.g., algebraic invariances such as commutativity or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present (wen2024sparse)).
+Unlike the previous type, this mechanism explains the generalization beyond the coverage by leveraging ‘global’ properties that hold across all possible inputs of a primitive, beyond what is actually observed.
+We interpret the Reversal Curse phenomenon (berglund2023reversal) as an example of the layered nature of challenges across multiple generalization types.
+Our framework predicts the failure of pattern matching on this problem, since training on “$A$ is BB” provides no functional equivalence evidence for “$B$ is $A^{-1}$”.
+An architectural modification to learn inverse mappings from the same training data to handle this problem (lv2024analysis) can be interpreted as a utilization of Type-II generalization to enable generalization beyond coverage.
+Type-III: Shared-operator generalization.
+This mechanism emerges through the reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$).
+Recurrent architectures (hochreiter1997long) exemplify the utilization of this through weight sharing across time steps, enabling processing of variable-length sequences (graves2014neural).
+Similarly, it has been reported in Transformers with inductive biases towards reuse of the same computation through parameter sharing (dehghani2018universal; csordas2021devil; wang2024grokked) can improve generalization on complex compositional tasks where the same primitive function can be reused in various contexts.
+We interpret this mechanism as exploiting structural repetition.
+Distinguishing mechanisms from phenomena.
+Compared to prior categorizations of generalization, which focus on observed phenomena (lake2018generalization; hupkes2020compositionality), we categorize the underlying mechanisms.
+As noted in Sec.˜1, many behavioral studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure.
+We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.
+Implications and future directions.
+Real compositional tasks typically involve combinations of all three types (and possibly more).
+While preliminary, we believe this taxonomy guides future research design on constructive characterization of neural networks’ generalization behaviors on discrete sequence tasks.
+In this broader context, this work can be understood as a characterization and formalization of pattern-matching generalization to clarify its specific boundaries.
+When models succeed beyond our coverage predictions, we view these as exploiting other generalization mechanisms, i.e., beyond pattern matching.
+Our focused study suggests that challenges to reliable generalization remain as long as models rely primarily on pattern matching, requiring methodological innovations that harness non-pattern-matching mechanisms, e.g., variable binding.
+We hope this preliminary taxonomy serves as a research program towards our better understanding of generalization, and confirming or refuting its utility is an empirical matter that we invite the community to explore.

### v3

@@
-Characterizing Pattern Matching and Its Limits on Compositional Task Structures
+arXiv:2505.20278v3[cs.LG] 02 Mar 2026
 Abstract
@@
 1 Introduction
-Despite the remarkable performance of Large Language Models (LLMs) (brown2020language; touvron2023llama), compositional generalization studies suggest that the core mechanism of generalization might be “pattern matching”, i.e., models learning local statistical regularities between input fragments and outputs in some cases (loula2018rearranging; johnson2017clevr; berglund2023reversal; wang2024grokked; mirzadeh2024gsm; keysers2019measuring; csordas2022ctl++).
+Despite the remarkable performance of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023), compositional generalization studies suggest that the core mechanism of generalization might be “pattern matching,” i.e., models learning local statistical regularities between input fragments and outputs in some cases (Loula et al., 2018; Johnson et al., 2017; Berglund et al., 2024; Wang et al., 2024a; Mirzadeh et al., 2025; Keysers et al., 2020; Csordás et al., 2022).
 However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), discussing pattern matching without a precise definition and diagnosing it post-hoc rather than characterizing it predictively.
-As a result, it remains unclear which behaviors should count as pattern matching and which should not, obscuring a constructive and testable account of its boundary.
-To make this notion precise, we (1) introduce a model-agnostic, data-centric formalism for pattern matching and (2) systematically study how modern architectures like decoder-only Transformers (vaswani2017attention) and Mamba (gu2023mamba) perform generalization through pattern matching.
-Specifically, we first propose a model-agnostic and data-centric definition of pattern matching by formalizing the substitution of input patterns observed to result in identical outputs in shared contexts as functional-equivalence (Sec.˜3; henceforth, we use pattern matching as equivalent to functional-equivalence-based generalization).
-This induces a coverage boundary: if learning relies only on such evidence, reliable prediction is expected only for test inputs reachable by these substitutions.
-Moreover, to isolate and study pattern-matching behaviors, we use controlled setups that deliberately remove other generalization sources and make functional equivalence the primary available mechanism.
-With this setting, our formalism yields predictive and quantitative insights about the limitations of pattern matching that, to our knowledge, have not been well characterized in prior works:
-  • Generalization success is well predicted by the number of supporting contexts that witness the relevant functional equivalence.
-Mechanistically, Transformers implement functional equivalence via clustered intermediate representations at specific layers/positions, with clustering strength aligning with evidence strength (Sec.˜5).
-  • We prove a tight sample complexity bound of pattern matching on a two-hop structure by identifying the exponent of the data scaling law (in terms of the token set size) which is necessary and sufficient for perfect in-domain generalization (Theorem˜6.1).
-The measured power-law exponent agrees with our theoretical bounds and remains stable under roughly $20\times$ parameter increase (from 68M to 1.5B) for GPT-2 (radford2019language), and also holds for Mamba (gu2023mamba) architecture (Sec.˜6).
+As a consequence, it remains unclear which behaviors should be counted as pattern matching and which should not, thereby obscuring a constructive and testable account of its boundary.
+To make this notion precise, we (1) introduce a data-centered formalism for pattern matching and (2) systematically study how modern architectures like decoder-only Transformers (Vaswani et al., 2017; Radford et al., 2018; 2019) and Mamba (Gu and Dao, 2024) perform generalization through pattern matching.
+Specifically, we first propose a model-agnostic and data-centric definition of pattern matching by formalizing the substitution of input patterns observed to result in identical outputs in shared contexts as functional-equivalence (Def.˜3.1; henceforth, we use pattern matching as equivalent to functional-equivalence-based generalization).
+This induces a coverage boundary (Def.˜3.2): if learning relies only on such evidence, there is no hope to expect any reliable predictions for the test inputs unreachable by these substitutions.
+Moreover, to isolate and study pattern-matching behaviors, we use controlled setups that deliberately remove other sources of generalization and make functional equivalence the primary available mechanism.
+With these setups, our formalism offers predictive and quantitative insights about the limitations of pattern matching.
+To the best of our knowledge, these have not been well characterized yet in prior works:
+  • The success in generalization is well predicted by the number of supporting contexts that witness the relevant functional equivalence.
+Mechanistically, Transformers seem to implement the functional equivalence by clustering intermediate representations at specific layers/positions.
+The strength of the clustering aligns with the evidence strength (Sec.˜5).
+  • We prove a tight sample complexity bound of pattern matching on a two-hop structure by identifying the exponent of the data scaling law (in terms of the token set size) which is necessary and sufficient to achieve perfect in-domain generalization (Theorem˜6.1).
+We find that the measured power-law exponent agrees with our theoretical bounds, remains stable under roughly $20\times$ parameter increase (from 68M to 1.5B) for GPT-2 (Radford et al., 2019), and also holds for Mamba (Gu and Dao, 2024) architecture (Sec.˜6).
   • When the same variable influences the output along multiple computational paths, models fail to form unified intermediate state representations.
-Analysis reveals that they instead develop context-dependent state representations, impairing both generalization and interpretability (Sec.˜7).
-  • Chain-of-Thought (CoT) supervision (wei2022chain) reduces data requirements yet does not resolve path ambiguity without seeing nearly exhaustive in-domain combinations (Sec.˜8).
+Our analysis reveals that they instead develop context-dependent state representations, impairing both generalization and interpretability (Sec.˜7).
+  • Chain-of-Thought (CoT) supervision (Wei et al., 2022) does not completely resolve the path ambiguity without observing nearly exhaustive in-domain combinations, although it reduces the data requirements to some extent (Sec.˜8).
 Finally, we situate this characterization of pattern matching within a mechanism-based taxonomy of generalization mechanisms, proposing two additional distinguishable mechanisms of generalization in compositional tasks: property-based and shared-operator generalization (Sec.˜9.2 and I).
@@
 Pattern matching behaviors of LLMs on compositional tasks.
-It is well perceived that pattern matching alone is inadequate for systematic generalization (fodor1988connectionism), and modern LLMs display generalization abilities that seem to be far beyond what pattern matching alone can do, as measured by their remarkable performance on complex benchmarks (achiam2023gpt).
-However, a growing body of work has consistently reported that LLMs still fall short on benchmarks designed to test compositionality (hupkes2020compositionality), including mathematical reasoning (mirzadeh2024gsm), multi-hop reasoning (yang2024large; wang2024grokked), and more (lake2018generalization; kim2020cogs; csordas2022ctl++; dziri2024faith).
+It is well perceived that pattern matching alone is inadequate for systematic generalization (Fodor and Pylyshyn, 1988), and modern LLMs display generalization abilities that seem to be far beyond what pattern matching alone can do, as measured by their remarkable performance on complex benchmarks (OpenAI, 2023).
+However, a growing body of work has consistently reported that LLMs still fall short on benchmarks designed to test compositionality (Hupkes et al., 2020), including mathematical reasoning (Mirzadeh et al., 2025), multi-hop reasoning (Yang et al., 2024; Wang et al., 2024a), and more (Lake and Baroni, 2018; Kim and Linzen, 2020; Csordás et al., 2022; Dziri et al., 2023).
 This gap between their capabilities and pattern-matching behaviors on compositional tasks calls for a principled framework to define what pattern matching is and to what extent a model’s behavior can be attributed to pattern matching, but it is mostly discussed with behavioral studies under the context of a specifically designed benchmark.
@@
 Mechanistic interpretability.
-Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors (elhage2021mathematical; olsson2022context; nanda2023progress; elhage2022toy).
-Recent work analyzes how Transformer components are causally related to certain behaviors (meng2022locating; hanna2023does; goldowsky2023localizing).
-In particular, it is reported that in-domain compositional generalization can emerge through grokking, with identifiable intermediate state representations inside Transformers (wang2024grokked).
+Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors (Elhage et al., 2021; Olsson et al., 2022; Nanda et al., 2023; Elhage et al., 2022).
+Recent work analyzes how Transformer components are causally related to certain behaviors (Meng et al., 2022; Hanna et al., 2023; Goldowsky-Dill et al., 2023).
+In particular, it is reported that in-domain compositional generalization can emerge through grokking, with identifiable intermediate state representations inside Transformers (Wang et al., 2024a).
 Our framework complements these works by providing mechanistic insights about pattern matching.
-Our findings also explain why standard interpretability techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) may fail to identify state representations in models trained on tasks with path ambiguities.
+Our findings also explain why standard interpretability techniques like logit lens (nostalgebraist, 2020; Belrose et al., 2023) may fail to identify state representations in models trained on tasks with path ambiguities.
 3 Formalizing Pattern Matching with Functional Equivalence
@@
 The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is chosen from a finite set $\mathcal{X}$}.
-[^1] For brevity, we use a shared token set $\mathcal{X}$} in the main text.
-A more general notion using position-specific domains (e.g., ${\mathcal{X}}_{i}$’) can be used; see Sec.
+[^1] For brevity, we use a shared token set $\mathcal{X}$} in most parts of this paper.
+A more general notion using position-specific domains (e.g., ${\mathcal{X}}_{i}$’) can be used; e.g., see Sec.
 F.1.
@@
 Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{\left|I\right|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ to be an $I$-co-occurrence of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ if it satisfies $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset S$ and ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[\ell]\setminus I}={\bm{x}}^{\prime}_{[\ell]\setminus I}$.
-Also, the subsequences $\bm{a}$} and $\bm{a}$$^{\prime}$ are said to be functionally $k$-equivalent at $I$ in $S$ and denoted by ${\bm{a}}\equiv^{I}_{S}{\bm{a}}^{\prime}$, if it satisfies:
+Also, the subsequences $\bm{a}$} and $\bm{a}$$^{\prime}$ are said to be functionally $k$-equivalent at $I$ in $S$ and denoted by ${\bm{a}}\equiv^{(D,k)}_{I}{\bm{a}}^{\prime}$, if it satisfies:
   1. 1.
@@
 To formalize this, we define substitution graph: Let ${\mathcal{G}}^{(D,k)}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{\ell}$ of all possible inputs.
-Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[\ell]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$.
+Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[\ell]$ such that ${\bm{x}}_{I}\equiv^{(D,k)}_{I}$ $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$.
 This process is illustrated on the right side of Fig.˜1, as a special case where $k=1$.
@@
 The $k$-coverage of $D\subset{\mathcal{X}}^{\ell}$, denoted by $\mathrm{Cover}_{k}(D)$, is the set of all inputs in $\mathcal{X}$$^{\ell}$ connected to an ${\bm{x}}\in D$ with a path in the substitution graph ${\mathcal{G}}^{(D,k)}$.
-Note that the notion of coverage is a stricter condition of the canonical definition of in-domain (ID), which is obtained by random train/test split (wang2024grokked) or taking combinations of observed internal computations (dziri2024faith).
+Note that the notion of coverage is a stricter condition of the canonical definition of in-domain (ID), which is obtained by random train/test split (Wang et al., 2024a) or taking combinations of observed internal computations (Dziri et al., 2023).
 In Sec.˜5, we demonstrate that learners may not necessarily generalize on data that are classified as ID in a canonical sense, but coverage can precisely explain when and why this occurs.
 We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms, and we demonstrate that the predictions made by our framework are invariant across model architecture and scale in Sec.˜6 and 7.
-Finally, $k$-coverage can be algorithmically determined for any fixed-length discrete sequence tasks (Alg.
-1), which we use for the analyses in the following sections.
+Finally, $k$-coverage can be algorithmically determined for any fixed-length discrete sequence tasks (Alg.˜1), which we use for the analyses in the following sections.
 Now, we formally define pattern matching as a kind of generalization that is done by substituting functionally $k$-equivalent fragments of inputs, whose boundary is precisely the $k$-coverage defined above.
@@
 We construct four synthetic tasks with different structures: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig.˜2).
-To isolate functional-equivalence-based generalizations, we create random mappings from a product space of token sets to control the generalization sources not attributable to compositional structures (i.e., commutativity).
+To isolate functional-equivalence-based generalizations, we randomly create ground-truth mappings from a product space of token sets to control the generalization sources not attributable to compositional structures (i.e., commutativity).
 We explain the dataset construction process using 2-Hop task (Fig.˜2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example.
-We construct training datasets by defining a token set with size $\left|{\mathcal{X}}\right|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$.
+We construct training datasets by defining a token set with size $\left|{\mathcal{X}}\right|$ and randomly defining two mappings for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$.
 We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset.
@@
 Training & evaluation.
-Following wang2022interpretability, we train randomly initialized GPT-2 (radford2019language) models with 8 layers, 12 heads, and 768 dimensions as a base model (see Sec.˜B.2 for details).
+Following Wang et al.
+(2023), we train randomly initialized GPT-2 (Radford et al., 2019) models with 8 layers, 12 heads, and 768 dimensions as a base model (see Sec.˜B.2 for details).
 We construct two evaluation sets, each with 2,000 instances: (1) ID Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{3})$ in 2-Hop task) are observed during training, but their specific combination was unseen.
@@
 Then, we analyze what fraction of ID test data of 2-Hop task with $\left|{\mathcal{X}}\right|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$.
-Fig.˜3 (Left) shows that at N=5N=5k, every ID test example is already covered with minimal evidence ($k=1$).
+Fig.˜3 (Left) shows that, at N=5N=5k, every ID test example is already covered with minimal evidence ($k=1$).
 Hence, in an ideal scenario where a single witness of functional equivalence suffices, training with the dataset as small as N=5N=5k will lead to perfect ID generalization.
@@
 The parameter $k$ effectively quantifies this evidence strength, directly impacting generalization speed and reliability.
-Second, while our experiments use uniformly sampled datasets, the results can explain why models struggle with generalizing long-tail distributions in imbalanced real-world data (mallen2022not; kandpal2023large; chang2024large).
+Second, while our experiments use uniformly sampled datasets, the results can explain why models struggle with generalizing long-tail distributions in imbalanced real-world data (Mallen et al., 2023; Kandpal et al., 2023; Chang et al., 2024).
 Despite technically being in-distribution, rare combinations naturally receive limited evidence of functional equivalence (low $k$), effectively and practically placing them outside the coverage.
@@
 The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually.
-We verify that the representation clusters play a causal role in pattern matching with causal tracing (goldowsky2023localizing; hanna2023does), a widely used technique to identify Transformer circuits (Fig.˜8).
-Our findings extend the previous insights from mechanistic interpretability literature (wang2024grokked) in several ways.
+We verify that the representation clusters play a causal role in pattern matching with causal tracing (Goldowsky-Dill et al., 2023; Hanna et al., 2023), a widely used technique to identify Transformer circuits (Fig.˜8).
+Our findings extend the previous insights from mechanistic interpretability literature (Wang et al., 2024a) in several ways.
 First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps.
-Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens nostalgebraist2020logitlens may fail to detect these functional equivalence representations despite their presence (see App.˜E).
+Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens (nostalgebraist, 2020) may fail to detect these functional equivalence representations despite their presence (see App.˜E).
 6 Data Scaling Law of Pattern Matching Behaviors
@@
 Intuitively, this requires the training set to support (strongly enough) the functional equivalence of every pair of inputs that shares the same intermediate state $b$.
-Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$.
+Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{(D,k)}_{\{1,2\}}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$.
 Assuming that generalization is constrained by the $k$-coverage, how should the train data size scale in the token set size $\left|{\mathcal{X}}\right|$ to achieve it completely?
@@
 For a uniformly randomly sampled train dataset $D$ of size $N$, consider a learner that generalizes within the $k$-coverage of $D$.
-Then, for large enough $n$, the learner achieves perfect ID generalization with high probability if $N\!\gtrsim\!n^{c}$ with $c=2.5-\frac{0.5}{k}$.
-In contrast, the learner (with $k\!\geq\!2$) does not achive perfect ID generalization with high probability for some 2-Hop task if $n^{2}\!\lesssim\!N\!\lesssim\!n^{c}$.
+Then, for sufficiently large $n$, the learner can achieve perfect ID generalization with high probability if $N\!\gtrsim\!n^{c}$ with $c=2.5-\frac{0.5}{k}$.
+In contrast, the learner (with $k\!\geq\!2$) cannot achieve perfect ID generalization with high probability for some 2-Hop task if $n^{2}\!\lesssim\!N\!\lesssim\!n^{c}$.
 Here, we ignore the polylogarithmic factors in $n$.
-Theorem 6.1 presents a tight sample complexity bound $\tilde{\Theta}(n^{c})$ for pattern-matching-based generalization, showing that the training dataset required to ensure full ID generalization grows polynomially in the token set size $n$, with an exponent $c\in[2,2.5)$.
+Theorem˜6.1 presents a tight sample complexity bound $\tilde{\Theta}(n^{c})$ for pattern-matching-based generalization, showing that the training dataset required to ensure full ID generalization grows polynomially in the token set size $n$, with an exponent $c\in[2,2.5)$.
 To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(\left|{\mathcal{X}}\right|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data (see App.˜G for the measurement details).
@@
 G).
-We note that our result aligns with the practical observation that parameter scaling does not significantly improve the multi-hop reasoning capability of LLMs (yang2024large) and the data-hungry nature of compositional tasks (lake2018generalization), suggesting that these could be partly attributed to pattern-matching behaviors.
+We note that our result aligns with the practical observation that parameter scaling does not significantly improve the multi-hop reasoning capability of LLMs (Yang et al., 2024) and the data-hungry nature of compositional tasks (Lake and Baroni, 2018), suggesting that these could be partly attributed to pattern-matching behaviors.
 We leave further analysis of the connection between scaling behavior and pattern matching as an exciting future research direction.
@@
 In contrast, grouping by $x_{2}$-conditioned intermediate state $(b,x_{2})$ leads to high IICG scores, showing the formation of context-dependent state representations.
-This context-dependence due to path ambiguity raises an interpretability concern, as standard linear probing-based techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) would not reliably identify intermediate states when a model relies on pattern matching.
-Hence, a generalization mechanism other than pattern matching will be required for a robust ID generalization on complex task structures that require the access and update of intermediate states through multiple paths (e.g., planning tasks (ruis2020benchmark; kambhampati2024position; valmeekam2023planbench)), where further characterization of this problem remains as an exciting future direction.
+This context-dependence due to path ambiguity raises an interpretability concern, as standard linear probing-based techniques like logit lens (nostalgebraist, 2020; Belrose et al., 2023) would not reliably identify intermediate states when a model relies on pattern matching.
+Hence, a generalization mechanism other than pattern matching will be required for a robust ID generalization on complex task structures that require the access and update of intermediate states through multiple paths (e.g., planning tasks (Ruis et al., 2020; Kambhampati et al., 2024; Valmeekam et al., 2023)), where further characterization of this problem remains as an exciting future direction.
 8 CoT Improves Data Efficiency but Path Ambiguity Persists
@@
 Models are trained for 100 epochs after reaching training accuracy > 0.99.
-CoT supervision (wei2022chain; kojima2022large) dramatically improves performance on multi-step reasoning tasks.
-We investigate how CoT interacts with our framework and whether it can address the challenges observed in Sections 6 and 7.
+The CoT supervision (Wei et al., 2022; Kojima et al., 2022) dramatically improves performance on multi-step reasoning tasks.
+We investigate how it interacts with our framework and whether it can address the challenges observed in Sec.˜6 and 7.
 Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example.
-This substantially improves data efficiency (Fig.˜7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in the 3-Hop task, aligning with previous studies on the sample efficiency of CoT (srivastava2022beyond; kim2024transformers; wen2024sparse).
+This substantially improves data efficiency (Fig.˜7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in the 3-Hop task, aligning with previous studies on the sample efficiency of CoT (Srivastava et al., 2023; Kim and Suzuki, 2025; Wen et al., 2025).
 The scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision.
@@
 Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec.˜7.
-Our analysis may explain why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data (stechly2024chain), where we leave further analysis as future work.
+Our analysis may explain why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data (Stechly et al., 2024), where we leave further analysis as future work.
 9 Discussion
@@
 Although the experiments in this work have focused on synthetic setups to dissect purely pattern-matching behaviors, our results provide valuable practical implications for understanding and improving LLMs on natural language tasks.
-First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization (lake2017building).
-For instance, our scaling-law analysis (Theorem˜6.1) provides a quantitative explanation for the observed behavior that the data demand for generalization on multi-hop natural language data dramatically increases with the number of hops (yao-etal-2025-language-models).
-Such a data-hungry nature of compositional tasks is also well observed in semantic parsing (dong-lapata-2016-language), where it has been shown that the diversity of data (i.e., the same component is shown in diverse contexts) is more important than the sheer size of the dataset for better generalization (keysers2019measuring).
-Based on these insights, we believe our results will motivate strategic data augmentation methods that seek to maximize coverage by ensuring diverse shared contexts for functionally equivalent components (andreas2019good).
-Second, our framework helps interpret various real-world failure cases observed in LLMs as originating from their reliance on pattern matching.
-For example, our framework aligns with the reversal curse phenomenon, which is the failure to automatically infer that ‘$A$ is a child of $B$’ by observing that ‘$B$ is a parent of $A$’ (berglund2023reversal).
-This is because pattern matching is fundamentally incapable of generalizing to the reversing relation without explicit functional equivalence evidence for that in the training dataset.
-Similarly, our framework can explain the notoriously hard problem of logical negation in LLMs (truong-etal-2023-language).
+First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization (Lake et al., 2017).
+For instance, our scaling-law analysis (Theorem˜6.1) provides a quantitative explanation for the observed behavior that the data demand for generalization on multi-hop natural language data dramatically increases with the number of hops (Yao et al., 2025).
+Such a data-hungry nature of compositional tasks is also well observed in semantic parsing (Dong and Lapata, 2016), where it has been shown that the diversity of data (i.e., the same component is shown in diverse contexts) is more important than the sheer size of the dataset for better generalization (Keysers et al., 2020).
+Based on these insights, we believe our results will motivate strategic data augmentation methods that seek to maximize coverage by ensuring diverse shared contexts for functionally equivalent components (Andreas, 2020).
+Second, our framework helps interpret various real-world failure cases observed in LLMs as stemming from their reliance on pattern matching.
+For example, our framework aligns with the reversal curse phenomenon, which is the failure to automatically infer that ‘$A$ is a child of $B$’ by observing that ‘$B$ is a parent of $A$’ (Berglund et al., 2024).
+This occurs because pattern matching inherently cannot generalize to the reversing relation without clear evidence of functional equivalence in the training dataset.
+Similarly, our framework can explain the notoriously hard problem of logical negation in LLMs (Truong et al., 2023).
 A purely pattern-matching learner cannot deduce that the negation rule (i.e., if a statement ‘$p$’ is true, then ‘not $p$’ is false and vice versa) should apply to any well-formed statement, including the ones that aren’t observed, through finite observations.
-Finally, difficulties in solving complex planning tasks with LLMs (valmeekam2023planbench; stechly2024chain; kambhampati2024position; wang2024planning) might be partially attributed to path ambiguities, since such tasks likely require the correct tracking of intermediate states, which can be affected by various computational paths.
+Finally, difficulties in solving complex planning tasks with LLMs (Valmeekam et al., 2023; Stechly et al., 2024; Kambhampati et al., 2024; Wang et al., 2024b) may partly result from path ambiguities, since such tasks likely require the correct tracking of intermediate states, which can be affected by multiple computational paths.
 Hence, we believe our framework and analyses can provide important practical insights, which future work can extend to investigate LLMs’ pattern-matching behaviors under more realistic setups.
 9.2 Towards a Taxonomy of Generalization Mechanisms
-Natural language tasks possess algebraic and structural properties that differ from the random mappings we studied.
+Natural language tasks exhibit algebraic and structural properties that distinguish them from the randomly generated mappings we studied.
 For example, the same knowledge can be used in any hop of the multi-hop reasoning in practice, unlike our 2-Hop task that used $f_{1}\neq f_{2}$.
@@
   • Shared-operator generalization leverages the reuse of the same computation across positions (e.g., when $f_{1}=f_{2}$ in a two-hop task), which may be important in compositional generalization.
-For example, it is known that Transformers with inductive biases towards computation reuse can improve generalization on compositional tasks (csordas2021devil).
+For example, it is known that Transformers with inductive biases towards computation reuse can improve generalization on compositional tasks (Csordás et al., 2021).
 We envision this taxonomy as a foundational diagnostic that quantifies when pattern matching suffices and when other mechanisms are required.
@@
 In this work, we formalized a framework for characterizing pattern matching.
-Our theoretical and experimental analyses yielded quantitative and predictive insights into modern neural networks’ pattern-matching behaviors, moving beyond post-hoc accounts of a model’s behavior on compositional tasks: (i) the alignment of instance-wise success with the strength of functional-equivalence evidence (Sec.˜5), (ii) the theoretical identification and empirical verification of a sharp sample complexity bound for complete ID generalization on the 2-Hop task through pattern matching (Sec.˜6), and (iii) the identification of the path ambiguity problem that impairs accuracy and interpretability even under high coverage and CoT supervision (Sec.˜7 and 8).
+Our theoretical and experimental analyses provided quantitative and predictive insights into modern neural networks’ pattern-matching behaviors, moving beyond post-hoc accounts of a model’s behavior on compositional tasks: (i) the alignment of instance-wise success with the strength of functional-equivalence evidence (Sec.˜5), (ii) the theoretical identification and empirical verification of a sharp sample complexity bound for complete ID generalization on the 2-Hop task through pattern matching (Sec.˜6), and (iii) the identification of the path ambiguity problem that impairs accuracy and interpretability even under high coverage and CoT supervision (Sec.˜7 and 8).
 We anticipate that future work will build on this foundation, towards a more complete and constructive understanding of compositional generalization and its failures.
@@
 Acknowledgements
-We thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
-Contents of the Appendix
+We thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their invaluable feedback.
+This work was supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)).
+Hoyeon Chang, Jinho Park, Miyoung Ko, Hyeonbin Hwang, & Minjoon Seo acknowledge support by the InnoCORE program of the Ministry of Science and ICT(N10250156).
+Hanseul Cho acknowledges support by IITP grant funded by the Korean government (MSIT) (No.
+RS-2024-00457882, National AI Research Lab Project).
 Appendix A Limitations
@@
 Function Construction
-For the 2-Hop task, we construct two primitive functions $f_{1}:{\mathcal{X}}^{2}\rightarrow{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\rightarrow{\mathcal{X}}$ by randomly mapping from their respective domains to the codomain $\mathcal{X}$}.
+For the 2-Hop task, we construct two primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ by randomly mapping from their respective domains to the codomain $\mathcal{X}$}.
 We create the domain by taking the Cartesian product of the token set with itself.
@@
 Test Set Construction
-We carefully construct test sets to evaluate the model’s generalization capabilities across different coverage conditions. Our test sets contain:
+We carefully construct test sets to evaluate the model’s generalization capabilities across coverage conditions. Our test sets contain:
   • In-Domain (ID) Test Set: Combinations unseen during training but where both primitive operations were observed in other contexts.
@@
 Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600.
-The implementation follows the codebase from (wang2024grokked).
+The implementation follows the codebase from (Wang et al., 2024a).
 We train all models using the AdamW optimizer with beta values of (0.9, 0.999) and epsilon of 1e-8.
@@
 Appendix C Implementation Details for the Coverage Determination Algorithm
-Input: Training examples $D=\{(x_{i},f(x_{i}))\}_{i=1}^{N}$, where $x_{i}\in{\mathcal{X}}^{n}$ and $f(x_{i})\in{\mathcal{X}}$
-Minimum evidence threshold $k\geq 1$
-Output: Coverage set $\text{Cover}(D)$
+In this section, we explain our algorithm of computing the $k$-coverage $\mathrm{Cover}_{k}(D)$ (see Def.˜3.2), given a train dataset $D\subseteq{\mathcal{X}}^{\ell}$ and a threshold $k$ to check functional equivalences (see Def.˜3.1).
+The algorithm works in three main stages, as described with a pseudo-code in Alg.˜1.
+[…]
+Output: Coverage set: $\mathrm{Cover}_{k}(D)$
+1
 /* STEP 1: Build behavior maps for each subsequence pattern */
-foreach nonempty subset $I\subsetneq[n]$ do
- $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in D\}$
-end foreach
-/* STEP 2: Identify functionally equivalent subsequences */
-foreach nonempty subset $I\subsetneq[n]$ do
- $UF_{I}\leftarrow\text{new UnionFind}()$
- foreach pair of subsequences $(\alpha,\beta)$ in $\text{Behavior}_{I}$ do
-    $\text{SharedComplements}\leftarrow$ complements observed with both $\alpha$ and $\beta$
-    if No contradictions in SharedComplements and matching evidence $\geq k$ then
-       $UF_{I}.\text{Union}(\alpha,\beta)$ ;
-       // Mark as functionally equivalent
-      end if
-   end foreach
- $\text{EquivClasses}_{I}\leftarrow UF_{I}$
-end foreach
-/* STEP 3: Build substitution graph */
-$G\leftarrow\text{empty graph}$ with nodes for all $x\in{\mathcal{X}}^{n}$
-foreach pair of inputs $(x,y)$ with $f(x)=f(y)$ do
- foreach subset II where $x$ and $y$ differ only on indices in II do
-    if $\text{EquivClasses}_{I}.\text{Find}(x_{I})=\text{EquivClasses}_{I}.\text{Find}(y_{I})$ then
-         Add edge $(x,y)$ to $G$
-       break
-      end if
-   end foreach
-end foreach
-/* STEP 4: Determine coverage */
-$\text{Cover}(D)\leftarrow\bigcup_{x\in D}$ ConnectedComponent(G,x)(G,x)
-return $\text{Cover}(D)$
+2 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav$$\leftarrow$ an empty hash table
+3foreach $(I,I^{\mathsf{c}}):\varnothing\neq I\subsetneq[\ell],~I^{\mathsf{c}}=[\ell]\setminus I$ do
+4
+5 ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I]\leftarrow$ an empty hash table
+6 foreach ${\bm{x}}\in D$ do
+7    if ${\bm{x}}_{I}\notin{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I].keys$ then
+8       ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{x}}_{I}]\leftarrow$ an empty hash table
+9      end if
+10    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{x}}_{I}][{\bm{x}}_{I^{\mathsf{c}}}]\leftarrow f({\bm{x}})$
+11   end foreach
+12
+13 end foreach
+14
+/* STEP 2: Build substitution graph using behavior maps */
+15 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$$\leftarrow$ a graph (vertex set $\mathcal{X}$$^{\ell}$; edge set $\varnothing$)
+16foreach $(I,I^{\mathsf{c}}):\varnothing\neq I\subsetneq[\ell],~I^{\mathsf{c}}=[\ell]\setminus I$ do
+17 foreach $({\bm{a}},{\bm{a}}^{\prime}):{\bm{a}},{\bm{a}}^{\prime}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I].keys$  and  ${\bm{a}}\neq{\bm{a}}^{\prime}$ do
+    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}].keys~\cap~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}^{\prime}].keys$ // II-co-occurrences
+18
+19    if $|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}|\geq k$  and  $\forall{\bm{b}}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}][{\bm{b}}]={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}^{\prime}][{\bm{b}}]$   then
+       /* Statisfies the functional $k$-equivalence (Def.˜3.1) */
+20       foreach ${\bm{b}}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}$ do
+21            Let ${\bm{x}},{\bm{x}}^{\prime}\in{\mathcal{X}}^{\ell}$ such that ${\bm{x}}_{I}={\bm{a}},~{\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime},$ and ${\bm{x}}_{I^{\mathsf{c}}}={\bm{x}}^{\prime}_{I^{\mathsf{c}}}={\bm{b}}$
+22           Add an edge $\{{\bm{x}},{\bm{x}}^{\prime}\}$ to $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$}
+23         end foreach
+24
+25      end if
+26
+27   end foreach
+28
+29 end foreach
+30
+/* STEP 3: Build $k$-coverage */
+31 $\mathrm{Cover}_{k}(D)\leftarrow\bigcup_{{\bm{x}}\in D}\mathrm{ConnectedComponent}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}};{\bm{x}})$
+return $\mathrm{Cover}_{k}(D)$
 Algorithm 1 $k$-Coverage Determination Algorithm
-Alg.˜1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
-Stage 1: Behavior mapping
-We first analyze the training data to create a mapping of behaviors for each possible subsequence of the input.
-For each subset of indices $I$, we record how different subsequences $x_{I}$ behave when paired with their complements $x_{[n]\setminus I}$, essentially mapping each subsequence to a function from complements to outputs.
-Stage 2: Equivalence class construction
-For each subset of indices $I$, we build equivalence classes of subsequences that exhibit functionally identical behavior.
-Two subsequences are considered equivalent only if: (1) they share at least $k$ distinct complements where they produce the same output, and (2) they never produce different outputs when given the same complement (no contradictions).
-We use a Union-Find data structure to efficiently track and merge these equivalence classes.
-The Union-Find (or Disjoint-Set) data structure efficiently maintains a collection of disjoint sets, supporting two key operations: (1) Find - determine which set an element belongs to, and (2) Union - merge two sets.
-Stage 3: Substitution Graph Construction
-We construct a graph where nodes represent input sequences from our training and test sets, rather than the entire domain space (which would be computationally prohibitive for large token sets).
-We add an edge between two inputs $x$ and $y$ if and only if: (1) they produce the same output, (2) they differ only in one subsequence position set $I$, and (3) their differing subsequences belong to the same equivalence class.
-This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output.
-Our implementation uses parallel processing to efficiently construct this graph, even for large datasets.
-Stage 4: Coverage computation
-Finally, we compute the coverage set by taking the union of all connected components in the substitution graph that contain at least one training example.
-This set comprises all inputs that are reachable from the training data through chains of equivalent subsequence substitutions.
+In the first stage (Alg.˜1, Lines 11), we first create a mapping of behaviors for all subsequences of every training data.
+To elaborate on the behaviors, for each subset of indices $I$, we record how each subsequence ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ at $I$ of ${\bm{x}}\in D$ behave when being paired with another subsequence ${\bm{x}}_{I^{\mathsf{c}}}$ at the complement of indices $I^{\mathsf{c}}:=[\ell]\setminus I$.
+This ends up creating a thrice-nested mapping $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav$}, as showcased in Line 1.
+It is particularly useful in the next stage, where we should collect the co-occurrences of each pair of input sequences and verify their consistencies.
+The second stage (Alg.˜1, Lines 11) is for establishing the substitution graph ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}}={\mathcal{G}}^{(D,k)}$ which is essential in the definition of $k$-coverage.
+To recall, ${\mathcal{G}}^{(D,k)}$ is an undirected graph with vertex set $\mathcal{X}$$^{\ell}$, having edges $\{{\bm{x}},{\bm{x}}^{\prime}\}$ when and only when there exists an index set $I$ such that ${\bm{x}}_{I}\equiv^{(D,k)}_{I}{\bm{x}}^{\prime}_{I}$.
+To verify the presence of an edge via functional $k$-equivalence, we need to figure out at least $k$ distinct $I$-co-occurrences ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc$} in the pseudocode) for some index set $I$ (Def.˜3.1(1)) and check whether all $I$-co-occurrences satisfy the consistency (Def.˜3.1(2)).
+It is done for every nonempty index set $I\subsetneq[\ell]$ and every distinct pair of subsequences $({\bm{a}},{\bm{a}}^{\prime})\in{\mathcal{X}}^{|I|}\times{\mathcal{X}}^{|I|}$, where we take advantage of the mapping $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav$} defined in stage 1; see Line 1 in particular.
+In the last stage, we generate the $k$-coverage (Alg.˜1, Line 1).
+Here we utilize a subroutine ${\rm ConnectedComponent}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}};{\bm{x}})$ which returns the set of all vertices of $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$} sitting on the same connected component as the given vertex $\bm{x}$}.
+Hence, Line 1 generates the set of all input sequences that are connected with a train datapoint by a path in the substitution graph $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$}, which is precisely the definition of $k$-coverage.
+In other words, this set comprises all inputs that are reachable from the training data through a chain of functionally equivalent subsequence substitutions.
 Appendix D Detailed Analysis for Representation Unification Experiments
 D.1 Causal Tracing Methodology
-To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference (goldowsky2023localizing; hanna2023does).
-Specifically, we measure the causal effect using the indirect effect metric defined in (sharma2024locating).
-This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization.
-We illustrate the measurement with 2-Hop task.
+To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference (Goldowsky-Dill et al., 2023; Hanna et al., 2023).
+Specifically, we measure the causal effect using the indirect effect metric defined in (Sharma et al., 2024).
+This methodology enables us to identify which components and positions in the model most strongly contribute to compositional generalization.
+We illustrate the measurement with the 2-Hop task.
 Our analysis begins by collecting three types of computational traces:
@@
 To quantify the causal effect of a specific hidden state $h_{i}^{(\ell)}$ on the model’s prediction, we measure the Indirect Effect (IE):
-$\text{IE}_{h_{i}^{(\ell)}}=\frac{p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})-p(t^{*})}{p^{*}(t^{*})-p(t^{*})}$ (4)
+$\mathrm{IE}_{h_{i}^{(\ell)}}=\frac{p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})-p(t^{*})}{p^{*}(t^{*})-p(t^{*})}$
 where:
-  • $p(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the clean run $G$
-  • $p^{*}(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the corrupted run $G^{*}$
-  • $p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the patched run $G[\leftarrow{h^{*}}_{i}^{(\ell)}]$
+  • $p(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the clean run $G$;
+  • $p^{*}(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the corrupted run $G^{*}$;
+  • $p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the patched run $G[\leftarrow{h^{*}}_{i}^{(\ell)}]$.
 This metric quantifies how much corruption in a particular state affects the overall outcome.
@@
 D.2 Causal Tracing Results for Each $k$-Cutoff Value in 2-Hop Task
-Figure 8 displays the causal tracing results for the 2-Hop task, broken down by different $k$-cutoff values.
+Fig.˜8 displays the causal tracing results for the 2-Hop task, broken down by different $k$-cutoff values.
 We observe that the causal patterns are similar across different $k$-cutoff values, with slight differences in where and how strongly the causal effects manifest in the model.
@@
 We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$.
-For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
+For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\mathrm{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
 Figure Fig.˜9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases.
@@
 We show that GPT-2 models trained on Parallel-2-Hop and 3-Hop tasks exhibit the same patterns: clustered functional equivalence representations of intermediate states at specific layers and positions, confirmed through cosine similarity analysis, with causal tracing analysis verifying their role in model predictions.
-For both tasks, we analyze with $|{\mathcal{X}}|=50$ and examine model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
-Figures 11 and 12 show the results for the Parallel-2-Hop task.
+For both tasks, we analyze with $|{\mathcal{X}}|=50$ and examine model checkpoints with training dataset size $N=\hat{N}_{\mathrm{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
+Figs.˜11 and 12 show the results for the Parallel-2-Hop task.
 The IICG patterns reveal strong representation clustering at mid-layers: at positions $x_{2}$ and $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{4}$ when grouped by $b_{2}=f_{2}(x_{3},x_{4})$.
@@
 Right: perturbation with different $(x_{3},x_{4})$ pair leading to a different $b_{2}$ value.
-Similarly, Figures 13 and 14 show results for the 3-Hop task.
+Similarly, Figs.˜13 and 14 show results for the 3-Hop task.
 The IICG patterns exhibit strong representation clustering at mid-layers: at position $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{3}$ when grouped by $b_{2}=f_{2}(b_{1},x_{3})$.
@@
   • In Sec.˜F.3, we provide (almost) all postponed proofs of lemmas used in Sec.˜F.2.
-We first prove that the connectivity of all $b$-evidence graphs (Def.˜F.5), which are the intermediate-state-specific induced subgraphs of the substitution graph, is a sufficient condition for the perfect in-domain generalization we want to show: see Lemma˜F.10.
+We first prove that the connectivity of all $b$-evidence graphs (Def.˜F.5), the intermediate-state-specific induced subgraphs of the substitution graph, is a sufficient condition for the perfect in-domain generalization we want to show: see Lemma˜F.10.
     • Sec.˜F.3.1 is dedicated to explaining the renowned Poissonization technique, rigorously proving one of our main lemmas, Lemma˜F.12.
@@
 Given a train dataset $D$ and for each ${\bm{x}}\in{\mathbb{X}}$, define $Z_{\bm{x}}$ as the sampled count of $\bm{x}$} (i.e., the number of identical duplicates equal to $\bm{x}$}) in $D$.
-Then, if $D$ is generated by with-replacement sampling $N$ times from $\mathbb{X}$}, the random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ follows a multinomial distribution with parameters $n=N$ and $p_{\bm{x}}=\frac{1}{\left|{\mathbb{X}}\right|}$ ($\forall{\bm{x}}\in{\mathbb{X}}$).
+In particular, if $D$ is uniformly randomly generated by with-replacement sampling $N$ times from $\mathbb{X}$}, the random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ follows a multinomial distribution with parameters $n=N$ and $p_{\bm{x}}=\frac{1}{\left|{\mathbb{X}}\right|}$ ($\forall{\bm{x}}\in{\mathbb{X}}$).
 That is, for all collections of integers $z_{\bm{x}}\geq 0$ ($\forall{\bm{x}}\in{\mathbb{X}}$) such that $\sum_{{\bm{x}}\in{\mathbb{X}}}z_{\bm{x}}=N$,
@@
 The Poissonization technique we use is summarized as the lemma below, which will be proved in Sec.˜F.3.1:
-Lemma F.12 (De-Poissonizaiton of Monotone Multinomial Events, $\downarrow$).
+Lemma F.12 (De-Poissonization of Monotone Multinomial Events, $\downarrow$).
 Fix any $n\geq 1$. Define
@@
 In particular, we will use only the $k\geq 2$” part (which tightly matches our assumption $m=\Omega(n_{b})$ ($\forall b\in B_{D}$)) below:
-Lemma F.14 (zhao2014k, Theorem 2; zhao2017connectivity, Theorem 1 & Remark 1).
+Lemma F.14 (Zhao et al., 2014, Theorem 2; Zhao et al., 2017, Theorem 1 & Remark 1).
 Fix any $k\geq 1$. Suppose that
@@
 Then, choosing all $n_{b}=\left|V_{b}\right|$ to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the bound (8).
-Note that a choice of $p$ larger than the threshold in Eq.˜8 will never change that ${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$.
+Note that a choice of $p$ larger than the threshold in Eq.˜7 will never change that ${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$.
 Thus, since we choose $\alpha_{n}=\ln n$, it suffices to have
@@
 Namely, we again observe that every ${\mathcal{G}}_{b}^{(D,1)}$ is a binomial random $1$-intersection graph ${\mathcal{G}}^{(1)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$ (see the definition of binomial random intersection graph in Sec.˜F.3.2).
-To guarantee the union bound in Eq.˜9, however, we use a different lemma: Lemma˜F.15 (instead of Lemma˜F.14).
+To guarantee the union bound in Eq.˜8, however, we use a different lemma: Lemma˜F.15 (instead of Lemma˜F.14).
 In particular, we will use only the $\rho\leq 1$” part (which tightly matches our assumption $m=O(n_{b})$ ($b\in B_{D}$)) of the lemma below:
-Lemma F.15 (singer1995random, Propositions 3.1–2, Theorem 3.3).
+Lemma F.15 (B., 1995, Propositions 3.1–2, Theorem 3.3).
 Let $k=1$. Suppose that $m=n^{\rho}$ for $\rho>0$ and
@@
 $D_{B3}=B\times{\mathcal{X}}_{3}$ $\iff U_{w}\geq 1\;\;(\forall w\in B\times{\mathcal{X}}_{3}).$
-Hence, to show Eq.˜10, it suffices to prove the following three inequalities:
+Hence, to show Eq.˜9, it suffices to prove the following three inequalities:
 ${\mathbb{P}}_{(Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\sim\mathrm{Multinomial}(N;(p_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}})}\left((Z_{{\bm{x}}})_{{\bm{x}}\in{\mathbb{X}}}\in{\mathcal{M}}_{1}\right)$ $\leq\frac{\delta}{2};$ (11)
@@
 ${\mathbb{P}}_{(U_{w})_{w\in B\times{\mathcal{X}}_{3}}\sim\mathrm{Multinomial}(N;(r_{w})_{w\in B\times{\mathcal{X}}_{12}})}\left(\exists w\in B\times{\mathcal{X}}_{3}\ \text{ such that }\ U_{w}=0\right)$ $\leq\frac{\delta}{4}.$ (13)
-A size $N$ of training dataset $D$ that is sufficient to ensure Eqs.˜13 and 14 can be characterized with the following lemma.
+A size $N$ of training dataset $D$ that is sufficient to ensure Eqs.˜12 and 13 can be characterized with the following lemma.
 Lemma F.19 (A Tail Bound for Coupon Collector’s Problem, $\downarrow$).
@@
 $N\geq\max\left\{\left|{\mathcal{X}}_{12}\right|\ln\left(\frac{4\left|{\mathcal{X}}_{12}\right|}{\delta}\right),\;\frac{\left|{\mathbb{X}}\right|}{\hat{n}}\ln\left(\frac{4\left|B\!\times\!{\mathcal{X}}_{3}\right|}{\delta}\right)\right\}$ (14)
-is enough to ensure both Eqs.˜13 and 14. The remaining task now is to determine a condition for $N$ that guarantees Eq.˜12.
-Observe that ${\mathcal{M}}_{1}$ (defined as Eq.˜11) is a non-decreasing property (Def.˜F.11) because of a similar reason for the monotonicity of the property $\mathcal{C}$} (defined as Eq.˜7; see the proof of Theorem˜F.7).
+is enough to ensure both Eqs.˜12 and 13. The remaining task now is to determine a condition for $N$ that guarantees Eq.˜11.
+Observe that ${\mathcal{M}}_{1}$ (defined as Eq.˜10) is a non-decreasing property (Def.˜F.11) because of a similar reason for the monotonicity of the property $\mathcal{C}$} (defined as Eq.˜6; see the proof of Theorem˜F.7).
 Thanks to the monotonicity of ${\mathcal{M}}_{1}$, we can apply the Poissonization technique (Lemma˜F.12) once again.
@@
 Poisson random variables $Z_{\bm{x}}\sim\mathrm{Poisson}\left(\frac{N+{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)$ ($\forall{\bm{x}}\in{\mathbb{X}}$).
-Let us take ${\varepsilon}=\sqrt{\frac{6N}{3-c}\ln\left(\frac{4}{\delta}\right)}$, which is smaller than $N$ by Eq.˜15 (since $\left|{\mathcal{X}}_{12}\right|\geq 6>\frac{6}{3-c}$) and ensures that
+Let us take ${\varepsilon}=\sqrt{\frac{6N}{3-c}\ln\left(\frac{4}{\delta}\right)}$, which is smaller than $N$ by Eq.˜14 (since $\left|{\mathcal{X}}_{12}\right|\geq 6>\frac{6}{3-c}$) and ensures that
 $\exp\left(-\frac{(3-c){\varepsilon}^{2}}{6N}\right)=\exp\left(-\ln\left(\frac{4}{\delta}\right)\right)=\frac{\delta}{4}.$
@@
 Then, choosing $\hat{n}$ greater than $n_{0}\left(\frac{\delta}{4}\right)$, we yield the bound (15).
-Observe that a choice of $p$ smaller than the threshold in Eq.˜8 will never change the limit ${\mathbb{P}}(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1)\to 0$.
+Observe that a choice of $p$ smaller than the threshold in Eq.˜7 will never change the limit ${\mathbb{P}}(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1)\to 0$.
 Thus, by our choice $\alpha_{n}=-(1-\frac{1}{k})\ln n$, it is enough to have
@@
 Before moving on to the postponed proofs of lemmas, we lastly remark that the same proof (of Theorem˜F.16) can hardly apply to the case of $k=1$ in general.
-This is because, for the sake of simplicity in applying the necessary condition for a perfect coverage (Lemma˜F.18), we first characterize a minimal data size to guarantee $\overline{D}={\mathbb{X}}$ with high probability as Eq.˜15, using the tail bound of the coupon collector’s problem (Lemma˜F.19).
+This is because, for the sake of simplicity in applying the necessary condition for a perfect coverage (Lemma˜F.18), we first characterize a minimal data size to guarantee $\overline{D}={\mathbb{X}}$ with high probability as Eq.˜14, using the tail bound of the coupon collector’s problem (Lemma˜F.19).
 Unfortunately, it already exceeds the sample complexity upper bound to ensure $\overline{D}\subset\mathrm{Cover}_{1}(D)$ (obtained as Theorem˜F.8) for $k=1$ with high probability, especially in the regime of balanced cardinalities assumed in Corollary˜F.17.
@@
 In both cases above, we obtain the same result that ${\mathbb{X}}\setminus\mathrm{Cover}_{k}(D)\neq\varnothing$. It concludes the proof of the lemma. ∎
-We also prove the tail probability bound for the coupon collector’s problem (newman1960double; erdHos1961classical) here. We again restate the lemma here for the sake of readability.
+We also prove the tail probability bound for the coupon collector’s problem (Newman, 1960; Erdős and Rényi, 1961) here. We again restate the lemma here for the sake of readability.
 See F.19
@@
 Moreover, to guarantee the tail bounds to be small, we often choose $N$ having a lot of mass concentrated around its mean (e.g., sub-Gaussian); then, we can take advantage of concentration inequalities associated with $N$ to have a better estimate of $a_{n}\approx\mathbb{E}[a_{N}]$.
-Poissonization (kac1949deviations; holst1986birthday; aldous1989probability; jacquet1998analytical; johansson1998longest; borodin2000asymptotics) can be thought of as a special case of randomization using Poisson random variables.
+Poissonization (Kac, 1949; Holst, 1986; Aldous, 1989; Jacquet and Szpankowski, 1998; Johansson, 1998; Borodin et al., 2000) can be thought of as a special case of randomization using Poisson random variables.
 A Poisson random variable $N\sim\mathrm{Poisson}(\lambda)$ with a parameter $\lambda>0$ is equipped with a probability mass function
@@
 The only things left to obtain a complete (de)-Poissonization lemma for upper/lower-bounding the sequence $a_{n}$ are the tail probability bounds for a Poisson distribution.
-The concentration inequalities for the Poisson distribution are already well-known, [^8] For example, see a document online: “A short note on Poisson tail bounds.” although we derive slightly different forms of them for our own purposes.
+The concentration inequalities for the Poisson distribution are already well-known, [^8] For example, see a document online: “A short note on Poisson tail bounds.” although we derive slightly different forms of them for our own purpose.
 Lemma F.26 ($\downarrow$).
@@
 ${\mathbb{P}}_{\mathrm{Multi}(n)}(Z_{1}=z_{1},\cdots,Z_{m}=z_{m})={\mathbb{P}}_{\mathrm{Po}(\lambda)}\left(Z_{1}=z_{1},\cdots,Z_{m}=z_{m}\;\middle|\;\sum\nolimits_{i=1}^{m}Z_{i}=n\right).$ (16)
-Moreover, it is sufficient for proving the first equation to study the case with non-negative integers $z_{1},\cdots,z_{m}$ such that $\sum_{i=1}^{m}z_{i}=n$; otherwise, both sides of Eq.˜17 are zero.
+Moreover, it is sufficient for proving the first equation to study the case with non-negative integers $z_{1},\cdots,z_{m}$ such that $\sum_{i=1}^{m}z_{i}=n$; otherwise, both sides of Eq.˜16 are zero.
 In this case, observe an inclusion between events
@@
 $={\mathbb{P}}_{\mathrm{Multi}(n)}(Z_{1}=z_{1},\cdots,Z_{m}=z_{m}).$
-Therefore, we have just proved Eq.˜17.
+Therefore, we have just proved Eq.˜16.
 […]
@@
 $={\mathbb{P}}_{\mathrm{Multi}(n)}\left((Z_{1},\cdots,Z_{m})\in{\mathcal{A}}\right).$
-This proves Eq.˜19, concluding the proof. ∎
+This proves Eq.˜18, concluding the proof. ∎
 Proof of Lemma˜F.26.
@@
 F.3.2 Binomial Random Intersection Graphs
-Let us explain the random $k$-intersection graph (singer1995random; karonski1999random; godehardt2003two).
+Let us explain the random $k$-intersection graph (B., 1995; Karoński et al., 1999; Godehardt and Jaworski, 2003).
 Consider a set $V$ of vertices and another set $W$ of items.
@@
 $p={\mathbb{P}}_{\mathrm{Po}(\lambda)}(Z_{(x_{12},x_{3})}\geq 1)=1-{\mathbb{P}}_{\mathrm{Po}(\lambda)}(Z_{(x_{12},x_{3})}=0)=1-\exp\left(-\frac{\lambda}{\left|{\mathbb{X}}\right|}\right).$
-In addition, observe that the set $S(x_{12},x^{\prime}_{12}\mid D)$ defined in Eq.˜6 is identical to the intersection $W_{x_{12}}\cap W_{x^{\prime}_{12}}$.
-Hence, by the definition of the edge set $E_{b}^{(D,k)}$ in Eq.˜5, two distinct vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if and only if $\left|W_{x_{12}}\cap W_{x^{\prime}_{12}}\right|\geq k$.
+In addition, observe that the set $S(x_{12},x^{\prime}_{12}\mid D)$ defined in Eq.˜5 is identical to the intersection $W_{x_{12}}\cap W_{x^{\prime}_{12}}$.
+Hence, by the definition of the edge set $E_{b}^{(D,k)}$ in Eq.˜4, two distinct vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if and only if $\left|W_{x_{12}}\cap W_{x^{\prime}_{12}}\right|\geq k$.
 In summary, every $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ is a binomial random $k$-intersection graph ${\mathcal{G}}^{(k)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-\lambda/\left|{\mathbb{X}}\right|)$.
@@
 Indeed, it is often studied in the literature on random graphs that a sufficient condition for the parameters (in terms of $n\to\infty$) to guarantee certain graph properties, at least asymptotically.
-Here, we review a few of the seminal results on the connectivity (as well as the disappearance of isolated vertices) of binomial random $k$-intersection graphs (singer1995random; zhao2014k; zhao2017connectivity; rybarczyk2011sharp; rybarczyk2017coupling).
+Here, we review a few of the seminal results on the connectivity (as well as the disappearance of isolated vertices) of binomial random $k$-intersection graphs (B., 1995; Zhao et al., 2014; 2017; Rybarczyk, 2011; 2017).
 [^9] Caveat: in the literature of (random) graph theory, the letter ‘$k$’ is often used for $k$-connectivity (i.e., being connected after removing less than $k$ vertices/edges).
-On the other hand, they study ‘random $s$-intersection graphs,’ using $s$ as the intersection constraint parameter.
+On the other hand, they study ‘random $s$-intersection graphs,’ using ss—instead of $k$ as this paper—as the intersection constraint parameter.
 The proofs are involved; thus, we omit them.
@@
 For each token set size $\left|{\mathcal{X}}\right|$ and task structure, we test multiple dataset sizes until we identify the threshold point where the model successfully generalizes to the ID test set.
-Specifically, our criterion for “reliable generalization” on ID is defined as:
-  • The model must reach ID test accuracy of 0.99 within 100 epochs after achieving training accuracy > 0.99.
+Specifically, our criterion for the “reliable generalization” on ID is defined as: The model must reach ID test accuracy of 0.99 within 100 epochs after achieving training accuracy > 0.99.
 This protocol balances several considerations:
@@
     Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion.
-This establishes that our measurement error is at most $\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
+This establishes that our measurement error is at most $\log(0.75)\approx 0.125$ in log scale, providing confidence in the derived power-law exponents.
 G.2 Measured Power-Law Scaling Exponents Across Task Structures and Model Sizes
@@
 For each task structure, we vary the token set size $\left|{\mathcal{X}}\right|$ from 50 to 200, allowing us to observe the scaling relationship.
-Table 2 presents the power-law exponents obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs.
+Tab.˜2 presents the power-law exponents obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs.
 $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$.
@@
 $R^{2}>0.99$ for all linear fitting.
-To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations.
-Figure 16 demonstrates that for the 2-Hop task with $\left|{\mathcal{X}}\right|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
+To verify that the power-law scaling relationship we observed is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations.
+Fig.˜16 demonstrates that for the 2-Hop task with $\left|{\mathcal{X}}\right|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
   1. 1.
-    Learning rate reduction: Halving the learning rate from 8e-4 to 4e-4
+    Learning rate reduction: Halving the learning rate from 8e-4 to 4e-4;
   2. 2.
-    Weight decay reduction: Decreasing weight decay by a factor of 10 (from 0.1 to 0.01)
+    Weight decay reduction: Decreasing weight decay by a factor of 10 (from 0.1 to 0.01);
   3. 3.
-    Generalization criteria modification: Requiring test accuracy > 0.95 within 10 epochs after training accuracy > 0.95
+    Generalization criteria modification: Requiring test accuracy > 0.95 within 10 epochs after training accuracy > 0.95.
 This robustness to hyperparameter variations suggests that the power-law relationship between token set size and required dataset size is primarily a property of the compositional generalization process, rather than an artifact of specific optimization settings.
@@
 Type-II: Function property-based generalization.
-This mechanism exploits intrinsic properties of individual primitive functions, e.g., algebraic invariances such as commutativity or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present (wen2024sparse)).
+This mechanism exploits intrinsic properties of individual primitive functions, e.g., algebraic invariances such as commutativity or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present (Wen et al., 2025)).
 Unlike the previous type, this mechanism explains the generalization beyond the coverage by leveraging ‘global’ properties that hold across all possible inputs of a primitive, beyond what is actually observed.
-We interpret the Reversal Curse phenomenon (berglund2023reversal) as an example of the layered nature of challenges across multiple generalization types.
-Our framework predicts the failure of pattern matching on this problem, since training on “$A$ is BB” provides no functional equivalence evidence for “$B$ is $A^{-1}$”.
-An architectural modification to learn inverse mappings from the same training data to handle this problem (lv2024analysis) can be interpreted as a utilization of Type-II generalization to enable generalization beyond coverage.
+We interpret the Reversal Curse phenomenon (Berglund et al., 2024) as an example of the layered nature of challenges across multiple generalization types.
+Our framework predicts the failure of pattern matching on this problem, since training on “$A$ is BB” provides no functional equivalence evidence for “$B$ is AA”.
+An architectural modification to learn inverse mappings from the same training data to handle this problem (Lv et al., 2024) can be interpreted as a utilization of Type-II generalization to enable generalization beyond coverage.
 Type-III: Shared-operator generalization.
 This mechanism emerges through the reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$).
-Recurrent architectures (hochreiter1997long) exemplify the utilization of this through weight sharing across time steps, enabling processing of variable-length sequences (graves2014neural).
-Similarly, it has been reported in Transformers with inductive biases towards reuse of the same computation through parameter sharing (dehghani2018universal; csordas2021devil; wang2024grokked) can improve generalization on complex compositional tasks where the same primitive function can be reused in various contexts.
+Recurrent architectures (Hochreiter and Schmidhuber, 1997) exemplify the utilization of this through weight sharing across time steps, enabling processing of variable-length sequences (Graves et al., 2014).
+Likewise, it has been reported in Transformers with inductive biases towards reuse of the same computation through parameter sharing (Dehghani et al., 2019; Csordás et al., 2021; Wang et al., 2024a) can improve generalization on complex compositional tasks where the same primitive function can be reused in various contexts.
+With a similar insight, the generalization capability of Transformers in terms of sequence length (i.e., length generalization) and its limitation has recently been studied a lot, especially in terms of arithmetic/algorithmic tasks (Anil et al., 2022; Zhou et al., 2024a; Shen et al., 2023; Zhou et al., 2024b; Cho et al., 2024; 2025).
 We interpret this mechanism as exploiting structural repetition.
 Distinguishing mechanisms from phenomena.
-Compared to prior categorizations of generalization, which focus on observed phenomena (lake2018generalization; hupkes2020compositionality), we categorize the underlying mechanisms.
+Compared to prior categorizations of generalization, which focus on observed phenomena (Lake and Baroni, 2018; Hupkes et al., 2020), we categorize the underlying mechanisms.
 As noted in Sec.˜1, many behavioral studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure.